A Complete Bibliography of Publications in Theory of Probability and its Applications: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

10 June 2022 Version 3.00

Title word cross-reference

2 [BO07, GPL08, HP09, Sha05]. ^{0} [Leb03b]. 2 [KN02]. A [Yak08, PS04]. α [Zie01]. B [Shi02]. α ∈ (0, 1) [Leb08]. χ [PS04]. d [Der05, Sof02, Sof05]. f [Gus08]. G [BNPA03]. H [KPK06b]. H^{d} [KS07a]. I_{0} [Neu00]. L [BN00, KN02, Leb03b]. L^{p} [Fat03, Fat09, Leb02]. L_{2} [Jaj06, NN05]. L_{q} [Shi02]. Lu = ψu [DK01]. MAP/G/1/∞ [Pec01]. p [Khr02, KY05]. p > 0 [Fat09]. q [Shi02]. R^{d} [Ben05]. R^{S} [Kho06]. S [Mik03]. σ [Kal04, VG08]. SO(n) [Jan03]. t [BP03, NK09]. U [Sha03, AB02, Bor01, Eic01, Gad04]. ε [Gap02]. x ↦ |x − θ| [Mat03].

-Adic [Khr02, KY05]. -Ancillarity [BN00].

-CUSUM [HP09]. -Deviatations [Gap02]. -Dimensional [Der05]. -Divergence [Gus08]. -Fields [VG08]. -function [KPK06b]. -Functionals [Fat09]. -Localization [Kal04]. -Martingales [Kal04]. -Mixingale [Shi02]. -Nonformation [BN00]. -Norm [Fat03]. -Optimal [Sof02, Sof05]. -patterns [Mik03]. -Permutation [Yak08]. -Process [PS04]. -Processes [Eic01]. -Small [NN05]. -Smoothness [Shi02]. -Stable [Zie01]. -Statistic [BP03]. -Statistics [Sha03, AB02, Bor01, Gad04]. -step [BO07]. -Sufficiency [BN00]. -Upcrossings [PS04]. -Valued [Leb03b, Leb02, Shi02].

2 [BK01, Zan04].

3 [BK02b].

41-45 [Ano02b].

60th [Kha08a].

70th [Ano00c, Ano01a, Ano07f, Kha08b, tbYGS05]. 75th [ABV^{+}07, Ano06c, CIM^{+}07, Kha08c].

80th [Ano02a, Ano04e, Ano05e]. 85th [Kha09].

A. [CIM^{+}07, Klo07b, Kol04a, PS09, tbYGS05]. Abdullovich [Ano03a, The03]. Abel [Bor08]. Abelian [Bor04a, Fel01, FG05, Gab01, ZN04]. Absolute [Now05, She07]. Absolutely [BK06a, Roz04b]. Absorption [Kal03]. Academician [Ano07f]. Academy [Shi00]. Accidents [Mas07]. Accuracy [KS06a, KS06b, Mik02, Zhu00a]. Action [Web06, Web07]. Actions [Ver04]. Adaptive [CGLT04, RR05]. Additive [DB04, EL00]. Additivity [Hol07]. Adic [Khr02, KY05]. Adjoint [Shu01]. Adjusted [BD04]. after [Mas07]. Akiva [Ht09]. Alarm [Tar09]. Aleksandrovich [Ano02e]. Algebraic [RR05]. Allocation [Gaa08, Tim01, Tim05]. Almost [BR06b, BR06c, CF04, IL00, VD06]. along [BG04]. Alternative [MO01]. Alternatives [HM06, HP09]. American [Chr04]. Analogue [Sen08]. Analogues [Osm09, PVK02, PVK06]. Analogy [Mas04a]. Analysis [Bor07, GP05, GT07b, IM03, KO02, KR00, Nas07, Pas05, Rom01]. Analytic [Ibr07, KL08]. Anatolii [Ano01b]. Ancillarity [BN00]. Ancillary [KR06]. Anderson [HT07]. Andre [Ano04a]. Andrei [Ano02a]. Anisotropic [KLP08, Ost02]. Anniversaries [Ano05a]. Anniversary [Ano02f]. Antanovich [Ano00c, Ano05b]. Anticipation [GPS01]. Application [BK05, CPP02, Di 04, GP05, Lev01, NK09]. Applications [Bob03, BK06a, BF00, Kor01, Leb00b, PT09, RS01, Zhd09]. Approach [AS09, Bel05b, Dar05, Dok04, FK01, MO01, Mol03]. Approximate [KR04]. Approximates [Yar07]. Approximation [AS02, AS05, BB05a, BP03, Bor04a, Bor03b, Bor01, GZ09, HM08, KS06a, KS06b, KL05, LN00, LRT00, Mik02, MS07b, PS08, RS01, Roo01, Roo03, Roo04, Sun08, VM00, Zhu00a]. Approximations [CCHH08, Cek00, Roo02, Ste03, Ver03]. Arbitrage [Rok08]. Arbitrary [BF00]. Arcsine [Afa07, Kru06]. Arithmetic [Gap07, Zhd09]. ARMA [BS05]. Array [Sha06b]. Arrays [CHLV08, CHL02, HLRV03]. Articles [Shi00]. Artin [Nir03]. Asian [CS04]. Assisted [Hol04]. Associated [Amo05, CHLV08, Che01, CHL02]. Asymmetric [KP05, Nag07a]. Asymptotic [ASVV07, BY04, BR06a, Bog03, BM01a, BFS09, Chi02b, Chi02a, Chi03, Eng04, Fro03, FJT02, Hol05, KS07b, Kor00, Kor01, Kru01, KM00, MP04, Nad07, Nag07a, PT09, Ryb01a, Sen07, Sen09, Smo00, Sob08, TV05, Tar09, Yar09, vRRvZ00]. Asymptotical [Rag06b]. Asymptotically [Chi02b, Chi02a, Chi03, Erm04, HM06, Ser00, Sof02, Sof05]. Asymptotics [Bak00, Bor03a, Bor04b, Ćwi09, Fat03, Fat04, Fat09, GS07, NN05, Roz04a, Roz05, Tim01, Yak03, Yak04]. Attaining [Afa08]. Attains [Uru05]. Attraction [Gri00, Roz04a, Roz05]. Author [Ano02b]. Automorphically [Gab01]. Averages [ABM08, GW00, Gap07, Gap09, GT07b, PS08]. Averaging [LY06, LN00, Mas04b, Mas05]. Averse [GS07]. Axioms [Mas04b].

B [Ano05e]. B. [Ano04e, Kha09]. Bachelier [Kam09]. Bachelier-Version [Kam09]. Backward [BER05, BE06, BE08, CFS08]. Bahadur [KR01a]. Bakry [Bob03]. Ball [NN05]. Balls [Roz09b]. Banach [DM08, Fat04, HLRV03, Kva01, Mat00, RS05, RS00, Sha05, Shi02, VK07]. Banach-Valued [DM08, Fat04]. Bartlett [MS07a, Sta02]. Basis [Leb02, RR05]. Bayes [LS03]. Bayesian [Bur09, RS01, Ryb01a, Sol00, TV05, Tar09]. Behavior [BY04, CSY03, Fro03, KS08, Koz02, Kru01, LY06, Nad07, Rag06b, VD06, Zap06]. Belong [IK08a]. Below [Roz03a]. Benford [KP05, KPK06b]. Bernoulli [Kru05]. Berry [Gad04, She07]. Bessel [CS04, Che01, Mis06a]. Best [Chi02b, Chi02a, Chi03, HP09, KR01b]. Between [Afa04, Kon02, Zub00, Noq00, Now05]. Bias [BT08a, BT08b]. Bibliography [Ano00d, Ano00e, Ano00f, Ano00g, Ano01f, Ano01g, Ano02g, Ano02h, Ano02i, Ano03d, Ano03e, Ano03f, Ano04g, Ano06d, Ano07g, Ano07h]. Binomial [HM08, PVK02, PVK06, Roo01]. Birthday [ABV^{+}07, Ano00c, Ano01a, Ano02a, Ano02c, Ano04e, Ano05e, Ano06a, Ano06c, Ano07f, CIM^{+}07, Kha08b, Kha08c, Kha08a, Kha09, tbYGS05]. Block [CGLT04]. Books [Ano04c, Ano07c, Pan06b, Pan06a]. Bootstrap [BP03]. Bootstrapped [VCH06]. Border [Kal03]. Borovkov [CIM^{+}07]. Bottlenecks [VD07]. Bound [Ben05, GZ05, MPW06, She07]. Boundaries [IK08a]. Boundary [BF00, Bor03a, Bor04b, BM09a, BM09b, DK01, Fat03, Rag06a, RS00]. Boundary-Value [Fat03]. Bounded [Sha05]. Bounds [AB02, BRS06, CG04, GZ09, Jan03, Klo07a, Mat09, Nag02b, Nag02a, RR00, Yar09]. Branching [Afa04, Afa07, AK04, Bor03b, Fak09, FVW03, GK01, GDLH03, Kal02, Kal03, KS07a, KS08, Lan07, Leb06, Shi01, Vat03, VD04, Vat04, VD05, VT05, VD07, VD08, VD09].

Bridges [Lev02]. Brownian

[DW06, BFS09, CS04, DSY00, GPS01, GSY07, HM06, HP09, Lev02, LS04, Lob06b, Lob06a, Mol00, Mol03, OE09, SC00, SY04, Shi09c, Sui02, Uru05, Yar07]. Brunk [Kru03]. Brunswick [Nir03]. Burgers [Bak02]. Burkholder [Pre09].

Càdlàg [CFS08]. Calculus [DD00, GV07]. Calibration [GHKT07]. Call [PS02]. Call-Put [PS02]. Capacities [Hol04, Pug09a]. Capacity [HS06, Shi08, Shi09a]. Case [BO03, BR06b, BR06c, BK01, BK02b, KLP08, KP04, KPK06a, LB04, Pet03, Rok06, Rok08, Sen09, Vat03]. Cases [Eic01]. Catalytic [VT05]. Catastrophe [GP05]. Categorical [KK06]. Cells [Tim01, Tim05]. Censored [Ibr07]. Central [Min01, Bak02, BG04, GS05, Gri03, Joh06, Kam07, RS05, Roz03a, Sen07, Sen09, Sha06a, Sob08, Su02, Yar09]. Certain [Bor00, Kam07, KP02, Lou03, Nas00, Pin06, Sch03]. Chain [Bor03a, Bor04b, CCHH08, Rot08, She00, Sze06]. Chains [BO07, Lin00, Mit06, Shu01]. Chains.Part [BK01, BK02b]. Change [BNS06, Bor09b, BFS09, HM06, KS02, PT09, Sof02, Sof05, TV05]. Change-Point [Bor09b, Sof02, Sof05]. Changed [PKJ07]. Changepoint [Tar09]. Channels [AH03, Hol04, HS06, Hol07, Shi08, Shi09a]. Chaos [DDP07]. Characteristic [Gam01, KP03, Roz00c, Rya01]. Characteristics [DSY00, Lob06a, Osm09, Rot08]. Characterization [Ali00]. Charged [PSV00]. Chebyshev [Kon02, Osm09, Sen08]. Chemical [MMZ07]. Chernoff [PVK02, PVK06, Roz09a]. Chernoff-Type [Roz09a]. Chibisov [Kha08b]. Choice [KR01b]. Choosing [MO01]. Chover [Pin06]. Chover-type [Pin06]. Chung [PL07]. Chung-Type [PL07]. Clark [JP09]. Class [Dom02, HT07, Klo07a, KS07b, Lin00, MT01, NK09, Neu00, PT09, Rag06b, Ser06, Zan04]. Classes [Gap09, Kam07, Nas00, Web04]. Classical [Shi08]. Classification [Ryb01a]. Close [GPS01, Ing01]. Closed [Sch03]. CLT [Lou02, RR00, Web06, Web07]. Clusters [Vys06]. Coefficient [BDK05]. Coefficients [Chu00, IK00, IK01, Vla05]. Coherent [Che08b, Kul08]. Coinciding [Kly03]. Collision [KS07c]. Collision-Free [KS07c]. Color [Mat04]. Combined [KM00]. Common [Roz04a, Roz04b, Roz05]. Communication [AH03, BR06b, BR06c]. Communications [Kru05]. Compact [MP08]. Compactness [Rog06]. Compensated [JP09]. Complementary [Hol07]. Complete [CHLV08, HLRV03]. Completely [KP05]. Component [KR06, Nov04, Vis01]. Components [PK08, Tim04]. Composite [BD08, BD09]. Compound [AS05, Cek00, Cek03, FMS08, Kor01, Mik02]. Concentration [BH08, Rog02, Rog06]. Concept [KY05]. Condition [BM07a, BM07b, Gri03, Kru08, Roz04b]. Conditional [Di 04, Hes04, KS07b, Koz02, Sha06b, Shi09c, Ste03]. Conditional-Extremal [Shi09c]. Conditionally [She00]. Conditions [Fak09, Mor00, Sho05]. Cone [Nos02, Yak04]. Conference [Ano04f]. Configuration [GT07b]. Congratulations [Ano02c]. Conjecture [Mat04]. Conjugate [ZN04]. Connected [Kol04b]. Consequences [Roz00a]. Consistency [LS03]. Consistent [Che02a]. Constant [IS02, Joh04, Pre09, Rog00, She07]. Constants [Chi02b, Chi02a, Chi03, Deb06]. Constrained [Hol04]. Constraint [Tar09]. Constraints [BAH01, BHA02, Rok05]. Constructed [Kor00]. Constructing [BB06]. Construction [Ruz02].

Constructive [Rok08]. Contemporary [Ano04f]. Context [Pet03]. Context-Free [Pet03]. Contextual [Khr07]. Continuity [BE08, Now05, Ost02]. Continuous [BR06c, BK06a, GT07b, HS06, Kol09, Mit06, Roz02, Roz04b, VK09]. Continuous-Time [Kol09, Mit06]. Contoured [Sha06b]. Contractive [CCHH08]. Contractivity [Mat03]. Control [BKP04, LRT00]. Controlled [She00]. Convergence [Ale06, CHLV08, Che01, Chu00, DM08, GT03, HLRV03, Joh04, KTT08, Kon02, KP02, KB02, Kru08, Kru09, Kul03, Kuz06, Kva01, Lou02, Mak00, Mal01b, MP01, Pug09b, RR00, Sho05, Tim07, VK07, VM00, Ver00, VCH06, Mor00]. Convergencies [Kru04]. Convex [BK06b, DW06, DT03, Kul08, KP03, Pug09a, Ryc03, Sui02]. Convolution [Li02, Roo04]. Convolutions [Kar05, Rog06]. Copolymer [dHW07]. Correctional [DW06]. Correlated [MMSS08]. Correlation [BDK05, Ruk07]. Correlations [PT01]. Costs [KPT01, Rok08]. Couette [MM09]. Coupling [Lev02]. Couplings [Bor04a]. Covariance [Luo09]. Coverage [ASVV07]. Cox [Kor01]. Cramer [Boi03, PK03, Roz04b, Bak00, BM09a, BM09b, BM07a, BM07b]. Criterion [Rok08, SS00]. Critical [Afa04, FV01, FVW03, GK01, NV06b, NV06a, Vat03, VT05, VWF08, Wac08]. Crossing [Bor03a, Bor04b, Rag06a]. Crossings [Bor09a]. Current [BDG^{+}07]. CUSUM [HM06, HP09]. Cycles [Yak08].

D [GPL08]. D-Oriented [GPL08]. D. [Kha08b]. Dalang [Rok05]. Darmois [FG05]. Darmous [Fel01]. Data [Bak02, KK06, PCMGM09, VK09]. Decaying [Bor04b, BB05b]. Decision [Ryb01a]. Decisions [FP06]. Decomposition [GS07, Kar05]. Decompositions [OE09, Sen09, She03].

Deconvolution [BT08a, BT08b]. Decreasing [Rog03, Roz08, Vla05]. Definition [Rog00]. Denoising [KLP08]. Densities [BRS08, BRS09, BM01a, Nag07a, Zhu00a, Rok09]. Density [BT08a, BT08b, CPP02, GY07, Har01, Ibr07, Nov00a, RR05, Yak03]. Dependency [Gri03, Rot08]. Dependency-Neighborhoods [Rot08]. Dependent [AK04, DM08, Eic01, Fak09, Kam07, KR04, Lou02, Lou03, Nag01, Sha03, Sha05, VCH06]. Derivative [GE02]. Detection [BFS09, HM06, HP09, PT09, Shi09c, Shi09d, Sof02, Sof05, TV05, Tar09]. Determined [Ryc03]. Deterministic [Rad07]. Development [IM03]. Deviatations [Gap02]. Deviation [Arc03, Arc04, ABM08, BK01, BK02b, Erm04, Erm07, Kor02, Nag02b, Roz08, Sze06, Tim01]. Deviations [AS02, AILvZ09, BB05a, BR02, BDG^{+}07, BM01b, BB02, BB05b, BM07a, BM07b, BM09a, BM09b, CG03, CG04, Eic01, Eic05, Fat03, Fat04, FV01, FMS08, GDLH03, GL00, Koz02, Kru05, MPW06, Nag02a, Nag05, Nag07a, Roz04b, Roz09a, Tim03, Ver03, Wac08, ZN04, Zhu00b, Zhu05]. Difference [Ver00]. Different [IM03, Tim05]. Differential [Bak03, BNS08, BER05, BE06, BE08, CFS08, Che02b, IK00, IK01, Leb03b, Mal01b, MS07b, Ryb01b, Ver00]. Diffusion [AS09, BRS08, BRS09, Bog01, Bor03b, Har01, KS08, KS07c, KN02, LRT00, MT01, Nec00, PSV00, Pug09b]. Diffusion-Type [Bor03b]. Diffusions [BL01, KS07a]. Dimension [FVW03, KS07a, Tem00]. Dimensional [BK01, BK02b, Der05, HS06, Mak00, Mak04, MF08, Nos02, PSV00, Ver04, VD05]. Direct [PT01]. Dirichlet [MmVT03, Pug09b, RS01]. Discipline [Pec01]. Disciplines [CM04]. Discontinuous [BM01a]. Discounted [KR01b]. Discounting [BL01]. Discrete [Ali00, BR06b, Esq08, FG05, KL08, KPT01, LB04, Mis06a, Pus03, Rok06, Rok08, Rok09]. Discrete-Time [Rok06]. Disorder [Bur09]. Dispersive [Zie01]. Dissimilarity [BR06a]. Distinguishable [Tim01]. Distributed [Bor06, BD07, Chu00, PL07, Zhu00a]. Distribution [AL08, Bel02, Bel05a, BK05, Bor09a, Bor00, BB02, DKF08, Har01, KN08, Kho02, Kho06, Kly03, Kor02, Kul03, KP04, KPK06a, KPK06b, KN02, Kuz06, Lan07, Lob06b, Mac06, Mik02, Mis06b, Nov00b, PN03, PT09, PU02, PVK06, Rog06, Roo01, Roo02, Roz00a, Roz04a, Roz04b, Roz05, Roz08, RR05, Ruz02, She03, Shi01, Smo00, Tim03, Vin07, Yak03]. Distributional [Jed08, Jed09, SC00]. Distributions [Ali00, BY04, BRS08, BK01, BK02b, BB05b, Bor09b, Cek03, Fat04, Fat09, Gaa08, Gab01, GP05, HM08, Kar05, KS07b, Kul03, KP03, Kuz07, Mau04, Min01, NK09, Noq00, Nos02, Nov09, RS01, Rog00, Rya01, Ryc03, Sha06b, Ste03, VD05, Yak04, ZN04, Zie01, Zub00]. Divergence [Gus08]. Divisibility [BNPA03]. Divisible [BY04, Yak03, Yak04]. Dmitriev [Ano02e]. Dobrushin [Bor04a]. Domain [MMSS08, Roz04a, Roz05]. Dominated [Rog03]. Dominating [BT08a, BT08b, Ser00]. Donsker [CSY03, Web04]. Doob [Roz00c]. Downfalls [DSY00, Lob06a]. Drift [BFS09, Lob06b, Lob06a, SC00, Ver01]. Drifts [BR01]. Driven [BK09, CFS08]. Driving [Leb03b]. Duality [Dok04, PS02, RR01]. Dvoretzky [FP06]. Dynamical [Mas07, OK02]. Dynamics [BNS08, Kol09, Mal01a, Mas04b]. Dynkin [Ano05e].

E. [Ano05e]. Econophysics [Mas05]. Edgeworth [Rot08]. Editor [Nir03]. editors [Ano06b, Ano07b, Ano01c, Ano04b, Ano05c, Ano07a, Pin09]. Effective [NS05]. Efficiency [KN02, vRRvZ00]. Efficient [Erm04]. Elements [Fat04, HLRV03]. Elliptically [Sha06b]. Emeritus [Nir03]. Empiric [KN02]. Empirical [BDG^{+}07, LS04, Min01]. Emulsion [dHW07]. Energy [Vys06]. Ensemble [GT03, KTT08]. Ensembles [HS06]. Entanglement [Hol04]. Entanglement-Assisted [Hol04]. Entropy [CM04, GW00, Gap07, Gap09, Mas01, VG08]. Environment [Afa04, Afa07, AK04, Fak09, GK01, Vat03, VD04, Vat04, VD05, VD07, VD08, VD09]. Environments [Bor03b]. Epidemic [MS07a, Sta02]. Equality [Di 04, Lip06]. Equation [Bak02, Bak03, BNS08, BE06, DK06, Eng04, Mas00, Yen05]. Equations [Bel05b, BRS06, BER05, BE08, CFS08, Che02b, Der05, Dok04, Fat03, Kop01, Leb03b, Mak00, Mak04, Mal01b, Mas04a, MS07b, Pus03, Ryb01b, Ver00]. Equations.Part [IK00, IK01]. Equilibrium [GHKT07, Pes09]. Equivalent [Mat04, Rok09]. Erdos [BO07, Koz02]. Ergodic [GW00, Gap00, Jaj06, KN02, VK09, Web04]. Ergodicity [Mat03]. Error [Bog03, BS05, Yar09]. Esseen [Gad04, She07]. Estimate [Bog03, GY07, IK08b, Kul03, Mik02, Roz03a]. Estimates [AL08, BRS08, BF00, BP01, DDP07, Gap07, Gap09, KNP03, Mak03, Mal01b, Mit06, Orl09, Sen07, Shi01]. Estimating [Vak02]. Estimation [BK03, Bor09b, BHA02, CGLT04, GE02, Hol05, IK00, IK01, Ibr04, Ibr07, KR06, KLP08, KL03, KR01a, Pre09, RR05, Ser00, Sol00]. Estimator [GY07, KM00, Mag00, MPW06, Nov00a]. Estimators [ASVV07, BR06a, Erm07, vRRvZ00]. Euler [BD04]. Evolution [Mas05, VD07]. Exact [CS01, Fat09, FJT02, Gap09, IS02, Roz04a, Roz05]. Example [GL00, Ver04, Zap06]. Examples [Gap02]. Excessive [Dyn00].

Exchangeable [CS01, Lev01, MP01].

Existence

[Bak03, BR01, DW06, Leb03b, Mau04]. Exit [Dyn00, Har01, PT09, Rag06b]. Expansion [ASVV07, Chi02b, Chi02a, Chi03, Roo01, Roo03, Sen08, Smo00]. Expansions [Bog03, MR00, Rot08, Sen07, Sob08, Yar09]. Expectation [FJT02, Ser00]. Expectations [Di 04, Kru07]. Expected [NZ09]. Explicit [Sen07]. Exponent [Roo03]. Exponential [BK01, GY07, Mor04, Roz08, Sel05, Zai01, Zai02a, Zai02b]. Exponentiality [PT09]. Exponentially [Bor04b, BB05b]. Extended [Rok05]. Extending [Kru03]. Extension [Gne06, Gus08, JP09, Roz02, SC00]. Extensions [BNPA03, CSY03, Khr02, Leb02]. Extinction [Fak09, Kal02, Pet03, Shi01]. Extrema [Kor01]. Extremal [PKJ07, Roz00a, Shi09c]. Extreme [KN03]. Extremes [Ans06, Leb03a, Leb08, TC03].

Factorial [Kho02]. Factorial-Power

[Kho02]. Factorized [Mas00]. False [Tar09]. Families [Che02a]. Fast [Bak00]. Feedback [BKP04]. Feller [Mas00, Roz00b]. Feynman [DDP07]. Field [GZ05, Nos02, PT01]. Fields [Fat03, Leb03a, Leb08, MY02, Ost02, RS05, Sun08, VG08]. Filtered [CU04, SS00, Zhd09]. Filtering [MR00]. Filters [VK09]. Filtrations [OE09, VG08]. Final [Lan07]. Financial [Kor01, Mas04b, Mas05, MO01]. Finite [BO03, BH08, CCHH08, Fel01, Kam09, KO02, LY06, Mit06, Rok06, Rok08, Roz03a, Ryb01b, She03, She00, Shi09a, Sze06, Tim04, VD05, Zai01, Zai02a, Zai02b]. Finite-dimensional [VD05]. Finite-Means [BO03]. Finite-Velocity [KO02]. Finitely [EL00, Joh06]. First [Har01, Nov04, Nov09, PT09, Rag06b, Shi00]. First-Passage [Nov04]. Fisher [Luo09]. Fission [KS07a]. Five [PS09]. FKG [Joh06]. Flow [MM09]. Form [Nas00, Har01]. Formalism [Kho02]. Forms [GS05, Pug09b]. Formula [JP09, Khr07, Vis01]. Foundation [Shi00]. Four [Mat04]. Fourier [MR00]. Fractals [Tem00]. Fractional [GE02, Kol09, KP04, KPK06a, Mol00, Mol03, NN05]. Fractionally [PSV08]. Framework [BNS08]. Free [KS07c, Pet03, Vis01]. Function [ALL^{+}03, BB06, CCHH08, Gam01, Gap00, IK08a, KL08, KL03, KS07c, Roz04a, Roz05, KPK06b]. Functional [Bak02, DB04, KP02, Smo00]. Functionals [Bor00, Fat09, GSY07, Mis06b, Roz02, Sch03, Sha03, SY04]. Functions [AL08, AB06, DR02, Dyn00, Esq08, GT07b, KL08, KP03, LS04, MT01, Neu00, Rog02, Rog03, Rog06, RS00, Roz00c, Rya01].

G [Ano07f]. Galton [Afa07, Afa08, FV01, NV06b, NV06a, VD04, VD05, VWF08, Wac08]. Game [BK09, GHKT07, Min00, Min01]. Game-Theoretic [Min01, Min00]. Games [Dom02, Mas04a, Pes09]. Gas [Der05, Mas07]. Gas-Dynamical [Mas07]. Gaussian [Ans06, AILvZ09, DT03, GT03, KTT08, Kob05, KP04, KPK06a, Kva01, MY02, NN05, PT01, PK08, Roz09b, VK07]. General [Bel05a, FK01, KL08, TV05, Yak08]. Generalization [BR01, Mla06]. Generalizations [Bob01]. Generalized [BK03, Bur09, Deb06, DB04, Kol09, Nov00a, Roo02]. Generalizing [MS07a, Sta02]. Generating [Esq08]. Geometric [CS04, Mac06]. Gibbs [Now05]. Given [Mau04]. Global [BRS06, Tar09]. Golden [BO07]. Goncharov [IM03]. Good [Leb02]. Gordon [DK06]. Governed [Bor07]. Gradient [Ver01]. Gradient-Type [Ver01]. Grammars [Pet03]. Gravitating [Vys06]. Group [Amo05, Mol03, She03, VT07]. Groups [Bor04a, Fel01, FG05, Gab01, Ver04]. Growth [Kru07]. GUE [Tim07].

H.F.D. [KPK06b]. Haan [Mla06].

Hammerstein [Fat03]. Harmonic [RS00]. Hausdorff [KS07a]. Haussmann [JP09]. Having [Roz04b, Wac08]. Heavy [Leb05, Bor03a]. Hedging [Nec00, Rad07]. Height [Lob06b]. Heritage [Kol04a]. Hermite [Kon02, MR00, Osm09]. Hidden [Bor07]. High [Afa08, Fat03, PS04, Roz03a]. High-Order [Fat03]. Hilbert [Bog03, Khr07, MP01, MS07b, Roz09b]. History [Shi00]. Hitting [Kol09]. Hölder [Di 04, KNP03, RS05]. Holevo [Shi09a]. Homeostasis [MMZ07]. Homogeneity [Rom01]. Homogeneous [Mit06, Smo00]. Horizon [Kam09, Nos02, Rok09]. Horizontal [CSY03]. Horizontal-Vertical [CSY03]. Hydrodynamic [LMS01]. Hyperbolic [Fat03, KS08]. Hypergeometric [Tim03]. Hypotheses [BD08, BD09, DS03, Ryb01a]. Hypothesis [AH03, BAH01, Ing01].

I. [Kha08c]. Ibragimov

[Ano03a, The03, Kha08c]. Idempotent [DD00]. Identical [Man09]. Identically [BD07, Chu00, PL07, Zhu00a]. II [Arc04, BR06c, BM01b, BB05b, BM07b, BM09b, BD09, BT08b, Chi02a, DD00, GSY07, IK00, Jed09, KS06b, KPK06a, Min01, RR00, Roz05, Smo00, VD05, Web07, Zai02a, Zhu05]. III [Chi03, IK01, Zai02b]. Ildar [Ano03a, The03]. Ill [CGLT04]. Ill-Posed [CGLT04]. Immigration [Li02]. Importance [Erm07]. Including [BM01b]. Incomplete [Mik03]. Increasing [Neu00, VD06]. Increment [KL05]. Increments [Amo05, Amo06, Fro03, Fro04, Fro05, Gra00]. Independence [Rom01]. Independent [BM07a, BM07b, BM09a, BM09b, BD07, Chu00, Fat04, Fro03, Fro04, Fro05, Gra00, KN03, Kor02, Kru07, Leb00a, Leb05, Lip05, Mat09, MF08, Nag02b, Nag02a, OE09, PL07, Pet02, Rag06a, Roz03a, Roz04a, Roz04b, Roz05, Roz08, She00, ZN04, Zhu00a]. Index [Ano02b, KLP08, Leb08]. Indices [IK08a]. Individuals [Afa04]. Induced [VT07]. Inequalities [AR06, Bob01, CS01, Gad04, Kol04b, Kru05, Lou03, Nag01, NV06a, Nag07b, Noq00, Ryc03]. Inequality [Bob03, Gam01, IS02, Kru06, KP03, Lev01, Pre09, PVK02, PVK06, Roz00c, She07]. Inference [Erm04, RS01]. Infinite [BNPA03, BS05, Bor06, HS06, Kru07, Ver04, VT07, Rok09]. Infinite-Dimensional [HS06, Ver04]. Infinitely [BY04, Yak03, Yak04]. Infinity [Yak03]. Information [BAH01, BHA02, Joh06, Luo09]. Information-Theoretic [Joh06]. Inhomogeneous [KM00]. Initial [Bak02, Har01, VK09]. Integers [Tsi00]. Integrability [BK06a, Gap00, Shu06]. Integrable [BR01]. Integral [BB06, CS04, GSY07, Mas04a, Rag06a, SY04, Tim03]. Integrals [Che01, KS02, Nas00, Nas07]. Integrals.Part [Zan04]. Integrated [PSV08]. Integrating [BD04]. Integro [BM01b]. Integro-Local [BM01b]. Intensity [KM00]. Interacting [BDG^{+}07, LMS01, Sta02, Yar07]. Interaction [Kal02, Kal03, MS07a, Nad07]. Interactions [Lan07]. Interference [Khr07]. Intermediate [Vat04]. International [Ano00h, Ano01h, Ano02j, Ano04f]. Interpolation [Klo07b]. Interpretations [Khr02]. Intervals [Kly03]. Intrinsically [Ma09]. Invariance [BD07, CSY03, GZ09, Lin00, Sha03, Zai01, Zai02a, Zai02b]. Invariant [BR01, LS02, Mal01a, Mal01b, Ver04]. Inverse [CGLT04, vRRvZ00]. Inversion [Vis01]. Investments [Bog01]. Irreversible [Bog01]. Isomorphisms [Zub00].

Isoperimetric [Bob03]. Itô [GV07, LY06]. Iterated [ALL^{+}03, Ćwi09, EL00, Kov02, Mat00, Min00, MP08, PL07, Pet02, Pin06, Roz03b, Sha05]. Ito [Dok04, Nas00]. Ivanovskaya [Ano02f].

J. [Roz00c]. James [ASVV07]. Joint [Gaa08, PN03, Ruz02, Vin07]. Jump [Bor07, Nov04]. Jumps [Bor03a, Bor06, Mor04, PK03, Rad07].

Kac [DDP07]. Kantorovich [Lip06].

Kantorovitch [RR01]. Kernel

[MPW06, Nov00a]. Kerstan [Roo03]. Khasminskii [BR01]. Khinchin [Bob01]. Khinchin-Type [Bob01]. Kinds [Kal02]. Klein [DK06]. Known [Tim04]. Kolmogorov [Ano04f, Kol04a, Mas00, PS09, Roz03b, Ver04]. Kolmogorovian [Khr07]. Kornya [Roo04]. Krawtchouk [Roo01, Roo02].

Lagrange [Vis01]. Laplace [BB05a, Fat04]. Laplace-Type [Fat04]. Large [AS02, Arc03, Arc04, ABM08, BB05a, BR02, BDG^{+}07, BM01b, BK01, BK02b, BB02, BB05b, BM07a, BM07b, BM09a, BM09b, Bor09b, DM08, Eic01, Eic05, Erm07, Fat03, FK01, Fro03, GDLH03, GL00, IK08a, KS08, Kol04b, Kor02, Koz02, Kru03, Kru05, Man09, Min00, MPW06, Nag02b, Nag02a, Nag05, Nag07a, Pet09, PK03, Roz04a, Roz05, Roz09a, Sha06b, Shi02, Sze06, Tim01, Tim03, VWF08, Ver03, Wac08, ZN04, Zhu00b, Zhu05]. Large-Deviation [BK01, BK02b, Kor02, Tim01]. Large-Time [KS08]. Lattice [Kar05]. Lattices [GPL08, Mat00]. Law [Afa07, AS02, AS05, Che02b, CHL02, Chu00, Ćwi09, DM08, EL00, FK01, GT03, Gri00, IK08a, KTT08, Kol04b, Kov02, Kru03, KP05, Mat00, Min00, MP08, PL07, Pet02, Pet09, Pin06, Roz03b, Roz04a, Roz05, Sha05, Sha06b, KPK06b]. Laws [CM03, Dyn00, Hes04, KP05, Mas05, RS01, Shi02, TC03]. Leads [Vin07]. Least [NZ09]. Lebesgue [Nas00]. Ledoux [Bob03]. Lee [Nec00]. Lemma [KN08]. Length [Lob06b]. Letter [Ano01c, Ano04b, Ano05c, Ano07a, Pin09]. Letters [Ano06b, Ano07b]. Level [Afa08, PS04]. Levy [Mor04, BK09, Mol01, Sel05, SC00]. Life [Ano00a, Ano00b, Ano01e, Ano03c, Ano05d, Ano07d, Ano07e, CB06, Che08a, SM08, Shi09b, Ano01d, Ano02d, Ano04d]. Lifetimes [PSV00]. Likelihood [BM01a]. Limit [Bak02, BG04, BM01b, CPP02, CSY03, CF04, FVY00, Fro04, Fro05, Gaa08, GS05, GT07a, Gri03, IL00, Joh06, KL08, Kam07, Kly03, Kop01, LMS01, Leb00b, Leb00a, Mac06, Mak04, Mik03, Mik09, Min01, MS07a, NV06b, PS08, PS04, RS05, Rag06a, Rog02, Roz03a, Sch03, Sen07, Sen09, Ser06, Sha06a, She00, Shu01, Shu06, Sob08, Su02, Tim03, Tim05, VD04, Vat04, VT05, VD08, Wac08, Yak08, Yar09, ZN04]. Limiting [KS07a, LY06, PN03, Min01]. Limits [DKF08]. Linear [Di 04, LB04, Mol03, MF08, Roz02, Ryb01b]. Lipschitz [ALL^{+}03]. Living [FVW03]. Local [BM01b, BD07, CPP02, FVY00, Mal01a, NV06b, She00, Vla05]. Localization [Bob03, Kal04, dHW07]. Locally [BR01, Pug09a]. Location [BR06a, KR06]. Location-scale [BR06a]. Logarithm [CHL02, Ćwi09, EL00, Kov02, Mat00, Min00, MP08, PL07, Pet02, Pin06, Roz03b, Sha05]. Logarithmic [NN05]. Long [Bog01, FVW03, PSV08]. Long-Living [FVW03]. Long-Term [Bog01]. Lower [DKF08, GZ05, Klo07a, LS04, Mat09, Nag02b, Nag02a]. LQG [BKP04]. LQG-Problem [BKP04]. Lundberg [Boi03]. Lyapunov [AB02, Ben05, Chi02b, Chi02a, Chi03, Kru08]. Lyapunov-Type [AB02, Ben05].

M [ABV^{+}07, Kha08b, Kha08a, Klo07b]. MA [HT07]. MacLaurin [BD04]. Macrodimension [Mal01a]. Majorization [KN08]. Many [BD08, BD09]. Map [Mat03]. Mappings [BK02a, DB04, Lip05, Tim04]. Marginal [BNPA03]. Marginals [Mau04]. Margins [HT07]. Market [MO01, Rok09]. Markov [Ber09, BO07, Bor07, BK01, BK02b, Bor03a, Bor04b, CCHH08, Eng04, KV05, Klo07a, KL05, Lin00, Mit06, PT09, Ruk07, She00, Shi01, Shu01, Sze06, VT07]. Markovian [Amo05, Amo06, Bak00, Bor00]. Martingale [Gne06, GT07a, Gra00, Rok06, Sel05, VM00]. Martingales [CFS08, Kal04, Nag07b, Nov04, Pre09, RR00, Sho05]. Maslov [DD00]. Mass [BK06a]. Masses [Pus03]. Matches [Mik03]. Matching [Mik02, Mik09]. Mathematical [BK05]. Mathematics [Ano04f, Kor01, Mas04b]. Matrices [GT07a, Ruk07]. Matrix [BNPA08, MP08]. Matrix-Normalized [MP08]. Matsumoto [KL08]. Max [TC03, HT07]. Max-AR [HT07]. Max-Semistable [TC03]. Maxima [CDT01, Chu00, Kor02, Leb00a, Leb05, PN03]. Maximal [Afa04, AR06, CS01, Gap00, Leb06, Lev01]. Maximization [CM04]. Maximum [BDK05, GPS01, HT07, Mol00, Mol01, Uru05, VWF08]. Maximums [Leb00b]. McKendrick [MS07a, Sta02]. Mean [GZ05, IS02, Kor02, Nec00, Ryc03, Zhd09]. Mean-Field [GZ05]. Mean-Variance [Nec00, Ryc03]. Means [BO03, VCH06, Web06, Web07]. Measurable [DR02]. Measure [GK05, Mal01b, Now05, Orl09, Roz09b, Su02, Ver04, Zhd09]. Measure-Valued [Su02]. Measures [Ale06, BO07, BR01, BK02a, BRS06, BK06a, BK06b, Che02a, CU04, GL09, Gra00, Kar05, Kul08, Leb02, Leb03b, LS02, Pug09a, Roo04, SS00, Sel05, Sha06b, SP06, Su02, VT07, Vla05, Rok09].

Medium [KS07c, PSV00]. Mejzler [Mla06]. Memoriam [Ano03b, Ano04a, Ht09, Nir03, Ano05b]. Memory [Bak03, PSV08]. Merging [CM03]. Method [IM03, Roo03]. Methods [Mas03, Pas05, Shi09d]. Metric [Gap07, Gap09, GL09, Tem00]. Meyer [Ano04a]. Micromodel [MM09]. Minimal [Gri03]. Minimax [BD08, BD09, BFS09, GE02, Sol00]. Minimax-Bayesian [Sol00]. Minimization [CM04]. Minimizing [Rad07]. Minimum [BR06a]. Minorants [Sui02]. Mises [Eic05, Sha03]. Mixed [Mor04]. Mixed-Exponential [Mor04]. Mixing [Jed08, Jed09, KV05, Klo07a, Rot08, Sun08, Ver00, Ver01]. Mixingale [Shi02]. Mixture [Kuz07]. Mixtures [RS01]. Mode [MPW06, Nov00b]. Model [Bog01, Boi03, BP01, HM06, HP09, Khr07, KM00, MM06, MS07a, Nec00, Rad07, Sta02]. Models [Ano01h, Ano02j, BS05, Bor07, DDP07, GZ05, HT07, IM03, Kon02, Sel05, Ser06, Shi09d, VM00, Rok09, Ano00h]. Moderate [CG03, CG04, Erm04, Erm07, Fat04]. Modification [DR02]. Moduli [AL08, Ost02]. Moiseevich [Ht09]. Moment [Cek00, Lou03, Mak03, MO01, Nag01, Nag07b, Shi01, Uru05]. Moments [GV07, Kon02, Roz03a, Zai01, Zai02a, Zai02b]. Monge [Lip06, RR01]. Monge-Kantorovitch [RR01]. Monin [Ano02a]. Monograph [PS09]. Monotone [Di 04, Roz02]. Morton [Rok05]. Mosco [Pug09b]. Motion [BFS09, CS04, DSY00, GPS01, GSY07, HM06, HP09, KO02, Lob06b, Lob06a, Mol00, Mol03, SC00, SY04, Shi09c, Sui02, Uru05]. Motions [Bak00, DW06, OE09]. Moving [ABM08, PS08]. Multi [KLP08]. Multi-Index [KLP08]. Multidimensional [Bak02, BB05a, BM01a, Gam01, GZ09, Kul08, Nag07a, PCMGM09, Rag06b, Sen09, Tim03, Yen05, Zai01, Zai02a, Zai02b]. Multifractal [GT07b]. Multimodal [GP05, Vak02]. Multinomial [Kho02, Roo02]. Multiple [HM06, IK08a, Ryb01a]. Multiplicative [LS02]. Multiplicativity [AH03]. Multistability [GDLH03]. Multitype [FV01]. Multivariate [Cek03, Ibr04, Ibr07, Osm09, PT01, Roo03]. Mutual [SS00].

N [Kol04a, PS09, tbYGS05]. Nash [GHKT07, Pes09]. Natascha [Nir03]. Necessary [Mor00, Sho05]. Necessity [Kru08]. Negative [BK09, Kor02, Mor04, PVK06]. Negatively [CHLV08, CHL02]. Neighborhood [Har01]. Neighborhoods [Rot08]. Networks [OK02]. Newman [Sha06a]. News [Ano00a, Ano00b, Ano01d, Ano01e, Ano02d, Ano03c, Ano04d, Ano05d, Ano07d, Ano07e, CB06, Che08a, SM08, Shi09b]. Nikolai [Ano02e]. No [Rok08]. No-Arbitrage [Rok08]. Noise [BB06, Leb03a, Leb08]. Non [Bor00]. Non-Markovian [Bor00]. Nonclassical [PL07, Yar09]. Noncorrelated [BG04]. Nonformation [BN00]. Nongradient [LMS01]. Nonhomogeneous [Vin07]. Nonidentically [Bor06]. Nonlinear [Bak03, Bel05b, BK06b, Fat03, KLP08, Mas04b, Mas05, MR00, Rag06a, Yar07]. Nonlinearity [Dok04]. Nonnegative [AR06, Leb06, Pet09]. Nonparametric [BK03, LL03, RS01]. Nonpredictable [Shi09c]. Nonrandom [BB06]. Nonrandomized [FP06]. Nonsmooth [Dok04]. Nonstationary [Ans06, PSV08]. Nonuniform [Bog03]. Nonzero [Pet03]. Norm [Fat03]. Normal [Bor01, Gri00, HM08, KS06a, KS06b, Kru04, Kru08, PN03, Sun08, Zhu00a]. Normalization [Roz09a]. Normalized [MP08, Nag05, Nov05]. Normed [Kon02, Kov02]. Note [Bor04a, Bor09a, Bor03b, Chr04, DW06, Lip06, PS02]. Notion [Gus08]. NPPs [Mas07]. Number [ALL^{+}03, AB06, Bor09a, Imh03, Kly03, Kop01, Lan07, Mik02, Mik03, Mik09, NZ09, Sze06, Tim04, Vin07, Yak08]. Numbers [Afa04, DM08, FK01, GW00, IK08a, Kol04b, Kru03, Lev01, Mas01, Min00, Pet09, Roz04a, Sha06b, Shi02, Roz05]. Numerical [Rom01].

Obituary [Ano02e]. Observable [Shi09c]. Observations [BG04, Sha03]. Observed [CDT01]. Occupation [Fat09, Mas01]. Ocone [JP09]. One [ALL^{+}03, BK01, BK02b, Gne06, Hes04, IK08b, Mak00, Mak04, Mat04, Nos02, PSV00, Rag06b, Sen08, Ser06, Ste03]. One-Dimensional [BK01, BK02b, Mak00, Mak04, Nos02, PSV00]. One-Sided [Ste03]. Online [KL03]. Open [BK02a]. Operator [Kov02, MY02]. Operator-Self-Similar [MY02]. Operators [Amo05, Di 04, Jaj06, Shu01]. Optimal [AS09, BL01, BK06a, Dom02, HM06, KR04, LRT00, MP04, MO01, NS05, Nov09, Pes09, Ruz02, Sal03, Shi09d, Sof02, Uru05, Sof05]. Optimality [BKP04, BR06b, BR06c, BFS09, BT08a, BT08b, Ryb01a, Tar09]. Optimization [Bog01]. Option [Kam09]. Optional [GS07]. Options [CS04, Chr04]. Order [BO03, BY04, Ber09, BP03, BFS09, Fat03, PCMGM09, Roz03a]. Order-Unit [Ber09]. Ordering [Zie01]. Orders [Ryc03]. Oriented [GPL08]. Ornstein [BN01, BR02, Fat09, Nov04]. Orthogonal [Kho02, KTT08, OE09]. Orthogonality [BB06]. Overshooting [BF00].

P. [SC00]. Pairs [Mik09]. Pairwise [Kru07, Lan07]. Parabolic [Bel05b, BRS06, Dok04, KNP03]. Parameter [DS03, Hol05, KR06, KR01a]. Parameters [BR06a, Pus03]. Parity [PS02]. Part [Arc03]. Partial [ABM08, Eic01, IK00, IK01, Koz02]. Particle [BDG^{+}07, DDP07, LMS01, PSV00]. Particles [FVW03, Gaa08, Kal02, Kal03, KS07c, Lan07, Man09, MS07a, Sta02, Tim01, Tim05, Vys06, Yar07]. Partitions [FVY00, Tsi00]. Parts [KP04, KPK06a]. Passage [Nov04, Nov09]. Path [MY02]. Pathways [MMZ07]. Pathwise [Che02b, Nas00]. Pattern [Ruk07]. Patterns [Mik02, Mik03]. Paul [Ano04a]. Payment [Vin07]. Pearson [CF04, Kru01]. Periodically [MMSS08]. Periodogram [PCMGM09]. Permutation [Yak08]. Permutations [IM03, Rya01]. Perturbations [Amo05, Amo06]. Phase [MM06, Mas04a]. Phenomena [Ale04, Bor06, Sak05]. phi [AR06]. phi-Inequalities [AR06]. Physics [Mas04a]. Pickands [Deb06]. Piterbarg [Rom01]. Planar [KO02]. Poetry [KR00]. Poincaré [Joh04]. Point [Bor09b, HM06, Har01, Sof02, Sof05]. Pointwise [Jaj06]. Poisson [AS02, AS05, Bor04a, Bur09, Cek00, Gaa08, JP09, Kar05, KL05, Kru04, KM00, Leb00b, LS02, Mik02, Mik03, Mik09, Roo01, Roo03, Roo04, Sho05]. Poisson-Type [Mik09, Mik03]. Polling [Ser06]. Polyhedrons [DW06]. Polynomial [NZ09, Sal03, Sen08, Ver00, Ver01, Zap06]. Polynomials [Bob01, Imh03, Kho02, Osm09, PK08, RR05]. Populations [VD07]. Portfolio [Rok05]. Posed [CGLT04]. Position [Har01]. Positive [BRS09, Tsi00]. Possible [GPS01]. Possibly [BS05]. Posterior [NK09]. Posteriori} [Sof05]. Power [BNS06, Kho02, Pin06, Wac08]. Precise [Fat04, Gap07]. Predictable [SS00]. Prediction [LB04]. Predictor [BD04]. Premium [Boi03]. Prescribed [Imh03]. Presence [KR06, Leb03a, Leb05, Leb08]. Presman [Roo04]. Pressureless [Der05]. Prestationary [BK01, BK02b]. Pretest [Mag00]. Price [Mas04b]. Pricing [Che08b, Chr04]. Principle [Arc03, Arc04, ABM08, BD07, CSY03, GZ09, Lin00, Sha03, Zai01, Zai02a, Zai02b]. Prize [Bor08]. Probabilistic [Bel05b, BP01, Mat04, Pas05]. Probabilistic-Statistical [Pas05]. Probabilities [BRS08, BRS09, BK01, BK02b, BB02, BB05b, Erm04, Erm07, FMS08, KY05, Kor02, Mor04, Nag02b, Nag02a, RS01, Roz08, Wac08]. Probability [ASVV07, Ans06, BR06b, BR06c, BK03, BK05, BK02a, BRS06, Bor03a, Bor04b, Che02a, DD00, DSY00, Esq08, GK01, GL09, Ibr07, IK08b, Kal02, Kal03, Khr02, Khr07, KR00, Kol04a, Kul03, Lob06a, Mas00, Mas03, Mau04, Nag01, NV06a, Nag07b, Nov00b, Pet03, Rya01, SS00, Shi00, Tar09, Tim07, Ver04, Zub00]. Problem [Bel02, BP01, BFS09, Bur09, DS03, Dar05, DK06, Gap02, GE02, GSY07, Hol07, Klo07b, Kob05, KR01b, MmVT03, NS05, PU02, Rok06, Ryb01a, SY04, BKP04]. Problems [Ano00h, Ano01h, Ano02j, AS09, BAH01, BHA02, CGLT04, Fat03, IK00, IK01, Kor01, Lip06, Mol03, Nag07a, RS00, Shi09c, Shi09d, Tar09, Uru05]. Procedures [LN00, LS03, Rom01, Ruk01]. Process [Afa04, Afa08, Amo05, BK09, BR02, Boi03, Bur09, DB04, GK01, GL00, Har01, Kal02, Kal03, Kob05, KM00, Lan07, Mas05, MS07a, Mis06a, NV06b, NV06a, PS04, Shi01, Sta02, Vat04, Vin07, Wac08, Yar07]. Processes [Afa07, AK04, Amo06, Ans06, Arc03, Arc04, ABM08, AS09, AILvZ09, BN01, BR06b, BR06c, BD04, BK03, Ber09, BRS08, BRS09, Bor07, Bor00, Bor03b, CS04, Che01, DT03, Eic01, Fak09, Fat09, FV01, Fro05, FMS08, Gap07, GV07, Gra00, GDLH03, JP09, Jed08, Jed09, KV05, Klo07a, Kor01, KL05, KN02, Leb06, Li02, LS04, MMSS08, MT01, Mas07, Mol01, Mor04, NN05, Nov04, PKJ07, PSV08, PT09, SP06, Su02, Vat03, VD04, VD05, VD07, VWF08, VD08, VD09]. Product [PT01]. Professor [Ano03b]. Progenies [VWF08]. Projections [Kho06]. Prokhorov [Ano01a, Ano06c, Bob01, CSY03, GL09, Kru03, Kru06]. Proof [Bob03]. Propagation [DDP07]. Properties [BR06a, Deb06, Jed08, Jed09, KN08, KS07b, Kor00, Kor01, KM00, MY02, MP04, Mis06a, Nag07a, PCMGM09, Shi08, SC00, ZN04, Kho06]. Property [BR06b, BR06c, Lip05, Mac06, Roz00a, Uru05]. Proportions [FV01]. Publication [PS09]. Publications [Shi00]. Purely [Min00, Min01]. Put [PS02].

Quadrant [Kal03]. Quadratic [GS05]. Quantiles [KS07b, Kor00, Ryc03]. Quantization [GL09]. Quantized [Mas01]. Quantum [AH03, Amo06, Hol04, Hol05, HS06, KS07c, Luo09, Mas03, Shi08]. Quasi [Gap07, LS02]. Quasi-Invariant [LS02]. Quasi-Stationary [Gap07]. Queue [Pec01]. Queues [GDLH03]. Queuing [Leb00b]. Quickest [BFS09, HP09, PT09, Shi09c, Shi09d, TV05].

Random [ALL^{+}03, Afa04, Afa07, AK04, Ale04, Bak02, BB05a, BL01, Bor09a, BF00, BM01b, BB02, Bor03b, BB05b, Bor06, BM07a, BM07b, BM09a, BM09b, BD07, CHLV08, CSY03, CS01, Chu00, DB04, Esq08, Fak09, Fat03, Fat04, FVY00, Fro03, Fro04, Gaa08, GK01, GT07a, GPL08, HLRV03, IS02, Imh03, IM03, Jan03, Kal03, Kam07, KN03, KY05, KO02, Kol09, Kop01, Kor00, Kor02, Kov02, KB02, Kru07, KP04, KPK06a, Kva01, Leb00a, Leb02, Lev01, Lou02, Lou03, MP01, MY02, Mat03, Mat09, MmVT03, Min00, Min01, Mis06b, MP08, Nad07, Nag01, Nag02b, Nag02a, NZ09, Neu00, Nos02, NS05, Ost02, PKJ07, PL07, Pet02, Pet03, Pet09, PK03, PK08, Pus03, RS05, Rag06b, RS01, Roz03a, Roz04a]. Random [Roz04b, Roz05, Roz08, Roz09a, Ruz02, Rya01, Sak05, Sch03, Sha05, She00, Smo00, SP06, Su02, Sui02, Sun08, TC03, Tem00, Tim04, Tsi00, Vat03, VD04, Vat04, VD05, VT05, VD07, VD08, VD09, Vis01, Vla05, Yak08, ZN04, Zap06, Zhu00a]. Randomized [Dom02]. Randomly [DKF08]. Randomness [Ruk01, Ruk07]. Range [BH08]. Rank [PT01]. Rapidly [VD06]. Rarely [CDT01]. Rate [GT03, GZ09, HLRV03, KTT08, KV05, Klo07a, Kon02, Kul03, Mal01b, PS08, RR00, Tim07, Ver00, VCH06]. Rates [DM08, Lou02]. Ratio [Afa04, BO07, BM01a, FJT02, Rog02, Shu01, Shu06]. Reaction [MMZ07]. Real [Esq08, Imh03, Lev01, NZ09]. Real-Valued [Esq08]. Rearrangements [DT03]. Reconstructing [Bel02, PU02]. Reconstruction [Bel05a]. Records [Ali00, LB04]. Recovery [Dar05, GK05]. Recurrence [AK04]. Recursive [KR01a]. Reduced [Vat03, VD08, VD09]. Refined [Eic05]. Reflecting [DW06]. Reflection [KO02]. Regina [Ano02f]. Regression [Ibr04, KL03, LL03]. Regular [BL01, BB05b]. Regularity [BRS06, Web04]. Regularly [Leb03a, Leb08, PCMGM09, BB02]. Related [BNPA08, Li02, Shi08]. Relation [Kon02, Mas01]. Relationship [GDLH03]. Remainder [Roz03a, Sen07]. Remark [Roz03b, Sze06]. Remarks [Klo07b]. Renewal [BO03, Eng04, FMS08, Yen05]. Renyi [Koz02]. Reparameterization [Zie01]. Replication [KPT01]. Representation [BK02a, BM01a, KS02, Khr07, Nas00, Zan04]. Representations [GSY07, SY04, VT07]. Reservoirs [LMS01]. Respect [Kho02, KY05, LS02]. Restrictions [AB06, Cek00, Tim01, Tim05]. Result [Zai01, Zai02a, Zai02b]. Results [Bob01, FJT02, MP01]. Retransmission [GDLH03]. Reviews [Ano00d, Ano00e, Ano00f, Ano00g, Ano01f, Ano01g, Ano02g, Ano02h, Ano02i, Ano03d, Ano03e, Ano03f, Ano04g, Ano06d, Ano07g, Ano07h]. Rhythm [KR00]. Risk [BK03, Che08b, GS07, KR01a, Kul08, Orl09, Vin07]. Robustness [BR06a]. Role [KN03]. Roots [NZ09]. Rosenthal [IS02]. Rotation [Web06, Web07]. Rotations [Jan03]. Rough [DK01]. Row [CHLV08, Ryb01b]. Row-Finite [Ryb01b]. Row-Wise [CHLV08]. Ruin [BK03, IK08b, Kob05, Mor04, Vin07]. Rule [HP09, Ryb01a]. Rules [HM06]. Russian [Kam09, KR00, Shi00].

Sakhanenko [Zai01, Zai02a, Zai02b].

Sample

[Bor09b, CDT01, Ibr07, Kor00, MY02]. Samples [BM01a, Kor00]. Sampling [Erm07]. Sazonov [Ano03b]. scale [BR06a]. Scaled [VG08]. Scheme [Gaa08, Kru05, Sha06b, Tim01]. Schemes [MS07b, Sal03]. Sciences [Shi00]. Scientific [Ano00a, Ano00b, Ano01d, Ano01e, Ano02d, Ano03c, Ano04d, Ano05d, Ano07d, Ano07e, CB06, Che08a, SM08, Shi09b]. SDE [Ver03]. SDEs [Ver01]. Second [BO03, BY04, BP03, BFS09, PCMGM09]. Second-Order [BO03, BY04, BP03]. Sector [IK08a]. Selection [Rok06]. Self [MY02, Nag05, Nov05, Shu01]. Self-Adjoint [Shu01]. Self-normalized [Nag05, Nov05]. Semicircle [KTT08]. Semicircular [GT03]. Semicontinuous [Neu00]. Semigroup [Ver03]. Semigroups [Jan03, Li02]. Semimartingale [DW06, MT01, MO01]. Semimartingales [Mor00, Zan04]. Seminar [Ano00h, Ano01h, Ano02j]. Semistable [CM03, TC03]. Separating [CU04]. Sequence [KY05, Noq00, PN03, Pet02]. Sequences [Gap02, Gap09, GS05, Gri00, Gri03, KR01b, KR04, Leb00b, Mik09, Min00, Min01, Nov09, Ruk07, Sal03, Shi02, TC03, VD06, Vla05]. Sequential [BD08, BD09, Ste03]. Sergeevich [Ano02a]. Series [Kva01, Pin06, VK07, Vak02]. Service [Pec01]. Set [Gap07, KS07a]. Sets [KP03, Sch03, Tim04, Zub00]. Setting [Bur09]. Sevastyanov [Ano04e, Kha09]. Seventy [PS09]. Seventy-Five [PS09]. Several [Sal03]. Severely [CGLT04]. Sharp [BB05a, BR02, BT08a, BT08b, CGLT04, DDP07, Ryc03]. Sharpening [She07]. Sheets [LS04]. Shepp [BK09]. Shift [Hol05]. Shiryaev [BK09, tbYGS05]. Short [Bog03, Kru05]. Shot [Leb03a, Leb08]. Shot-Noise [Leb03a, Leb08]. Sided [HP09, Ste03]. Sign [BS05]. Signed [Kar05, Roo04]. Signs [Rya01]. Similar [MY02]. Simple [Ing01, Mis06b, SP06]. Simplest [MmVT03]. Simulation [Erm07]. Simultaneous [Ans06]. Sinai [Ano07f]. Since [PS09]. Singularity [SS00]. Size [Tim01, Tim05]. Sizes [Kor00]. Skew [Li02]. Skip [Vis01]. Skip-Free [Vis01]. Skitovich [Fel01, FG05]. Skorokhod [Ano01b, BK02a]. Slow [Bak00]. Small [AILvZ09, FMS08, Har01, NN05]. Smooth [AILvZ09, BD04, KL03]. Smoothness [AL08, Shi02]. Solution [Bak03, Bel05b, BE06, NS05]. Solutions [Bak02, BRS06, BER05, BE08, DK01, Kop01, KNP03, Leb03b, Mak00]. Some [Bel05a, Bob01, Bob03, BAH01, BHA02, Che01, CSY03, Deb06, GW00, Gap02, Gap09, GZ05, Klo07b, Kol04b, Kor01, Lin00, Lou02, MP01, Mis06b, Osm09, Pas05, RR00, Roz00a, Sen09, SC00, Uru05, Vla05, Web04, NN05]. Space [Bog03, GT07b, HLRV03, KL08, KS08, Khr07, Ma09, Mas00, MS07b, RS05, Roz09b, SS00, Sha05, Shi02, Yen05]. Spaced [PCMGM09]. Spaces [Ber09, CU04, Kva01, MP01, Pug09a, RS00, Tem00, VK07]. Sparse [KLP08]. Spatial [BD04]. SPDEs [KNP03]. Special [Bor07]. Spectra [GT07a, Tim07, Vak02]. Spectral [MMSS08]. Spectrally [BK09]. Spectrum [GZ05]. Sphere [Kho06]. Spheres [PT01]. Square [FJT02, Gap00]. Squares [FJT02].

SRPT [Pec01]. Stability [Ano00h, Ano01h, Ano02j, Mit06, OK02, Yar07]. Stable [BH08, Gab01, Jed08, Jed09, KP05, MP01, Mol01, Mor00, Nag07a, Roz04a, Roz05, Sha06b, Smo00, VD06, Zan04, Zie01]. Stable-Type [BH08]. Standard [DSY00]. Star [Ryc03]. State [AK04, Bor07, Fak09, Hol05, PSV00]. State-Dependent [AK04, Fak09]. States [Sze06]. Stationary [Amo05, Amo06, AILvZ09, Bak03, BRS08, CHL02, Gap02, Gap07, Gap09, GS05, Gri00, Gri03, GL00, Kob05, Ma09, PSV00, PN03, TC03, Tsi00]. Statistic [BP03, CG03, CG04, CF04, Kru01, Nov05]. Statistical [Erm04, FP06, GP05, Mas04a, Pas05, Ruk01]. Statistics [AB02, Bel05a, BK05, Bor01, Eic05, Gad04, Mas03, MF08, Sha03]. Statulevicius [Ano00c, Ano05b]. Stein [ASVV07]. Step [AB06, BO07]. Sticky [Vys06]. Stochastic [Amo05, Amo06, Ano00h, Ano01h, Ano02j, Arc03, Arc04, Bak03, BNS08, BK09, BKP04, Boi03, BB06, BER05, BE06, BE08, CFS08, Che02b, Dar05, DR02, Gra00, GSY07, IK00, IK01, KN08, KS02, LMS01, LN00, Leb02, Leb03b, LRT00, Mak00, Mak04, MM09, Mal01b, Man09, Mas04a, MS07b, Nas00, Nas07, OK02, RS00, Ryb01b, SY04, Shi09d, VM00, Ver00, Vys06, Yar07]. Stock [Mas04b]. Stopped [DKF08]. Stopping [AS09, BL01, Dom02, GPS01, HP09, KR04, MP04, NS05, Nov09, Pes09, Uru05]. Strategies [MO01]. Strengthening [Kru06]. Strict [FVY00]. Strictly [CHL02, Smo00]. String [Pus03]. Strip [Bor09a]. Strong [BE06, FK01, GZ09, Gri00, IK08a, Kol04b, Kru03, Min01, Pet09, Rot08, Sha06b, Shi02, Shu06, Min00]. Strongly [Sun08]. Structural [Eng04]. Structure [GT07a, KK06]. Structures [Rot08]. Student [BK05, BP03, CG03, CG04, KN08, Nov05].

Sub [VK07]. Sub-Gaussian [VK07]. Subcritical [Vat04]. Subexponential [BY04, BK02b, KV05, Kor02, Mal01b, Rog00]. Submartingales [AR06, Kru09]. Subordination [Kol09]. Subordinators [BNPA08, VD06]. Subsequences [BG04]. Sufficiency [BN00]. Suite [Ruk01]. Sum [ABM08, Bel02, FJT02, Nag05, PU02, Roz03a, Roz04b]. Summands [Bel02, KN03, PU02]. Sums [BB05a, BM01b, BM07a, BM07b, BM09a, BM09b, DKF08, Eic01, Fat04, Fro03, Fro04, HLRV03, IK08a, KN03, Kon02, Kor02, Kov02, Koz02, KB02, Kru07, LY06, Leb00a, Leb05, Lou03, Mat09, MP08, Nag02b, Nag02a, Nov05, PS08, PN03, Rag06a, Roz04a, Roz05, Roz08, Roz09a, Rya01, ZN04, Zhu00a]. Super [KPT01]. Super-Replication [KPT01]. Supercritical [Pet03]. Superlarge [BM07a, BM07b, Roz04b, Roz08]. supermartingale [Rok09]. Supermartingales [Nag07b, Pre09]. Superposition [BN01]. Superprocesses [Dyn00]. Sure [CF04, IL00, VD06]. Surely [BR06b, BR06c]. Surface [Pug09a]. Survey [Ver04]. Survival [GK01]. Susceptible [Joh06]. Switching [KL05, MO01]. Symmetric [HM08, KS07b, Mat09, Nas07, Shu01, VT07, Zie01]. Symmetrization [GK05]. Symmetry [LL03]. Synchronization [MM06, Man09]. Syracuse [BP01]. System [Bak00, Kop01, LMS01, Man09, Yar07]. Systems [BDG^{+}07, CS01, FVW03, Joh06, OK02, Ryb01b, Vys06].

Tableaux [VT07]. Tails [BB02, Bor03a, Bor04b, DKF08, Leb03a, Leb05, Leb08, Mat09, Wac08]. Tauberian [Rog03, Yak04]. Technical [Pas05]. Temperature [DK06]. Term [Bog01, HT07, Khr07]. Terms [Sen07]. Test [LL03, Sof02, Sof05]. Testing [BAH01, DS03, Ing01, Rom01, Ruk01]. Tests [BS05, BD08, BD09, Erm07, Ruk07, Sal03]. Their [Bel02, BF00, Khr02, Kor01, Leb00b, PU02, Shi08, VM00, Zie01]. Theorem [BO03, BR01, BK02a, Bor04a, CPP02, Chi02a, Chi03, FP06, FG05, FVY00, Jaj06, KL08, Mak04, Mik09, Min01, NV06b, PS04, RR01, Rog02, Rok05, Roz00b, SC00, Sze06, Vat04, VT05, Yak08, Bak02, BG04, Fel01, GS05, Gri03, Joh06, Kam07, Mla06, Roz03a, Sen07, Sen09, Sha06a, Sob08, Yar09]. Theorems [BM01b, CF04, Eng04, Fro04, Fro05, GT07a, IL00, Kop01, Leb00b, Leb00a, Mik03, MS07a, PS08, Rag06a, Rog03, Sch03, Ser06, She00, Shu01, Shu06, Tim03, Tim05, VD04, VD08, Wac08, Yak04, ZN04, RS05, Su02]. Theoretic [Joh06, Min01, Min00]. Theorm [Chi02b]. Theory [BK05, Gap00, GP05, KR00, Kol04a, Leb00b, Mas03, Shi00, Sol00, TV05]. There [SS00]. Thermo [Mas07]. Thermo- [Mas07]. Thresholding [CGLT04]. Time [BNS06, BR06b, BR06c, Fat09, GT07b, KS02, Kam09, KS08, KPT01, Kol09, Ma09, MM06, MP04, PKJ07, PSV08, Rag06b, Rag06a, Rok06, Rok08, Vak02, VK09, Vin07, Mit06, Rok09]. Time-Changed [PKJ07]. Time-Varying [PSV08]. Times [CU04, Dom02, Kol09, Nov04, Nov09, PT09, Rad07, Shi09c, Ste03]. Timewise [BNS08]. Toeplitz [GS05]. Topologies [RS05]. Torus [MF08]. Total [Afa04, Khr07, Noq00, VWF08]. Trace [DK01]. Traditional [Mag00]. Trajectory [Bor03a, Bor04b]. Transaction [KPT01, Rok08]. Transform [CPP02]. Transformations [Ale06, BNPA08, BK06a, BK06b, Lan07, LS02, OE09]. Transformed [Bak02]. Transience [AK04]. Transient [Ale04, Bor06, GPL08, Sak05]. Transition [BRS08, BRS09, dHW07]. Transitions [MM06, Mas04a]. Translate [Noq00, Now05]. Transportation [BK06a, OK02]. Trend [Lob06b, Nos02]. Trials [Sal03]. Triangular [Ale06, Sha06b]. Turbulent [BNS08]. Twentieth [Ano01h, Ano02j]. Two [Ans06, DK06, Gaa08, HP09, Kal02, MS07a, MF08, Orl09, OE09, Ruz02, SS00, Sta02]. Two-Dimensional [MF08]. Two-Sided [HP09]. Two-Temperature [DK06]. Two-Type [Gaa08]. Type [AB02, BN01, BNPA03, Bob01, Bor03b, BH08, Fat04, FV01, Gaa08, Mas00, Mik09, PL07, Roz09a, Sha05, Ver01, Ben05, Mik03, Pin06]. Types [MS07a, Sta02]. Typical [FV01]. Tyurin [Rom01].

Uhlenbeck [BN01, BR02, Fat09, Nov04]. Ultimate [GPS01]. UMVUEs [KK06]. Unbounded [Jaj06]. Unconditional [Kva01, VK07]. Unified [Fro05]. Uniform [Kho06, Kul03, KP03, Kuz06, Roz00a, She03, Shu06]. Unimodal [Kuz07]. Unimprovability [Mak03]. Unions [Sch03]. Uniqueness [Bak03, Che02b, DW06]. Unit [Ber09]. Unitary [Amo05, GT03]. Unknown [Bor09b, DS03, Nov00b]. Unusual [Zap06]. Upcrossings [PS04]. Upper [MPW06, Neu00, She07]. Upsilon [BNPA08].

V [Ano06c]. V. [ABV^{+}07]. Value [Fat03, Ser00]. Valued [DM08, Esq08, Fat04, HLRV03, KY05, Leb02, Leb03b, RS05, Shi02, Su02]. Values [Leb06, Sha05]. Varadhan [CCHH08]. Variable [KS07a]. Variables [BD07, CHLV08, CS01, Chu00, DM08, Esq08, Fro03, Fro04, IS02, Kam07, KN03, Kor02, Kru07, Lev01, Lou02, Lou03, MP01, Mat09, Nag01, Nag02b, Nag02a, PL07, Pet02, Pet09, Rag06a, Roz03a, Roz04a, Roz04b, Roz05, Roz08, Ruz02, Sha05, She00, Vla05, Zhu00a]. Variance [BS05, Nec00, Rad07, Ryc03]. Variance-Minimizing [Rad07]. Variances [Bor06]. Variation [BNS06, Noq00, Rog03].

Variational [AS09]. Variograms [Ma09]. Various [CM04]. Varying [BB02, Leb03a, Leb08, PSV08]. Vasilievich [Ano01a, Ano03b]. Vector [GT07b, KR01a, Smo00, Su02]. Vectors [BB05a, BM01b, BM07a, BM07b, BM09a, BM09b, Kov02, KP04, KPK06a, MP08, PK08, Ser00, ZN04, Zai01, Zai02a, Zai02b]. Velocities [BNS08]. Velocity [KO02]. Version [EL00, Kam09, Rok05, Zai01, Zai02a, Zai02b]. Vertical [CSY03]. via [BP01, GV07, Roo04]. Viewpoint [Mas03]. Visitation [Vla05]. Vitautas [Ano00c, Ano05b]. Vladimirovich [Ano01b]. Volumes [Ano02b]. Vyacheslav [Ano03b].

Wald [FP06]. Walk

[Ale04, Bor09a, BF00, CSY03, Kal03, Mis06b, Nad07, PK03, Rag06b, Vis01]. Walks [BB02, BB05b, Bor06, GPL08, Kol09, MmVT03, NS05, Sak05, Sui02, VT05]. Watson [Afa07, Afa08, FV01, NV06b, NV06a, VD04, VD05, VWF08, Wac08]. Wave [Pus03]. Wavelet [GY07]. Waves [VD09]. Weak [Ali00, BER05, BE08, Gri03, KP02, KB02, Leb03b, LB04, Roz04a, Roz05]. Weakly [Lou02, Sha03, Sha05, VK07]. Weighted [HLRV03]. Which [Ing01, Uru05, Vin07]. Whose [IK08a]. Wiener [DB04, Fat03]. Wigner [KS07c]. Willinger [Rok05]. Window [VWF08]. Wise [CHLV08]. Wishart [GV07]. within [VWF08]. Without [GPS01, BB06, BE06]. Wolfowitz [FP06]. Wrong [VK09].

XIXth [Ano00h].

Ya [Ano07f]. Yaglom [Ht09, Klo07b]. Years [PS09]. Young [VT07]. Yu. [Ano06c]. Yurii [Ano01a].

Zero [Hes04, IS02]. Zero-One [Hes04]. Zeros [Imh03, Zap06]. Zeta [KL08]. Zeta-Function [KL08]. Zolotarev [ABV^{+}07]. Zone [BM09a, BM09b]. Zubkov [Kha08a].

References

Alberlink:2002:LTB

[AB02] I. B. Alberlink and V. Bentkus.

Lyapunov-type bounds for Ustatistics. Theory of Probability and its Applications, 46(4):571– 588, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97929.

Ahlswede:2006:NSF

[AB06] R. Ahlswede and V. Blinovsky.

On the number of step functions with restrictions. Theory of Probability and its Applications, 50(4):537–560, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arkashov:2008:LDP

[ABM08] N. S. Arkashov, I. S. Borisov, and A. A. Mogul’skii. Large deviation principle for partial sum processes of moving averages. Theory of Probability and its Applications, 52(2):181–208, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasyeva:2007:BVM

[ABV^{+}07] L. G. Afanasyeva, E. V. Bulinskaya, O. V. Viskov, I. E. Ibragimov, R. I. Ivanovskaya, V. Yu. Korolev, V. M. Kruglov, Yu. V.

Prokhorov, B. A. Sevastyanov, V. V. Senatov, M. V. Khatuntseva, S. Khokhlov, D. M. Chibisov, and A. N. Shiryaev. On the 75th birthday of V. M. Zolotarev. Theory of Probability and its Applications, 51 (4):680–682, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasyev:2004:RBM

[Afa04] V. I. Afanasyev. On the ratio between the maximal and total numbers of individuals in a critical branching process in a random environment. Theory of Probability and its Applications, 48(3):384–399, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98050.

Afanasyev:2007:ALB

[Afa07] V. I. Afanasyev. Arcsine law for branching processes in a random environment and Galton– Watson processes. Theory of Probability and its Applications, 51(3):401–414, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Afanasyev:2008:GWP

[Afa08]  V. I. Afanasyev.  Galton–

Watson process attaining a high level. Theory of Probability and its Applications, 52(3):509–515, ???? 2008. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Amosov:2003:MHQ

[AH03] G. G. Amosov and A. S. Holevo.

On the multiplicativity hypothesis for quantum communication channels. Theory of Probability and its Applications, 47(1): 123–127, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97950.

Aurzada:2009:SDS

[AILvZ09] F. Aurzada, I. A. Ibragimov, M. A. Lifshits, and J. H. van Zanten. Small deviations of smooth stationary Gaussian processes. Theory of Probability and its Applications, 53 (4):697–707, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Albeverio:2004:RTS

[AK04] S. Albeverio and M. V. Kozlov. On the recurrence and transience of state-dependent branching processes in random environment. Theory of Probability and its Applications, 48(4):575–591, December 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98068.

Adell:2008:EMS

[AL08] J. A. Adell and A. Lekuona. Estimates for moduli of smoothness of distribution functions. Theory of Probability and its Applications, 52(1):148–152, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aleskeviciene:2004:TPR

[Ale04] A. A. Aleskeviciené. Transient phenomena in a random walk. Theory of Probability and its Applications, 48(1):1– 18, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98030.

Alexandrova:2006:CTT

[Ale06] D. E. Alexandrova. Convergence of triangular transformations of measures. Theory of Probability and its Applications, 50(1): 113–118, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Aliev:2000:NCD

[Ali00] F. A. Aliev. New characterization of discrete distributions through weak records. Theory of Probability and its Applications, 44(4):756–761, December 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97797.

Abrams:2003:IRF

[ALL^{+}03] A. Abrams, H. Landau, Z. Landau, J. Pommersheim, and E. Zaslow. An iterated random function with Lipschitz number one. Theory of Probability and its Applications, 47(2): 190–201, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97964.

Amosov:2005:MPG

[Amo05] G. G. Amosov. On Markovian perturbations of the group of unitary operators associated with a stochastic process with stationary increments. Theory of Probability and its Applications, 49(1):123–132, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98090.

Amosov:2006:MPQ

[Amo06] G. G. Amosov. On Markovian perturbations of quantum stochastic processes with stationary increments. Theory of Probability and its Applications, 50(4):650–658, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:2000:NSLa

[Ano00a] Anonymous. News of scientific life. Theory of Probability and its Applications, 44(3):622– 631, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97785.

Anonymous:2000:NSLc

[Ano00b] Anonymous. News of scientific life. Theory of Probability and its Applications, 44(4):806– 807, December 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97800.

Anonymous:2000:BVA

[Ano00c]  Anonymous.  On the 70th

birthday of Vitautas Antanovich Statulevicius. Theory of Probability and its Applications, 44(4):755, December 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97792.

Anonymous:2000:RBa

[Ano00d] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 44(1): 222–224, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97751.

Anonymous:2000:RBb

[Ano00e] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 44(2): 428–431, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97769.

Anonymous:2000:RBc

[Ano00f] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 44(3):613– 621, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97784.

Anonymous:2000:RBd

[Ano00g] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 44(4):798– 805, December 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97799.

Anonymous:2000:XIS

[Ano00h] Anonymous. The XIXth International Seminar on Stability Problems for Stochastic Models. Theory of Probability and its Applications, 44(1): 140–196, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97747.

Anonymous:2001:BYV

[Ano01a] Anonymous. 70th birthday of Yurii Vasilievich Prokhorov. Theory of Probability and its Applications, 45(1):1, March 2001. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97801.

Anonymous:2001:AVS

[Ano01b] Anonymous. Anatolii Vladimirovich Skorokhod. Theory of Probability and its Applications, 45(4): 643, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97856.

Anonymous:2001:LE

[Ano01c] Anonymous. Letter to the Editors. Theory of Probability and its Applications, 45 (2):362, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97833.

Anonymous:2001:NSLa

[Ano01d] Anonymous. News of scientific Life. Theory of Probability and its Applications, 45(1): 173–174, March 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97815.

Anonymous:2001:NSLb

[Ano01e] Anonymous. News of scientific life. Theory of Probability and its Applications, 45(3): 546, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97847.

Anonymous:2001:RBa

[Ano01f] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 45(2): 359–361, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97832.

Anonymous:2001:RBb

[Ano01g] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 45(4):725– 730, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97869.

Anonymous:2001:TIS

[Ano01h] Anonymous. Twentieth international seminar on stability problems for stochastic models. Theory of Probability and its Applications, 45(4):699– 724, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97868.

Anonymous:2002:BAS

[Ano02a] Anonymous. 80th birthday of Andrei Sergeevich Monin. Theory of Probability and its Applications, 46(3):515, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97913.

Anonymous:2002:AIV

[Ano02b]  Anonymous.  Author index

to volumes 41-45. Theory of Probability and its Applications, 46(3):554–570, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97922.

Anonymous:2002:BC

[Ano02c] Anonymous. Birthday congratulations. Theory of Probability and its Applications, 46 (2):386, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97904.

Anonymous:2002:NSL

[Ano02d] Anonymous. News of scientific Life. Theory of Probability and its Applications, 46(4):753– 756, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97940.

Anonymous:2002:NAD

[Ano02e]  Anonymous.  Nikolai Alek-

sandrovich Dmitriev (obituary). Theory of Probability and its Applications, 46(3):549– 553, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97921.

Anonymous:2002:RIA

[Ano02f] Anonymous. Regina Ivanovskaya’s anniversary. Theory of Probability and its Applications, 46 (2):335, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97895.

Anonymous:2002:RBa

[Ano02g] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 46(1): 185–191, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97886.

Anonymous:2002:RBb

[Ano02h] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 46(2): 380–385, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97903.

Anonymous:2002:RBc

[Ano02i] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 46(4):750– 752, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97939.

Anonymous:2002:TIS

[Ano02j] Anonymous. Twentieth international seminar on stability problems for stochastic models. Theory of Probability and its Applications, 46(1):171– 184, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97885.

Anonymous:2003:IAI

[Ano03a] Anonymous. Ildar Abdullovich Ibragimov. Theory of Probability and its Applications, 47 (3):488, ???? 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:2003:MPV

[Ano03b] Anonymous. In memoriam: Professor Vyacheslav Vasilievich Sazonov. Theory of Probability and its Applications, 47 (2):299, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97967.

Anonymous:2003:NSL

[Ano03c] Anonymous. News of scientific life. Theory of Probability and its Applications, 47(4):742– 744, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98015.

Anonymous:2003:RBa

[Ano03d] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 47(1): 184–187, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97959.

Anonymous:2003:RBb

[Ano03e] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 47(2): 375–376, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97979.

Anonymous:2003:RBc

[Ano03f] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 47(3):556– 565, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97996.

Anonymous:2004:MPA

[Ano04a] Anonymous. In memoriam:

Paul Andre Meyer. Theory of Probability and its Applications, 48(1):190, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98032.

Anonymous:2004:LE

[Ano04b] Anonymous. Letter to the Editors. Theory of Probability and its Applications, 48(4): 752, December 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98082.

Anonymous:2004:NB

[Ano04c] Anonymous. New books. Theory of Probability and its Applications, 48(1):189, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98029.

Anonymous:2004:NSL

[Ano04d] Anonymous. News of scientific Life. Theory of Probability and its Applications, 48(2): 379–381, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98049.

Anonymous:2004:BBS

[Ano04e] Anonymous. On the 80th birthday of B. A. Sevastyanov. Theory of Probability and its Applications, 48(4):697–702, December 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98075.

Anonymous:2004:ICK

[Ano04f] Anonymous. On the International Conference “Kolmogorov and Contemporary Mathematics”. Theory of Probability and its Applications, 48(3): 383, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98066.

Anonymous:2004:RB

[Ano04g] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 48(3):569– 573, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98065.

Anonymous:2005:A

[Ano05a]  Anonymous.  Anniversaries.

Theory of Probability and its Applications, 49(2):372, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98113.

Anonymous:2005:MVA

[Ano05b] Anonymous. In memoriam:

Vitautas Antanovich Statulevicius. Theory of Probability and its Applications, 49(1): 187, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98098.

Anonymous:2005:LE

[Ano05c] Anonymous. Letter to the Editors. Theory of Probability and its Applications, 49(2): 373–375, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98114.

Anonymous:2005:NSL

[Ano05d] Anonymous. News of scientific life. Theory of Probability and its Applications, 49(3):553– 559, September 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98127.

Anonymous:2005:BBD

[Ano05e] Anonymous. On the 80th birthday of E. B. Dynkin. Theory of Probability and its Applications, 49(2):310, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98105.

Anonymous:2006:B

[Ano06a] Anonymous. Birthday. Theory of Probability and its Applications, 50(4):724, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:2006:LE

[Ano06b] Anonymous. Letters to the editors. Theory of Probability and its Applications, 50(1):170, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:2006:BYV

[Ano06c] Anonymous. On the 75th birthday of Yu. V. Prokhorov. Theory of Probability and its Applications, 50(1):112, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:2006:RB

[Ano06d] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 50 (1):169, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:2007:LEb

[Ano07a] Anonymous. Letter to the Editors. Theory of Probability and its Applications, 51(3):561–563, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:2007:LEa

[Ano07b] Anonymous. Letters to the editors. Theory of Probability and its Applications, 51(2):398–400, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:2007:NB

[Ano07c] Anonymous. New books. Theory of Probability and its Applications, 51(3):564–565, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:2007:NSLa

[Ano07d] Anonymous. News of scientific life. Theory of Probability and its Applications, 51(2):397, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:2007:NSLb

[Ano07e] Anonymous. News of scientific life. Theory of Probability and its Applications, 51(3):558–560, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:2007:BAY

[Ano07f] Anonymous. On the 70th birthday of Academician Ya. G. Sinai. Theory of Probability and its Applications, 51(1):1, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:2007:RBa

[Ano07g] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 51 (3):554–557, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anonymous:2007:RBb

[Ano07h] Anonymous. Reviews and bibliography. Theory of Probability and its Applications, 51 (4):737–738, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Anshin:2006:PSE

[Ans06] A. B. Anshin. On the probability of simultaneous extremes of two Gaussian nonstationary processes. Theory of Probability and its Applications, 50(3): 353–366, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Alsmeyer:2006:MPI

[AR06] G. Alsmeyer and U. Rosler.

Maximal phi-inequalities for nonnegative submartingales. Theory of Probability and its Applications, 50(1):118– 128, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Arcones:2003:LDP

[Arc03] M. A. Arcones. The large deviation principle for stochastic processes. Part I. Theory of Probability and its Applications, 47(4):567–583, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98003.

Arcones:2004:LDP

[Arc04] M. A. Arcones. The large deviation principle for stochastic processes. II. Theory of Probability and its Applications, 48 (1):19–44, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98028.

Aleskeviciene:2002:LDA

[AS02] A. Aleskeviciene and V. Statulvicius. Large deviations in the approximation by the Poisson law. Theory of Probability and its Applications, 46(4):589– 602, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97923.

Aleskeviciene:2005:ACP

[AS05] A. K. Aleskeviciene and V. A. Statulevicius. Approximation by the compound Poisson law. Theory of Probability and its Applications, 49(3):519– 526, September 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98122.

Arkin:2009:VAO

[AS09] V. I. Arkin and A. D. Slastnikov. A variational approach to optimal stopping problems for diffusion processes. Theory of Probability and its Applications, 53 (3):467–480, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ahmed:2007:AEC

[ASVV07] S. E. Ahmed, A. K. Md. E. Saleh, A. I. Volodin, and I. N. Volodin. Asymptotic expansion of the coverage probability of James–Stein estimators. Theory of Probability and its Applications, 51(4):683–695, ???? 2007.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Burnashev:2001:STH

[BAH01] M. V. Burnashev, Sh. Amari, and T. S. Han. On some testing of hypothesis problems with information constraints. Theory of Probability and its Applications, 45(4):558–568, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97850.

Bakhtin:2000:CAS

[Bak00] V. I. Bakhtin. Cramér asymptotics in a system with slow and fast Markovian motions. Theory of Probability and its Applications, 44(1):1–17, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97731.

Bakhtin:2002:FCL

[Bak02] Yu. Yu. Bakhtin. A functional Central Limit Theorem for transformed solutions of the multidimensional Burgers equation with random initial data. Theory of Probability and its Applications, 46(3):387– 405, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97906.

Bakhtin:2003:EUS

[Bak03] Yu. Bakhtin. Existence and uniqueness of a stationary solution of a nonlinear stochastic differential equation with memory. Theory of Probability and its Applications, 47(4):684– 688, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98005.

Borovkov:2002:PLD

[BB02] A. A. Borovkov and K. A. Borovkov. On probabilities of large deviations for random walks. I. Regularly varying distribution tails. Theory of Probability and its Applications, 46(2):193–213, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97887.

Barbe:2005:SLD

[BB05a] Ph. Barbe and M. Broniatowski. On sharp large deviations for sums of random vectors and multidimensional Laplace approximation. Theory of Probability and its Applications, 49(4):561–588, December 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98134.

Borovkov:2005:PLD

[BB05b] A. A. Borovkov and K. A.

Borovkov. On probabilities of large deviations for random walks. II. Regular exponentially decaying distributions. Theory of Probability and its Applications, 49(2): 189–206, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98099.

Borisov:2006:CSI

[BB06] I. S. Borisov and A. A. Bystrov.

Constructing a stochastic integral of a nonrandom function without orthogonality of the noise. Theory of Probability and its Applications, 50 (1):53–74, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Benhenni:2004:AEM

[BD04] K. Benhenni and R. Drouilhet.

Adjusted Euler–MacLaurin predictor for integrating smooth spatial processes. Theory of Probability and its Applications, 48(3):506–520, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98060.

Breton:2007:LIP

[BD07] J.-C. Breton and Yu. A. Davydov. Local invariance principle for independent and identically distributed random variables. Theory of Probability and its Applications, 51(2):256–278, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Brodsky:2008:MST

[BD08] B. E. Brodsky and B. S. Darkhovsky. Minimax sequential tests for many composite hypotheses. I. Theory of Probability and its Applications, 52 (4):565–579, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Brodsky:2009:MST

[BD09] B. E. Brodsky and B. S. Darkhovsky. Minimax sequential tests for many composite hypotheses. II. Theory of Probability and its Applications, 53 (1):1–12, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bertini:2007:LDE

[BDG^{+}07] L. Bertini, A. De Sole, D. Gabrielli, G. Jona-Lasinio, and C. Landim. Large deviations of the empirical current in interacting particle systems. Theory of Probability and its Applications, 51 (1):2–27, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bryc:2005:MCC

[BDK05] W. Bryc, A. Dembo, and A. Kagan. On the maximum correlation coefficient. Theory of Probability and its Applications, 49(1):132–138, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98096.

Buckdahn:2006:BSD

[BE06] R. Buckdahn and H. J. Engelbert. A backward stochastic differential equation without strong solution. Theory of Probability and its Applications, 50 (2):284–289, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Buckdahn:2008:CWS

[BE08] R. Buckdahn and H.-J. Engelbert. On the continuity of weak solutions of backward stochastic differential equations. Theory of Probability and its Applications, 52(1):152–160, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belomestnyi:2002:PRS

[Bel02] D. V. Belomestnyi. On the problem of reconstructing a summands distribution by their sum. Theory of Probability and its Applications, 46(2):336– 341, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97896.

Belomestnyi:2005:RGD

[Bel05a] D. V. Belomestnyi. Reconstruction of the general distribution by the distribution of some statistics. Theory of Probability and its Applications, 49(1): 1–15, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98083.

Belopolskaya:2005:PAS

[Bel05b] Ya. Belopolskaya. A probabilistic approach to a solution of nonlinear parabolic equations. Theory of Probability and its Applications, 49(4):589– 611, December 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98128.

Bentkus:2005:LTB

[Ben05] V. Bentkus. A Lyapunovtype bound in R^{d}. Theory of Probability and its Applications, 49(2):311–323, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98112.

Buckdahn:2005:WSB

[BER05] R. Buckdahn, H. J. Engelbert, and A. Ruascanu. On weak solutions of backward stochastic differential equations. Theory of Probability and its Applications, 49(1):16–50, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98087.

Berdikulov:2009:MPO

[Ber09] M. A. Berdikulov. Markov processes on order-unit spaces. Theory of Probability and its Applications, 53(1):136–144, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2000:EOA

[BF00] A. A. Borovkov and S. G.

Foss. Estimates for overshooting an arbitrary boundary by a random walk and their applications. Theory of Probability and its Applications, 44(2): 231–253, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97753.

Burnaev:2009:AOS

[BFS09] E. V. Burnaev, E. A. Feinberg, and A. N. Shiryaev. On asymptotic optimality of the second order in the minimax quickest detection problem of drift change for Brownian motion. Theory of Probability and its Applications, 53(3):519–536, ???? 2009. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Bobkov:2004:CLT

[BG04] S. G. Bobkov and F. Götze. On the Central Limit Theorem along subsequences of noncorrelated observations. Theory of Probability and its Applications, 48(4):604–621, December 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98073.

Breton:2008:FRS

[BH08] J.-C. Breton and C. Houdré. On finite range stable-type concentration. Theory of Probability and its Applications, 52 (4):543–564, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Burnashev:2002:SEP

[BHA02] M. V. Burnashev, T. S. Han, and S. Amari. On some estimation problems with information constraints. Theory of Probability and its Applications, 46(2): 214–225, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97888.

Borovkov:2001:LDP

[BK01] A. A. Borovkov and D. A. Korshunov. Large-deviation probabilities for one-dimensional Markov Chains.Part 2: Prestationary distributions in the exponential case. Theory of Probability and its Applications, 45(3):379–405, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97835.

Bogachev:2002:OMP

[BK02a] V. I. Bogachev and A. V.

Kolesnikov. Open mappings of probability measures and the Skorokhod representation theorem. Theory of Probability and its Applications, 46(1): 20–38, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97870.

Borovkov:2002:LDP

[BK02b] A. A. Borovkov and D. A. Korshunov. Large-deviation probabilities for one-dimensional Markov Chains.Part 3: Prestationary distributions in the subexponential case. Theory of Probability and its Applications, 46(4):603–618, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97924.

Bening:2003:NER

[BK03] V. E. Bening and V. Yu. Korolev. Nonparametric estimation of the ruin probability for generalized risk processes. Theory of Probability and its Applications, 47(1): 1–16, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97941.

Bening:2005:ASD

[BK05] V. E. Bening and V. Yu. Korolev. On an application of the Student distribution in the theory of probability and mathematical statistics. Theory of Probability and its Applications, 49(3):377–391, September 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98115.

Bogachev:2006:IAC

[BK06a] V. I. Bogachev and A. V. Kolesnikov. Integrability of absolutely continuous transformations of measures and applications to optimal mass transportation. Theory of Probability and its Applications, 50(3): 367–385, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bogachev:2006:NTC

[BK06b] V. I. Bogachev and A. V. Kolesnikov. Nonlinear transformations of convex measures. Theory of Probability and its Applications, 50(1):34–52, January 2006. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Baurdoux:2009:SSS

[BK09] E. J. Baurdoux and A. E. Kyprianou. The Shepp–Shiryaev stochastic game driven by a spectrally negative Lévy process. Theory of Probability and its Applications, 53(3):481–499, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belkina:2004:SOF

[BKP04] T. A. Belkina, Yu. M. Kabanov, and E. L. Presman. On a stochastic optimality of the feedback control in the LQG-Problem. Theory of Probability and its Applications, 48(4):592– 603, December 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98069.

Beibel:2001:OSR

[BL01] M. Beibel and H. R. Lerche.

Optimal stopping of regular diffusions under random discounting. Theory of Probability and its Applications, 45(4):547– 557, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97852.

Borisov:2001:ARL

[BM01a] I. S. Borisov and D. V. Mironov.

An asymptotic representation for the likelihood ratio for multidimensional samples with discontinuous densities. Theory of Probability and its Applications, 45(2):289–300, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97823.

Borovkov:2001:ILL

[BM01b] A. A. Borovkov and A. A. Mogulskii. Integro-local limit theorems including large deviations for sums of random vectors. II. Theory of Probability and its Applications, 45(1): 3–22, March 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97802.

Borovkov:2007:LSDa

[BM07a] A. A. Borovkov and A. A. Mogulskii. On large and superlarge deviations of sums of independent random vectors under Cramér’s condition. I. Theory of Probability and its Applications, 51(2):227–255, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2007:LSDb

[BM07b] A. A. Borovkov and A. A. Mogulskii. On large and superlarge deviations of sums of independent random vectors under Cramér’s condition. II. Theory of Probability and its Applications, 51(4):567–594, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2009:LDSa

[BM09a] A. A. Borovkov and A. A.

Mogulskii. On large deviations of sums of independent random vectors on the boundary and outside of the Cramér zone. I. Theory of Probability and its Applications, 53(2):301–311, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2009:LDSb

[BM09b] A. A. Borovkov and A. A.

Mogulskii. On large deviations of sums of independent random vectors on the boundary and outside of the Cramér zone. II. Theory of Probability and its Applications, 53(4):573–593, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Barndorff-Nielsen:2000:NAS

[BN00] O. E. Barndorff-Nielsen. Lnonformation, L-ancillarity, and L-sufficiency. Theory of Probability and its Applications, 44(1):197–201, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97749.

Barndorff-Nielsen:2001:SOU

[BN01] O. E. Barndorff-Nielsen. Superposition of Ornstein–Uhlenbeck type processes. Theory of Probability and its Applications, 45(2):175–194, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97816.

Barnsdorff-Nielsen:2003:ETM

[BNPA03] O. E. Barnsdorff-Nielsen and V. Perez-Abreu. Extensions of type G and marginal infinite divisibility. Theory of Probability and its Applications, 47(2): 202–218, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97965.

Barndorff-Nielsen:2008:MSR

[BNPA08] O. E. Barndorff-Nielsen and V. Pérez-Abreu. Matrix subordinators and related upsilon transformations. Theory of Probability and its Applications, 52(1):1–23, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Barndorff-Nielsen:2006:PVT

[BNS06] O. E. Barndorff-Nielsen and N. Shephard. Power variation and time change. Theory of Probability and its Applications, 50(1):1–15, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Barndorff-Nielsen:2008:SDE

[BNS08] O. E. Barndorff-Nielsen and J. Schmiegel. A stochastic differential equation framework for the timewise dynamics of turbulent velocities. Theory of Probability and its Applications, 52 (3):372–388, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Baltrunas:2003:SOR

[BO03] A. Baltrunas and E. Omey.

Second-order renewal theorem in the finite-means case. Theory of Probability and its Applications, 47(1):127–132, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97956.

Bezhaeva:2007:EMG

[BO07] Z. I. Bezhaeva and V. I. Oseledets. Erdős measures for the Golden ratio and 2-step Markov chains. Theory of Probability and its Applications, 51 (1):28–41, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bobkov:2001:SGP

[Bob01] S. G. Bobkov. Some generalizations of Prokhorov’s results on Khinchin-type inequalities for polynomials. Theory of Probability and its Applications, 45(4):644–647, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97858.

Bobkov:2003:LPB

[Bob03] S. G. Bobkov. Localization proof of the Bakry–Ledoux isoperimetric inequality and some applications. Theory of Probability and its Applications, 47(2):308–314, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97968.

Boguslavskaya:2001:OLT

[Bog01] E. V. Boguslavskaya. On optimization of long-term irreversible investments in a diffusion model. Theory of Probability and its Applications, 45(4):647–658, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97859.

Bogatyrev:2003:NEE

[Bog03] S. A. Bogatyrev. A nonuniform estimate for the error in short asymptotic expansions in Hilbert space. Theory of Probability and its Applications, 47(4):689–692, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98006.

Boikov:2003:CLM

[Boi03] A. V. Boikov. The Cramer– Lundberg model with stochastic premium process. Theory of Probability and its Applications, 47(3):489–493, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97987.

Borodin:2000:DFC

[Bor00] A. N. Borodin. Distribution of functionals of certain non-Markovian processes. Theory of Probability and its Applications, 44(3):433–452, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97770.

Borovskikh:2001:NAS

[Bor01] Yu. V. Borovskikh. On a normal approximation of Ustatistics. Theory of Probability and its Applications, 45(3):406– 423, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97836.

Borovkov:2003:ACP

[Bor03a] A. A. Borovkov. Asymptotics of crossing probability of a boundary by the trajectory of a Markov chain. Heavy tails of jumps. Theory of Probability and its Applications, 47(4):584– 608, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97998.

Borovkov:2003:NDT

[Bor03b]  K. Borovkov.  A note on

diffusion-type approximation to branching processes in random environments. Theory of Probability and its Applications, 47(1):132–138, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97957.

Borisov:2004:NDT

[Bor04a] I. S. Borisov. A note on Dobrushin’s theorem and couplings in Poisson approximation in Abelian groups. Theory of Probability and its Applications, 48(3):521–528, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98057.

Borovkov:2004:ACP

[Bor04b] A. A. Borovkov. Asymptotics of crossing probability of a boundary by the trajectory of a Markov chain. exponentially decaying tails. Theory of Probability and its Applications, 48(2): 226–242, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98036.

Borovkov:2006:TPR

[Bor06]  A. A. Borovkov.  Tran-

sient phenomena for random walks with nonidentically distributed jumps with infinite variances. Theory of Probability and its Applications, 50(2): 199–213, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borisov:2007:SAH

[Bor07] A. V. Borisov. State analysis of hidden Markov models governed by special jump processes. Theory of Probability and its Applications, 51(3):518–528, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2008:AP

[Bor08] K. A. Borovkov. The Abel prize. Theory of Probability and its Applications, 52(3):371, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borisov:2009:NDN

[Bor09a] I. S. Borisov. A note on the distribution of the number of crossings of a strip by a random walk. Theory of Probability and its Applications, 53(2):312–316, ???? 2009. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2009:LSC

[Bor09b] A. A. Borovkov. Large sample change-point estimation when distributions are unknown. Theory of Probability and its Applications, 53(3):402–418, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Borovkov:2001:ESP

[BP01] K. A. Borovkov and D. Pfeifer. Estimates for the Syracuse problem via a probabilistic model. Theory of Probability and its Applications, 45(2): 300–310, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97824.

Bloznelis:2003:SOB

[BP03] M. Bloznelis and H. Putter. Second-order and bootstrap approximation to Student’s tstatistic. Theory of Probability and its Applications, 47(2): 300–307, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97974.

Bogachev:2001:GKT

[BR01] V. Bogachev and M. Röckner. A generalization of Khasminskii’s theorem on the existence of invariant measures for locally integrable drifts. Theory of Probability and its Applications, 45(3):363–378, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97834.

Bercu:2002:SLD

[BR02] B. Bercu and A. Rouault.

Sharp large deviations for the Ornstein–Uhlenbeck process. Theory of Probability and its Applications, 46(1):1–19, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97873.

Bassetti:2006:APR

[BR06a] F. Bassetti and E. Regazzini.

Asymptotic properties and robustness of minimum dissimilarity estimators of location-scale parameters. Theory of Probability and its Applications, 50(2): 171–186, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belkina:2006:OPAa

[BR06b] T. A. Belkina and V. I. Rotar. On optimality in probability and almost surely for processes with a communication property. I. The discrete time case. Theory of Probability and its Applications, 50 (1):16–33, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Belkina:2006:OPAb

[BR06c] T. A. Belkina and V. I. Rotar. On optimality in probability and almost surely for processes with a communication property. II. The continuous time case. Theory of Probability and its Applications, 50 (2):187–198, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bogachev:2006:GRB

[BRS06] V. I. Bogachev, M. Rockner, and S. V. Shaposhnikov. Global regularity and bounds for solutions of parabolic equations for probability measures. Theory of Probability and its Applications, 50(4):561–581, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bogachev:2008:EDS

[BRS08] V. I. Bogachev, M. Röckner, and S. V. Shaposhnikov. Estimates of densities of stationary distributions and transition probabilities of diffusion processes. Theory of Probability and its Applications, 52(2):209–236, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Bogachev:2009:PDT

[BRS09] V. I. Bogachev, M. Röckner, and S. V. Shaposhnikov. Positive densities of transition probabilities of diffusion processes. Theory of Probability and its Applications, 53(2):194–215, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Boldin:2005:STA

[BS05] M. V. Boldin and W. Stute.

On sign tests in ARMA models with possibly infinite error variance. Theory of Probability and its Applications, 49(3):392– 413, September 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98116.

Butucea:2008:SODa

[BT08a] C. Butucea and A. B. Tsybakov. Sharp optimality in density deconvolution with dominating bias. I. Theory of Probability and its Applications, 52 (1):24–39, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Butucea:2008:SODb

[BT08b] C. Butucea and A. B. Tsybakov. Sharp optimality in density deconvolution with dominating bias. II. Theory of Probability and its Applications, 52 (2):237–249, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Burnaev:2009:DPP

[Bur09] E. V. Burnaev. Disorder problem for Poisson process in generalized Bayesian setting. Theory of Probability and its Applications, 53(3):500–518, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Baltrunas:2004:SOA

[BY04] A. Baltrunas and A. L. Yakymiv. Second-order asymptotic behavior of subexponential infinitely divisible distributions. Theory of Probability and its Applications, 48(4):703– 710, December 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98076.

Cherny:2006:NSL

[CB06] A. S. Cherny and Translated By. News of scientific life. Theory of Probability and its Applications, 50(2):348–350, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cavazos-Cadena:2008:CAV

[CCHH08] R. Cavazos-Cadena and D. Hernández-Hernández. Contractive approximations for the Varadhan’s function on a finite Markov chain. Theory of Probability and its Applications, 52(2):315–323, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

CantoECastro:2001:ROS

[CDT01] L. Canto E. Castro, L. De Haan, and M. G. Temido. Rarely observed sample maxima. Theory of Probability and its Applications, 45(4):658– 662, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97865.

Cekanavicius:2000:CPA

[Cek00] V. Cekanavicius. On compound Poisson approximations under moment restrictions. Theory of Probability and its Applications, 44(1):18–28, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97734.

Cekanavivcius:2003:MCD

[Cek03] V. Cekanavivcius. On multivariate compound distributions. Theory of Probability and its Applications, 47(3):493– 506, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97992.

Chuprunov:2004:ASL

[CF04] A. Chuprunov and I. Fazekas.

Almost sure limit theorems for the Pearson statistic. Theory of Probability and its Applications, 48(1):140–147, March 2004. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98021.

Carbone:2008:BSD

[CFS08] R. Carbone, B. Ferrario, and M. Santacroce. Backward stochastic differential equations driven by Càdlàg martingales. Theory of Probability and its Applications, 52(2):304–314, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:2003:MDS

[CG03] G. P. Chistyakov and F. Gotze. Moderate deviations for Student’s statistic. Theory of Probability and its Applications, 47(3):415–428, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97984.

Chistyakov:2004:BMD

[CG04] G. P. Chistyakov and F. Götze. On bounds for moderate deviations for Student’s statistic. Theory of Probability and its Applications, 48(3):528– 535, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98061.

Cavalier:2004:BTS

[CGLT04] L. Cavalier, Y. Golubev, O. Lepski, and A. Tsybakov. Block thresholding and sharp adaptive estimation in severely illposed inverse problems. Theory of Probability and its Applications, 48(3):426–446, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98055.

Cherny:2001:CSI

[Che01] A. S. Cherny. Convergence of some integrals associated with Bessel processes. Theory of Probability and its Applications, 45(2):195–209, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97817.

Cherny:2002:FCP

[Che02a] A. S. Cherny. Families of consistent probability measures. Theory of Probability and its Applications, 46(1):118–121, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97880.

Cherny:2002:ULP

[Che02b] A. S. Cherny. On the uniqueness in law and the pathwise uniqueness for stochastic differential equations. Theory of Probability and its Applications, 46(3):406– 419, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97909.

Cherny:2008:NSL

[Che08a] A. Cherny. News of scientific life. Theory of Probability and its Applications, 52(3):539–542, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Cherny:2008:PCR

[Che08b] A. S. Cherny. Pricing with coherent risk. Theory of Probability and its Applications, 52 (3):389–415, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Chistyakov:2002:NAEb

[Chi02a] G. P. Chistyakov. A new asymptotic expansion and asymptotically best constants in Lyapunov’s theorem. II. Theory of Probability and its Applications, 46(3):516–522, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97917.

Chistyakov:2002:NAEa

[Chi02b] G. P. Chistyakov. A new asymptotic expansion and asymptotically best constants in Lyapunov’s theorm. I. Theory of Probability and its Applications, 46(2):226–242, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97893.

Chistyakov:2003:NAE

[Chi03] G. P. Chistyakov. A new asymptotic expansion and asymptotically best constants in Lyapunov’s theorem. III. Theory of Probability and its Applications, 47(3):395–414, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97982.

Chun:2002:LLS

[CHL02] Su Chun, Taizhong Hu, and Hanying Liang. On the logarithm law for strictly stationary and negatively associated arrays. Theory of Probability and its Applications, 46(2): 369–379, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97902.

Chen:2008:CCA

[CHLV08] P. Chen, T.-C. Hu, X. Liu, and A. Volodin. On complete convergence for arrays of rowwise negatively associated random variables. Theory of Probability and its Applications, 52 (2):323–328, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Christopeit:2004:NPA

[Chr04] N. Christopeit. A note on the pricing of American options. Theory of Probability and its Applications, 48(1):131– 140, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98022.

Chuprunov:2000:CLM

[Chu00] A. N. Chuprunov. On convergence in law of maxima of independent identically distributed random variables with random coefficients. Theory of Probability and its Applications, 44(1): 93–97, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97743.

Chibisov:2007:BB

[CIM^{+}07] D. M. Chibisov, I. A. Ibragimov, V. P. Maslov, B. A. Sevastyanov, A. N. Shiryaev, N. N. Vakhania, I. S. Borisov, V. I. Lotov, and A. A. Mogul’skii. On the 75th birthday of A. A. Borovkov. Theory of Probability and its Applications, 51 (2):225–226, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Csorgo:2003:MSL

[CM03] S. Csorgo and Z. Megyesi. Merging to semistable laws. Theory of Probability and its Applications, 47(1):17–33, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97947.

Cherny:2004:MME

[CM04] A. S. Cherny and V. P. Maslov.

On minimization and maximization of entropy in various disciplines. Theory of Probability and its Applications, 48(3):447– 464, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98052.

Cacoullos:2002:ADT

[CPP02] T. Cacoullos, N. Papadatos, and V. Papathanasiou. An application of a density transform and the local limit theorem. Theory of Probability and its Applications, 46(4):699– 707, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97936.

Chobanyan:2001:EMI

[CS01] S. Chobanyan and H. Saleh. Exact maximal inequalities for exchangeable systems of random variables. Theory of Probability and its Applications, 45(3):424– 435, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97837.

Carr:2004:BPI

[CS04] P. Carr and M. Schröder. Bessel processes, the integral of geometric Brownian motion, and Asian options. Theory of Probability and its Applications, 48(3):400–425, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98054.

Cherny:2003:LBH

[CSY03] A. S. Cherny, A. N. Shiryaev, and M. Yor. Limit behavior of the “horizontal-vertical” random walk and some extensions of the Donsker–Prokhorov invariance principle. Theory of Probability and its Applications, 47(3):377–394, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97983.

Cherny:2004:STM

[CU04] A. S. Cherny and M. A. Urusov. Separating times for measures on filtered spaces. Theory of Probability and its Applications, 48(2):337–347, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98048.

Cwiklinska:2009:ALI

[Ćwi09] M. B. Ćwiklińska. Asymptotics in the law of the iterated logarithm. Theory of Probability and its Applications, 53(4):707–711, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Darkhovskii:2005:NAS

[Dar05] B. S. Darkhovskii. A new approach to the stochastic recovery problem. Theory of Probability and its Applications, 49 (1):51–64, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98085.

Dorogovtsev:2004:RMG

[DB04] A. A. Dorogovtsev and V. V.

Bakun. Random mappings and a generalized additive functional of a Wiener process. Theory of Probability and its Applications, 48(1):63–79, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98026.

DelMoral:2000:MIP

[DD00] P. Del Moral and M. Doisy.

Maslov idempotent probability calculus. II. Theory of Probability and its Applications, 44(2): 319–332, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97759.

DelMoral:2007:SPC

[DDP07] P. Del Moral, A. Doucet, and G. W. Peters. Sharp propagation of chaos estimates for Feynman–Kac particle models. Theory of Probability and its Applications, 51(3):459–485, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Debicki:2006:SPG

[Deb06] K. Debicki. Some properties of generalized Pickands constants. Theory of Probability and its Applications, 50(2):290– 298, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dermoune:2005:DPG

[Der05] A. Dermoune. d-dimensional pressureless gas equations. Theory of Probability and its Applications, 49(3):540–545, September 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98125.

denHollander:2007:LTC

[dHW07] F. den Hollander and S. G. Whittington. Localization transition for a copolymer in an emulsion. Theory of Probability and its Applications, 51(1): 101–141, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

DiNunno:2004:HEC

[Di 04] G. Di Nunno. Hölder equality for conditional expectations with application to linear monotone operators. Theory of Probability and its Applications, 48(1):177–181, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98018.

Dynkin:2001:RBT

[DK01] E. B. Dynkin and S. E.

Kuznetsov. Rough boundary trace for solutions of Lu = ψu. Theory of Probability and its Applications, 45(4):662– 667, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97857.

Dudnikova:2006:TTP

[DK06] T. V. Dudnikova and A. I.

Komech. On a two-temperature problem for the Klein–Gordon equation. Theory of Probability and its Applications, 50(4): 582–611, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Denisov:2008:LLD

[DKF08] D. E. Denisov, D. A. Korshunov, and S. G. Foss. Lower limits for distribution tails of randomly stopped sums. Theory of Probability and its Applications, 52(4):690–699, ???? 2008.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dedecker:2008:CRL

[DM08] J. Dedecker and F. Merlevède. Convergence rates in the law of large numbers for Banach-valued dependent variables. Theory of Probability and its Applications, 52(3):416–438, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dokuchaev:2004:PIE

[Dok04] N. G. Dokuchaev. Parabolic Ito equations with nonsmooth nonlinearity and duality approach. Theory of Probability and its Applications, 48(1): 45–62, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98027.

Domansky:2002:ROS

[Dom02] V. K. Domansky. Randomized optimal stopping times for a class of stopping games. Theory of Probability and its Applications, 46(4):708–717, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97931.

DiNunno:2002:MMS

[DR02] G. Di Nunno and Yu. A. Rozanov. On measurable modification of stochastic functions. Theory of Probability and its Applications, 46(1):122– 127, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97883.

Darkhovski:2003:HTP

[DS03] B. S. Darkhovski and M. Staroswiecki.

The hypotheses testing problem under unknown parameter. Theory of Probability and its Applications, 47(4):609– 622, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/99999.

Douady:2000:PCD

[DSY00] R. Douady, A. N. Shiryaev, and M. Yor. On probability characteristics of “downfalls” in a standard Brownian motion. Theory of Probability and its Applications, 44(1): 29–38, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97730.

Davydov:2003:CRG

[DT03] Yu. Davydov and E. Thilly.

Convex rearrangements of Gaussian processes. Theory of Probability and its Applications, 47(2):219–235, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97960.

Dai:2006:CNE

[DW06] J. G. Dai and R. J. Williams. Correctional note to “Existence and Uniqueness of Semimartingale Reflecting Brownian Motions in Convex Polyhedrons”. Theory of Probability and its Applications, 50(2): 346–347, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Dynkin:2000:ELE

[Dyn00] E. B. Dynkin. Exit laws and excessive functions for superprocesses. Theory of Probability and its Applications, 44(4):762– 767, December 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97798.

Eichelsbacher:2001:LDP

[Eic01] P. Eichelsbacher. Large deviations for partial sums Uprocesses in dependent cases. Theory of Probability and its Applications, 45(4):569–588, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97853.

Eichelsbacher:2005:RLD

[Eic05] P. Eichelsbacher. Refined large deviations for von Mises statistics. Theory of Probability and its Applications, 49(1):139– 148, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98097.

Epifani:2000:FAV

[EL00] I. Epifani and A. Lijoi. A finitely additive version of the law of the iterated logarithm. Theory of Probability and its Applications, 44(4):633–649, December 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97788.

Engibaryan:2004:AST

[Eng04] N. B. Engibaryan. Asymptotic and structural theorems for the Markov renewal equation. Theory of Probability and its Applications, 48(1):80– 92, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98025.

Ermakov:2004:AES

[Erm04] M. S. Ermakov. On asymptotically efficient statistical inference for moderate deviation probabilities. Theory of Probability and its Applications, 48(4):622–641, December 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98070.

Ermakov:2007:ISS

[Erm07] M. S. Ermakov. Importance sampling for simulation of large and moderate deviation probabilities of tests and estimators. Theory of Probability and its Applications, 51(2):279–290, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Esquivel:2008:PGF

[Esq08] M. L. Esquível. Probability generating functions for discrete real-valued random variables. Theory of Probability and its Applications, 52(1):40–57, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fakhrutdinova:2009:CES

[Fak09] T. K. Fakhrutdinova. On conditions of extinction for the statedependent branching processes in a random environment. Theory of Probability and its Applications, 53(4):711–716, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Fatalov:2003:ALD

[Fat03] V. R. Fatalov. Asymptotics of large deviations for Wiener random fields in L^{p}-norm, nonlinear Hammerstein equations, and high-order hyperbolic boundary-value problems. Theory of Probability and its Applications, 47(4):623–636, December 2003. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98002.

Fatalov:2004:PLT

[Fat04] V. R. Fatalov. Precise Laplacetype asymptotics for moderate deviations of the distributions of sums of independent Banach-valued random elements. Theory of Probability and its Applications, 48(4):642– 663, December 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98072.

Fatalov:2009:OTE

[Fat09] V. R. Fatalov. Occupation time and exact asymptotics of distributions of L^{p}-functionals of the Ornstein–Uhlenbeck processes, p > 0. Theory of Probability and its Applications, 53 (1):13–36, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feldman:2001:MSD

[Fel01] G. M. Feldman. More on the Skitovich–Darmous Theorem for finite Abelian groups. Theory of Probability and its Applications, 45(3):507–511, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97845.

Feldman:2005:SDT

[FG05] G. M. Feldman and P. Graczyk. On the Skitovich–Darmois theorem for discrete Abelian groups. Theory of Probability and its Applications, 49(3):527–531, September 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98123.

Fuchs:2002:ERS

[FJT02] A. Fuchs, A. Joffe, and J. Teugels. Expectation of the ratio of the sum of squares to the square of the sum: Exact and asymptotic results. Theory of Probability and its Applications, 46(2):243–255, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97891.

Fazekas:2001:GAS

[FK01] I. Fazekas and O. Klesov. A general approach to the strong law of large numbers. Theory of Probability and its Applications, 45(3):436–449, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97838.

Frolov:2008:PSD

[FMS08] A. N. Frolov, A. I. Martikainen, and J. Steinebach. On probabilities of small deviations for compound renewal processes. Theory of Probability and its Applications, 52(2):328–337, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Feinberg:2006:DWW

[FP06] E. A. Feinberg and A. B. Piunovskiy. On the Dvoretzky– Wald–Wolfowitz theorem on nonrandomized statistical decisions. Theory of Probability and its Applications, 50(3): 463–466, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Frolov:2003:ABL

[Fro03] A. N. Frolov. On the asymptotic behavior of the large increments of sums of independent random variables. Theory of Probability and its Applications, 47(2): 315–323, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97973.

Frolov:2004:LTI

[Fro04] A. N. Frolov. Limit theorems for increments of sums of independent random variables. Theory of Probability and its Applications, 48(1):93– 107, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98024.

Frolov:2005:ULT

[Fro05] A. N. Frolov. Unified limit theorems for increments of processes with independent increments. Theory of Probability and its Applications, 49(3):531– 540, September 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98124.

Fleischmann:2001:DTT

[FV01] K. Fleischmann and V. A. Vatutin. Deviations from typical type proportions in critical multitype Galton–Watson processes. Theory of Probability and its Applications, 45(1): 23–40, March 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97803.

Fleischmann:2003:BSL

[FVW03] K. Fleischmann, V. A. Vatutin, and A. Wakolbinger. Branching systems with long-living particles at the critical dimension. Theory of Probability and its Applications, 47(3):429– 454, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97980.

Freiman:2000:LLT

[FVY00] G. Freiman, A. M. Vershik, and Yu. V. Yakubovich. A local limit theorem for random strict partitions. Theory of Probability and its Applications, 44(3):453– 468, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97771.

Gaas:2008:PLJ

[Gaa08] V. V. Gaas. Poisson limit joint distributions for a random allocation scheme of two-type particles. Theory of Probability and its Applications, 52(2):337–345, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gabriyelyan:2001:ASD

[Gab01] S. S. Gabriyelyan. On the automorphically stable distributions on Abelian groups. Theory of Probability and its Applications, 45(3):512–517, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97843.

Gadasina:2004:BEI

[Gad04]  L. V. Gadasina.  Berry–

Esseen inequalities for Ustatistics. Theory of Probability and its Applications, 48(1): 147–152, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98020.

Gamkrelidze:2001:IMC

[Gam01] N. G. Gamkrelidze. An inequality for a multidimensional characteristic function. Theory of Probability and its Applications, 45(1):133–135, March 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97811.

Gaposhkin:2000:IMS

[Gap00] V. F. Gaposhkin. On the integrability of a maximal square function in ergodic theory. Theory of Probability and its Applications, 44(2):386–394, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97762.

Gaposhkin:2002:SEP

[Gap02] V. F. Gaposhkin. Some examples of the problem of εdeviatations for stationary sequences. Theory of Probability and its Applications, 46(2): 341–346, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97897.

Gaposhkin:2007:PEM

[Gap07] V. F. Gaposhkin. Precise estimates of the metric entropy for the set of arithmetic averages of quasi-stationary processes. Theory of Probability and its Applications, 51(4):695–704, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gaposhkin:2009:EEM

[Gap09] V. F. Gaposhkin. Exact estimates of the metric entropy of the averages for some classes of stationary sequences. Theory of Probability and its Applications, 53(1):37–58, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Grishechkin:2003:MQR

[GDLH03] S. Grishechkin, M. Devetsikiotis, I. Lambadaris, and C. Hobbs. Multistability in queues with retransmission and its relationship with large deviations in branching processes. Theory of Probability and its Applications, 47(1):139– 150, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97958.

Golubev:2002:MEP

[GE02] G. G. Golubev and F. N. Enikeeva. On the minimax estimation problem of a fractional derivative. Theory of Probability and its Applications, 46(4):619– 635, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97925.

Glonti:2007:NEG

[GHKT07] O. Glonti, P. Harremoës, Z. Khechinashvili, and F. Topsøe. Nash equilibrium in a game of calibration. Theory of Probability and its Applications, 51 (3):415–426, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Geiger:2001:SPC

[GK01] J. Geiger and G. Kersting. The survival probability of a critical branching process in a random environment. Theory of Probability and its Applications, 45(3):517–525, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97844.

Gushchin:2005:RMS

[GK05] A. A. Gushchin and U. Küchler. On recovery of a measure from its symmetrization. Theory of Probability and its Applications, 49(2):323–333, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98106.

Gulinsky:2000:ELD

[GL00] O. V. Gulinsky and R. Sh. Liptser. An example of large deviations for a stationary process. Theory of Probability and its Applications, 44(1):201– 217, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97748.

Graf:2009:QPM

[GL09] S. Graf and H. Luschgy. Quantization for probability measures in the Prokhorov metric. Theory of Probability and its Applications, 53(2):216–241, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gnedenko:2006:OEM

[Gne06] B. D. Gnedenko. On one extension of a martingale. Theory of Probability and its Applications, 50(4):659–662, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Glukhova:2005:ACT

[GP05] S. I. Glukhova and E. A.

Palkin. Application of catastrophe theory for multimodal distributions statistical analysis. Theory of Probability and its Applications, 49(3):414– 428, September 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98117.

Guillotin-Plantard:2008:TRW

[GPL08] N. Guillotin-Plantard and A. Le Ny. Transient random walks on 2D-oriented lattices. Theory of Probability and its Applications, 52(4):699–711, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Graversen:2001:SBM

[GPS01] S. E. Graversen, G. Peskir, and A. N. Shiryaev. Stopping Brownian motion without anticipation as close as possible to its ultimate maximum. Theory of Probability and its Applications, 45(1): 41–50, March 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97807.

Grandits:2000:MMS

[Gra00] P. Grandits. On martingale measures for stochastic processes with independent increments. Theory of Probability and its Applications, 44(1): 39–50, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97735.

Grin:2000:SAS

[Gri00] A. G. Grin’. On strong attraction of stationary sequences to a normal law. Theory of Probability and its Applications, 44(4):768–775, December 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97793.

Grin:2003:MCW

[Gri03] A. G. Grin’. On the minimal condition of weak dependency in the Central Limit Theorem for stationary sequences. Theory of Probability and its Applications, 47(3):506–510, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97988.

Ginovyan:2005:CLT

[GS05] M. S. Ginovyan and A. A.

Sahakyan. On the Central Limit Theorem for Toeplitz quadratic forms of stationary sequences. Theory of Probability and its Applications, 49(4):612– 628, December 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98129.

Grandits:2007:RAA

[GS07] P. Grandits and C. Summer.

Risk averse asymptotics and the optional decomposition. Theory of Probability and its Applications, 51(2):325–334, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Graversen:2007:PSI

[GSY07] S. Graversen, A. N. Shiryaev, and M. Yor. On the problem of stochastic integral representations of functionals of the Brownian motion. II. Theory of Probability and its Applications, 51 (1):65–77, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gotze:2003:RCS

[GT03] F. Gotze and A. N. Tikhomirov. Rate of convergence to the semicircular law for the Gaussian unitary ensemble. Theory of Probability and its Applications, 47(2):323–330, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97975.

Gotze:2007:LTS

[GT07a] F. Götze and A. N. Tikhomirov. Limit theorems for spectra of random matrices with martingale structure. Theory of Probability and its Applications, 51 (1):42–64, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gurevich:2007:MAT

[GT07b] B. M. Gurevich and A. A. Tempelman. Multifractal analysis of time averages for continuous vector functions on configuration space. Theory of Probability and its Applications, 51 (1):78–91, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gushchin:2008:END

[Gus08] A. A. Gushchin. On an extension of the notion of fdivergence. Theory of Probability and its Applications, 52 (3):439–455, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Graczyk:2007:MWP

[GV07] P. Graczyk and L. Vostrikova.

The moments of Wishart processes via Itô calculus. Theory of Probability and its Applications, 51(4):609–625, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Gamet:2000:ENS

[GW00] C. Gamet and M. Weber. Entropy numbers of some ergodic averages. Theory of Probability and its Applications, 44(4):650– 668, December 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97789.

Gama:2007:EEW

[GY07] J. M. R. Gama and V. Yurinsky. An exponential estimate for a wavelet density estimator. Theory of Probability and its Applications, 51(4):595–608, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Granovsky:2005:LBS

[GZ05] B. L. Granovsky and A. I. Zeifman. On the lower bound of the spectrum of some mean-field models. Theory of Probability and its Applications, 49(1): 148–155, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98092.

Gotze:2009:BRS

[GZ09] F. Götze and A. Yu. Zaitsev. Bounds for the rate of strong approximation in the multidimensional invariance principle. Theory of Probability and its Applications, 53(1):59–80, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Harlamov:2001:DDF

[Har01] B. P. Harlamov. On the distribution density of the first exit point of a diffusion process form a small neighborhood of its initial position. Theory of Probability and its Applications, 45(3):450–465, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97839.

Hess:2004:CZO

[Hes04] K. T. Hess. Conditional zeroone laws. Theory of Probability and its Applications, 48(4):711– 718, December 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98080.

Hu:2003:RCC

[HLRV03] T. C. Hu, D. Li, A. Rosalsky, and A. I. Volodin. On the rate of complete convergence for weighted sums of arrays of Banach space valued random elements. Theory of Probability and its Applications, 47(3):455– 468, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97985.

Hadjiliadis:2006:OAO

[HM06] O. Hadjiliadis and V. Moustakides. Optimal and asymptotically optimal CUSUM rules for change point detection in the Brownian motion model with multiple alternatives. Theory of Probability and its Applications, 50(1):75–85, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hipp:2008:NAS

[HM08] C. Hipp and L. Mattner. On the normal approximation to symmetric binomial distributions. Theory of Probability and its Applications, 52(3):516–523, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Holevo:2004:EAC

[Hol04] A. S. Holevo. Entanglementassisted capacities of constrained quantum channels. Theory of Probability and its Applications, 48(2):243–255, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98041.

Holevo:2005:AES

[Hol05] A. S. Holevo. Asymptotic estimation of a shift parameter of a quantum state. Theory of Probability and its Applications, 49(2):207–220, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98104.

Holevo:2007:CCA

[Hol07] A. S. Holevo. Complementary channels and the additivity problem. Theory of Probability and its Applications, 51 (1):92–100, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hadjiliadis:2009:BCS

[HP09] O. Hadjiliadis and H. V. Poor. On the best 2-CUSUM stopping rule for quickest detection of two-sided alternatives in a Brownian motion model. Theory of Probability and its Applications, 53(3):537–547, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Holevo:2006:CEC

[HS06] A. S. Holevo and M. E. Shirokov. Continuous ensembles and the capacity of infinitedimensional quantum channels. Theory of Probability and its Applications, 50(1):86–98, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Hall:2007:MTM

[HT07] A. Hall and M. G. Temido. On the maximum term of MA and Max-AR models with margins in Anderson’s class. Theory of Probability and its Applications, 51(2):291–304, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Holevo:2009:MAM

[Ht09] A. S. Holevo and translator. In memoriam: Akiva Moiseevich Yaglom (1921–2007). Theory of Probability and its Applications, 53(1):134–135, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:2004:EMR

[Ibr04] I. A. Ibragimov. Estimation of multivariate regression. Theory of Probability and its Applications, 48(2):256–272, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98038.

Ibragimov:2007:CSE

[Ibr07] I. A. Ibragimov. On censored sample estimation of a multivariate analytic probability density. Theory of Probability and its Applications, 51(1):142–154, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:2000:EPC

[IK00] I. A. Ibragimov and R. V. Khas’minskii. Estimation problems for coefficients of stochastic partial differential Equations.Part II. Theory of Probability and its Applications, 44(3):469–494, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97772.

Ibragimov:2001:EPC

[IK01] I. A. Ibragimov and R. Z. Khas’minskii. Estimation problems for coefficients of stochastic partial differential Equations.Part III. Theory of Probability and its Applications, 45(2):210–232, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97818.

Indlekofer:2008:SLL

[IK08a] K.-H. Indlekofer and O. I. Klesov. Strong law of large numbers for multiple sums whose indices belong to a sector with function boundaries. Theory of Probability and its Applications, 52(4):711–719, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ivanova:2008:OER

[IK08b] G. P. Ivanova and V. Ya. Kondratiev. On one estimate of the ruin probability. Theory of Probability and its Applications, 52(2):345–350, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ibragimov:2000:ASL

[IL00] I. A. Ibragimov and M. A. Lifshits. On almost sure limit theorems. Theory of Probability and its Applications, 44(2): 254–272, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97756.

Ivchenko:2003:GMD

[IM03] G. I. Ivchenko and Yu. I.

Medvedev. Goncharov’s method and its development in the analysis of different models of random permutations. Theory of Probability and its Applications, 47(3):518–526, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97989.

Imhof:2003:RPP

[Imh03] L. A. Imhof. Random polynomials with a prescribed number of real zeros. Theory of Probability and its Applications, 47(3):511– 518, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97994.

Ingster:2001:THW

[Ing01] Yu. I. Ingster. On testing a hypothesis which is close to a simple hypothesis. Theory of Probability and its Applications, 45(2):310–323, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97825.

Ibragimov:2002:ECR

[IS02] R. Ibragimov and Sh. Sharakhmetov. The exact constant in the Rosenthal inequality for random variables with mean zero. Theory of Probability and its Applications, 46(1):127– 132, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97876.

Jajte:2006:PET

[Jaj06] R. Jajte. Pointwise ergodic theorem for unbounded operators in L_{2}. Theory of Probability and its Applications, 50(4): 662–676, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Janvresse:2003:BSR

[Jan03] E. Janvresse. Bounds on semigroups of random rotations on SO(n). Theory of Probability and its Applications, 47(3):526– 532, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97995.

Jedidi:2008:SPM

[Jed08] W. Jedidi. Stable processes, mixing, and distributional properties. I. Theory of Probability and its Applications, 52(4): 580–593, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jedidi:2009:SPM

[Jed09] W. Jedidi. Stable processes, mixing, and distributional properties. II. Theory of Probability and its Applications, 53 (1):81–105, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Johnson:2004:CPC

[Joh04] O. Johnson. Convergence of the Poincaré constant. Theory of Probability and its Applications, 48(3):535–541, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98062.

Johnson:2006:ITC

[Joh06] O. Johnson. An informationtheoretic Central Limit Theorem for finitely susceptible FKG systems. Theory of Probability and its Applications, 50(2): 214–224, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Jaoshvili:2009:EOH

[JP09] V. Jaoshvili and O. Purtukhia. An extension of the Ocone– Haussmann–Clark formula for the compensated Poisson processes. Theory of Probability and its Applications, 53(2):316–321, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kalinkin:2002:EPB

[Kal02] A. V. Kalinkin. On the extinction probability of a branching process with two kinds of interaction of particles. Theory of Probability and its Applications, 46(2):347–352, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97898.

Kalinkin:2003:APB

[Kal03]  A. V. Kalinkin.  Absorp-

tion probability at the border of a random walk in a quadrant and a branching process with interaction of particles. Theory of Probability and its Applications, 47(3):469– 487, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97981.

Kallsen:2004:LM

[Kal04] J. Kallsen. σ-localization and σmartingales. Theory of Probability and its Applications, 48(1): 152–163, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98031.

Kaminski:2007:CLT

[Kam07] M. Kaminski. Central Limit Theorem for certain classes of dependent random variables. Theory of Probability and its Applications, 51(2):335–342, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kamenov:2009:BVR

[Kam09] A. A. Kamenov. Bachelierversion of Russian option with a finite time horizon. Theory of Probability and its Applications, 53(3):548–557, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Karymov:2005:DLD

[Kar05] D. N. Karymov. On the decomposition of lattice distributions into convolutions of Poisson signed measures. Theory of Probability and its Applications, 49(3):545–552, September 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98126.

Kruglov:2002:WCR

[KB02] V. M. Kruglov and Zhang Bo. Weak convergence of random sums. Theory of Probability and its Applications, 46(1): 39–57, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97871.

Khatuntseva:2008:BMZ

[Kha08a] M. V. Khatuntseva. On the 60th birthday of OF A. M. Zubkov. Theory of Probability and its Applications, 52(2):303, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2008:BDM

[Kha08b] M. V. Khatuntseva. On the 70th birthday of D. M. Chibisov. Theory of Probability and its Applications, 52(1):147, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2008:BI

[Kha08c] M. V. Khatuntseva. On the 75th birthday of I. A. Ibragimov. Theory of Probability and its Applications, 52(4):689, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khatuntseva:2009:BBS

[Kha09] M. V. Khatuntseva. On the 85th birthday of B. A. Sevastyanov. Theory of Probability and its Applications, 53(4):696, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khokhlov:2002:POR

[Kho02] V. I. Khokhlov. Polynomials orthogonal with respect to the multinomial distribution and the factorial-power formalism. Theory of Probability and its Applications, 46(3):529– 536, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97919.

Khokhlov:2006:UDS

[Kho06] V. I. Khokhlov. The uniform distribution on a sphere in R^{S}. Properties of projections. I. Theory of Probability and its Applications, 50(3): 386–399, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khrennikov:2002:IPT

[Khr02] A. Yu. Khrennikov. Interpretations of probability and their p-adic extensions. Theory of Probability and its Applications, 46(2):256–273, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97892.

Khrennikov:2007:FTP

[Khr07] A. Yu. Khrennikov. A formula of total probability with the interference term and the Hilbert space representation of the contextual Kolmogorovian model. Theory of Probability and its Applications, 51(3): 427–441, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kagan:2006:SUC

[KK06] A. Kagan and M. Konikov. The structure of the UMVUEs from categorical data. Theory of Probability and its Applications, 50(3):466–473, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Khasminskii:2003:OES

[KL03] R. Khasminskii and R. Liptser. Online estimation of a smooth regression function. Theory of Probability and its Applications, 47(3):541–550, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97990.

Korolyuk:2005:PAI

[KL05] V. S. Korolyuk and N. Limnios. Poisson approximation of increment processes with Markov switching. Theory of Probability and its Applications, 49(4):629– 644, December 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98133.

Kacinskaite:2008:GDL

[KL08] R. Kacinskaite and A. Laurincikas. A general discrete limit theorem in the space of analytic functions for the Matsumoto zeta-function. Theory of Probability and its Applications, 52(3):523–531, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klokov:2007:LBM

[Klo07a] S. A. Klokov. Lower bounds of mixing rate for a class of Markov processes. Theory of Probability and its Applications, 51(3):528–535, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klotz:2007:SRI

[Klo07b] L. Klotz. Some remarks on an interpolation problem of A. M. Yaglom. Theory of Probability and its Applications, 51 (2):342–350, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kerkyacharian:2008:NEA

[KLP08] G. Kerkyacharian, O. Lepski, and D. Picard. Nonlinear estimation in anisotropic multiindex denoising. sparse case. Theory of Probability and its Applications, 52(1):58–77, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klykova:2003:LDN

[Kly03] N. V. Klykova. Limit distribution of a number of coinciding intervals. Theory of Probability and its Applications, 47(1): 151–156, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97951.

Kukush:2000:API

[KM00] A. G. Kukush and Yu. S. Mishura. Asymptotic properties of an intensity estimator of an inhomogeneous Poisson process in a combined model. Theory of Probability and its Applications, 44(2): 273–292, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97757.

Kutoyants:2002:EED

[KN02] Yu. A. Kutoyants and I. Negri. On L_{2} efficiency of an empiric distribution for ergodic diffusion processes. Theory of Probability and its Applications, 46(1): 140–146, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97881.

Khamdamov:2003:RES

[KN03] I. M. Khamdamov and A. V.

Nagaev. On the role of extreme summands in sums of independent random variables. Theory of Probability and its Applications, 47(3):533– 541, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97991.

Kagan:2008:LSM

[KN08] A. Kagan and A. V. Nagaev. A lemma on stochastic majorization and properties of the Student distribution. Theory of Probability and its Applications, 52(1):160–164, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuksin:2003:HES

[KNP03] S. B. Kuksin, N. S. Nadirashvili, and A. L. Piatnitski. Hölder estimates for solutions of parabolic SPDEs. Theory of Probability and its Applications, 47(1):157–163, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97952.

Kolesnik:2002:AFV

[KO02] A. D. Kolesnik and E. Orsingher. Analysis of a finitevelocity planar random motion with reflection. Theory of Probability and its Applications, 46(1):132–140, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97877.

Kobelkov:2005:RPS

[Kob05] S. G. Kobelkov. The ruin problem for the stationary Gaussian process. Theory of Probability and its Applications, 49(1): 155–163, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98093.

Kolmogorov:2004:HKT

[Kol04a]  A. N. Kolmogorov.  From

the heritage of A. N. Kolmogorov: The theory of probability. Theory of Probability and its Applications, 48(2): 191–220, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98033.

Kolmogorov:2004:SIC

[Kol04b] A. N. Kolmogorov. On some inequalities connected to the strong law of large numbers. Theory of Probability and its Applications, 48(2):221–225, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98035.

Kolokoltsov:2009:GCT

[Kol09] V. N. Kolokoltsov. Generalized continuous-time random walks, subordination by hitting times, and fractional dynamics. Theory of Probability and its Applications, 53(4):594–609, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kondratenko:2002:RBR

[Kon02] A. E. Kondratenko. The relation between a rate of convergence of moments of normed sums and the Chebyshev–Hermite models. Theory of Probability and its Applications, 46(2):352– 355, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97900.

Kopytsev:2001:LTN

[Kop01] V. A. Kopytsev. Limit theorems for the number of solutions of a system of random equations. Theory of Probability and its Applications, 45(1): 51–68, March 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97804.

Korolev:2000:APS

[Kor00]  V. Yu. Korolev.  Asymp-

totic properties of sample quantiles constructed from samples with random sizes. Theory of Probability and its Applications, 44(2):394–399, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97763.

Korolev:2001:APE

[Kor01]  V. Yu. Korolev.  Asymp-

totic properties of extrema of compound Cox processes and their applications to some problems of financial mathematics. Theory of Probability and its Applications, 45(1):136– 147, March 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97813.

Korshunov:2002:LDP

[Kor02]  D. A. Korshunov.  Large-

deviation probabilities for maxima of sums of independent random variables with negative mean and subexponential distribution. Theory of Probability and its Applications, 46(2): 355–366, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97901.

Koval:2002:LIL

[Kov02] V. A. Koval’. On the law of the iterated logarithm for operator normed sums of random vectors. Theory of Probability and its Applications, 46(2): 366–368, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97899.

Kozlov:2002:ERP

[Koz02] M. V. Kozlov. On the Erdős– Renyi partial sums: Large deviations, conditional behavior. Theory of Probability and its Applications, 46(4):636– 651, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97926.

Kruglov:2002:WCC

[KP02] V. M. Kruglov and G. N. Petrovskaya. Weak convergence of a certain functional. Theory of Probability and its Applications, 46(4):721–727, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97932.

Kulikova:2003:UDC

[KP03] A. A. Kulikova and Yu. V.

Prokhorov. Uniform distributions on convex sets: Inequality for characteristic functions. Theory of Probability and its Applications, 47(4):700– 701, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98009.

Kulikova:2004:DFP

[KP04] A. A. Kulikova and Yu. V.

Prokhorov. Distribution of the fractional parts of random vectors: The Gaussian case. I. Theory of Probability and its Applications, 48(2): 355–359, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98045.

Kulikova:2005:CAS

[KP05] A. A. Kulikova and Yu. V. Prokhorov. Completely asymmetric stable laws and Benford’s law. Theory of Probability and its Applications, 49(1): 163–169, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98094.

Kulikova:2006:DFP

[KPK06a] A. A. Kulikova, Yu. V. Prokhorov, and V. I. Khokhlov. Distribution of the fractional parts of random vectors: The Gaussian case. II. Theory of Probability and its Applications, 50(4):685–687, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kulikova:2006:HFD

[KPK06b] A. A. Kulikova, Yu. V. Prokhorov, and V. I. Khokhlov. H.F.D. (H-function distribution) and Benford’s Law. I. Theory of Probability and its Applications, 50(2):311–315, January 2006. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Koehl:2001:SRD

[KPT01] P. F. Koehl, H. Pham, and N. Touzi. On super-replication in discrete time under transaction costs. Theory of Probability and its Applications, 45(4):667– 673, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97866.

Kolmogorov:2000:RPR

[KR00] A. N. Kolmogorov and N. G.

Rychkova. Russian poetry rhythm analysis and probability theory. Theory of Probability and its Applications, 44(2): 375–385, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97761.

Korostelev:2001:REV

[KR01a] A. P. Korostelev and A. L.

Rukhin. Recursive estimation of a vector parameter under Bahadur risk. Theory of Probability and its Applications, 45(4):589– 603, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97851.

Kuhne:2001:BCP

[KR01b] R. Kühne and L. Rüschendorf.

On a best choice problem for discounted sequences. Theory of Probability and its Applications, 45(4):673–677, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97867.

Kuhne:2004:AOS

[KR04] R. Kühne and L. Rüschendorf. Approximate optimal stopping of dependent sequences. Theory of Probability and its Applications, 48(3):465–480, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98056.

Kagan:2006:ELP

[KR06] A. Kagan and C. R. Rao. On estimation of a location parameter in the presence of an ancillary component. Theory of Probability and its Applications, 50(1): 129–133, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:2001:ABP

[Kru01] V. M. Kruglov. The asymptotic behavior of the Pearson statistic. Theory of Probability and its Applications, 45(1): 69–92, March 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97805.

Kruglov:2003:EBP

[Kru03] V. M. Kruglov. On extending the Brunk–Prokhorov strong law of large numbers. Theory of Probability and its Applications, 47(2):330–333, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97969.

Kruglov:2004:NPC

[Kru04] V. M. Kruglov. Normal and Poisson convergencies. Theory of Probability and its Applications, 48(2):347–355, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98044.

Kruglov:2005:SCI

[Kru05] V. M. Kruglov. Short communications on inequalities for large deviations in the Bernoulli scheme. Theory of Probability and its Applications, 49(4):695– 700, December 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98137.

Kruglov:2006:SPA

[Kru06] V. M. Kruglov. Strengthening of Prokhorov’s arcsine inequality. Theory of Probability and its Applications, 50(4): 677–684, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:2007:GSP

[Kru07] V. M. Kruglov. Growth of sums of pairwise independent random variables with infinite expectations. Theory of Probability and its Applications, 51(2):359–362, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:2008:NLC

[Kru08] V. M. Kruglov. On the necessity of the Lyapunov condition for normal convergence. Theory of Probability and its Applications, 52(1):164–166, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kruglov:2009:CS

[Kru09] V. M. Kruglov. On convergence of submartingales. Theory of Probability and its Applications, 53(2):321–329, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kallsen:2002:TCR

[KS02] J. Kallsen and A. N. Shiraev. Time change representation of stochastic integrals. Theory of Probability and its Applications, 46(3):522– 528, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97918.

Korolev:2006:ANAa

[KS06a] V. Yu. Korolev and I. G.

Shevtsova. On the accuracy of the normal approximation. I. Theory of Probability and its Applications, 50(2): 298–310, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Korolev:2006:ANAb

[KS06b] V. Yu. Korolev and I. G.

Shevtsova. On the accuracy of the normal approximation. II. Theory of Probability and its Applications, 50(3): 473–482, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kelbert:2007:BDV

[KS07a] M. Kelbert and Yu. M. Suhov.

Branching diffusions on H^{d} with variable fission: The Hausdorff dimension of the limiting set. Theory of Probability and its Applications, 51(1):155–167, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Knutova:2007:APC

[KS07b] E. M. Knutova and S. Ya.

Shatskikh. Asymptotic properties of conditional quantiles for a class of symmetric distributions. Theory of Probability and its Applications, 51(2):350–358, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kozlov:2007:WFD

[KS07c] V. V. Kozlov and O. G. Smolyanov. Wigner function and diffusion in a collisionfree medium of quantum particles. Theory of Probability and its Applications, 51(1):168–181, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kelbert:2008:LTB

[KS08] M. Kelbert and Y. Suhov. Large-time behavior of a branching diffusion on a hyperbolic space. Theory of Probability and its Applications, 52 (4):594–613, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kholopov:2008:RCS

[KTT08] A. A. Kholopov, A. N. Tikhomirov, and D. A. Timushev. Rate of convergence to the semicircle law for the Gaussian orthogonal ensemble. Theory of Probability and its Applications, 52 (1):171–177, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kulikova:2003:ERC

[Kul03]  A. A. Kulikova.  Estimate

of the rate of convergence of probability distributions to a uniform distribution. Theory of Probability and its Applications, 47(4):693–699, December 2003. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98008.

Kulikov:2008:MCC

[Kul08] A. V. Kulikov. Multidimensional coherent and convex risk measures. Theory of Probability and its Applications, 52 (4):614–635, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsova:2006:CUD

[Kuz06] A. Ya. Kuznetsova. On convergence to the uniform distribution. Theory of Probability and its Applications, 50(4): 687–689, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Kuznetsova:2007:MUD

[Kuz07] A. Ya. Kuznetsova. On a mixture of unimodal distributions. Theory of Probability and its Applications, 51(3):535–536, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Klokov:2005:SMR

[KV05] S. A. Klokov and A. Yu.

Veretennikov. On subexponential mixing rate for Markov processes. Theory of Probability and its Applications, 49(1):110– 122, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98084.

Kvaratskhelia:2001:UCG

[Kva01] V. V. Kvaratskhelia. Unconditional convergence of Gaussian random series in Banach spaces. Theory of Probability and its Applications, 45(1): 147–152, March 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97812.

Khrennikov:2005:CRS

[KY05] A. Khrennikov and S. Yamada. On the concept of random sequence with respect to padic valued probabilities. Theory of Probability and its Applications, 49(1):65–76, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98086.

Lange:2007:DNF

[Lan07] A. M. Lange. On the distribution of the number of final particles in a branching process with transformations and pairwise interactions. Theory of Probability and its Applications, 51 (4):704–714, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lopez-Blazquez:2004:LPW

[LB04]  F. López-Blázquez.  Linear

prediction of weak records:

The discrete case. Theory of Probability and its Applications, 48(4):718–723, December 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98081.

Lebedev:2000:LTMb

[Leb00a] A. V. Lebedev. Limit theorems for maxima of independent random sums. Theory of Probability and its Applications, 44(3):558– 561, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97777.

Lebedev:2000:LTMa

[Leb00b] A. V. Lebedev. Limit theorems on maximums of Poisson sequences and their applications to queuing theory. Theory of Probability and its Applications, 44(2):400–404, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97764.

Lebedev:2002:VRM

[Leb02] V. A. Lebedev. L^{p}-valued random measures and good extensions of a stochastic basis. Theory of Probability and its Applications, 46(3):536– 542, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97914.

Lebedev:2003:ESN

[Leb03a] A. V. Lebedev. Extremes of shot-noise fields in the presence of regularly varying tails. Theory of Probability and its Applications, 47(4):702–706, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98010.

Lebedev:2003:EWS

[Leb03b] V. A. Lebedev. On the existence of weak solutions for stochastic differential equations with driving L^{0}-valued measures. Theory of Probability and its Applications, 47(4):637– 648, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98000.

Lebedev:2005:MIS

[Leb05] A. V. Lebedev. Maxima of independent sums in the presence of heavy tails. Theory of Probability and its Applications, 49(4):700–703, December 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98138.

Lebedev:2006:MBP

[Leb06]  A. V. Lebedev.  Maximal

branching processes with nonnegative values. Theory of Probability and its Applications, 50 (3):482–488, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lebedev:2008:ESN

[Leb08] A. V. Lebedev. Extremes of shot-noise fields in the presence of regularly varying tails with index α ∈ (0, 1). Theory of Probability and its Applications, 52(2):350–353, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Levental:2001:MIR

[Lev01] S. Levental. A maximal inequality for real numbers with application to exchangeable random variables. Theory of Probability and its Applications, 45(3):525– 532, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97846.

Leventhal:2002:CBB

[Lev02] S. Leventhal. On coupling of Brownian bridges. Theory of Probability and its Applications, 46(1):146–153, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97882.

Li:2002:SCS

[Li02] Zeng-Hu Li. Skew convolution semigroups and related immigration processes. Theory of Probability and its Applications, 46(2):274–297, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97889.

Lin:2000:IPS

[Lin00] Zhengyan Lin. The invariance principle for some class of Markov chains. Theory of Probability and its Applications, 44(1):136–139, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97746.

Lipchius:2005:PIM

[Lip05] A. A. Lipchius. On a property of independent mappings. Theory of Probability and its Applications, 49(2):333–336, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98107.

Lipchius:2006:NEM

[Lip06] A. A. Lipchius. A note on the equality in the Monge and Kantorovich problems. Theory of Probability and its Applications, 50(4):689–693, January 2006. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

LeBlanc:2003:TSN

[LL03] F. LeBlanc and O. V. Lepski. Test of symmetry in nonparametric regression. Theory of Probability and its Applications, 47(1):34–52, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97948.

Landim:2001:HLN

[LMS01] C. Landim, M. Mourragui, and S. Sellami. Hydrodynamic limit for a nongradient interacting particle system with stochastic reservoirs. Theory of Probability and its Applications, 45(4):604– 623, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97854.

LeBreton:2000:SAP

[LN00] A. Le Breton and A. A. Novikov.

On stochastic approximation procedures with averaging. Theory of Probability and its Applications, 44(3):591–604, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97782.

Lobanov:2006:PCD

[Lob06a] S. Lobanov. Probability characteristics of downfalls of Brownian motion with drift. Theory of Probability and its Applications, 50(3):489–497, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lobanov:2006:DTH

[Lob06b]  S. N. Lobanov.  Distribu-

tion of trend height and length for Brownian motion with a drift. Theory of Probability and its Applications, 50(4):694– 700, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Louhichi:2002:RCC

[Lou02] S. Louhichi. Rates of convergence in the CLT for some weakly dependent random variables. Theory of Probability and its Applications, 46(2): 297–315, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97894.

Louhichi:2003:MIS

[Lou03] S. Louhichi. Moment inequalities for sums of certain dependent random variables. Theory of Probability and its Applications, 47(4):649–664, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98004.

Lipster:2000:DAO

[LRT00] R. Lipster, W. J. Runggaldier, and M. Taksar. Diffusion approximation and optimal stochastic control. Theory of Probability and its Applications, 44(4):669–698, December 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97786.

Lifshits:2002:PMQ

[LS02] M. F. Lifshits and E. Yu. Shmileva. Poisson measures quasiinvariant with respect to multiplicative transformations. Theory of Probability and its Applications, 46(4):652–666, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97927.

Lo:2003:CBP

[LS03] A. Y. Lo and V. V. Sazonov.

On consistency of Bayes procedures. Theory of Probability and its Applications, 47(2): 236–256, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97961.

Lifshits:2004:LFE

[LS04] M. A. Lifshits and Z. Shi. Lower functions of empirical processes and Brownian sheets. Theory of Probability and its Applications, 48(2):273–287, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98039.

Luo:2009:CQF

[Luo09] S. Luo. On covariance and quantum Fisher information. Theory of Probability and its Applications, 53(2):329–334, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Lazakovich:2006:LBI

[LY06] N. V. Lazakovich and A. L. Yablonski. The limiting behavior of Itô’s finite sums with averaging. Theory of Probability and its Applications, 50(4): 612–630, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ma:2009:ISV

[Ma09] C. Ma. Intrinsically stationary variograms in space and time. Theory of Probability and its Applications, 53(1):145–155, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Macutek:2006:LPG

[Mac06] J. Macutek. A limit property of the geometric distribution. Theory of Probability and its Applications, 50(2):316–319, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Magnus:2000:TPE

[Mag00] J. R. Magnus. The traditional pretest estimator. Theory of Probability and its Applications, 44(2):293–308, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97760.

Makhno:2000:CSO

[Mak00] S. Ya. Makhno. Convergence of solutions of one-dimensional stochastic equations. Theory of Probability and its Applications, 44(3):495–510, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97773.

Makrushin:2003:UME

[Mak03] A. V. Makrushin. he unimprovability of moment estimates. Theory of Probability and its Applications, 47(1):164– 171, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97953.

Makhno:2004:LTO

[Mak04] S. Ya. Makhno. Limit theorem for one-dimensional stochastic equations. Theory of Probability and its Applications, 48(1): 164–169, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98019.

Malyshev:2001:MIL

[Mal01a] V. A. Malyshev. Macrodimension: An invariant of local dynamics. Theory of Probability and its Applications, 45(2): 323–329, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97826.

Malyshkin:2001:SER

[Mal01b] M. N. Malyshkin. Subexponential estimates of the rate of convergence to the invariant measure for stochastic differential equations. Theory of Probability and its Applications, 45(3):466– 479, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97840.

Manita:2009:SSL

[Man09] A. D. Manita. Stochastic synchronization in a large system of identical particles. Theory of Probability and its Applications, 53(1):155–161, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Maslov:2000:EKF

[Mas00]  V. P. Maslov.  An equa-

tion of the Kolmogorov–Feller type in a factorized probability space. Theory of Probability and its Applications, 44(2):

309–318, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97758.

Maslov:2001:QER

[Mas01] V. P. Maslov. Quantized entropy and its relation to occupation numbers. Theory of Probability and its Applications, 45(4):678–680, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97860.

Maslov:2003:QSM

[Mas03] V. P. Maslov. Quantum statistics methods from the viewpoint of probability theory. I. Theory of Probability and its Applications, 47(4):665–683, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98001.

Maslov:2004:IEP

[Mas04a] V. P. Maslov. Integral equations and phase transitions in stochastic games. an analogy with statistical physics. Theory of Probability and its Applications, 48(2):359–367, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98046.

Maslov:2004:NAA

[Mas04b] V. P. Maslov. Nonlinear averaging axioms in financial mathematics and stock price dynamics. Theory of Probability and its Applications, 48(4):723– 733, December 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98077.

Maslov:2005:NFA

[Mas05] V. P. Maslov. Nonlinear financial averaging, the evolution process, and laws of econophysics. Theory of Probability and its Applications, 49(2): 221–244, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98101.

Maslov:2007:TGD

[Mas07] V. P. Maslov. Thermo- and gasdynamical processes in NPPs after accidents. Theory of Probability and its Applications, 51 (3):513–517, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Matsak:2000:LIL

[Mat00] I. K. Matsak. On the law of the iterated logarithm in Banach lattices. Theory of Probability and its Applications, 44(4):775– 784, December 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97796.

Mattingly:2003:CER

[Mat03]  J. C. Mattingly.  Contrac-

tivity and ergodicity of the random map x ↦ |x − θ|. Theory of Probability and its Applications, 47(2):333– 343, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97976.

Matiyasevich:2004:OPE

[Mat04] Yu. V. Matiyasevich. One probabilistic equivalent of the four color conjecture. Theory of Probability and its Applications, 48(2):368–372, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98047.

Mattner:2009:LBT

[Mat09] L. Mattner. Lower bounds for tails of sums of independent symmetric random variables. Theory of Probability and its Applications, 53(2):334–339, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mauch:2004:EPD

[Mau04] F. Mauch. On the existence of probability distributions with given marginals. Theory of Probability and its Applications, 48(3):541–548, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98063.

Myronyuk:2008:ILS

[MF08] M. V. Myronyuk and G. M. Feldman. Independent linear statistics on the twodimensional torus. Theory of Probability and its Applications, 52(1):78–92, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikhailov:2002:EAC

[Mik02] V. G. Mikhailov. Estimate of the accuracy of the compound Poisson approximation for the distribution of the number of matching patterns. Theory of Probability and its Applications, 46(4):667–675, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97928.

Mikhailov:2003:PTL

[Mik03] V. G. Mikhailov. Poisson-type limit theorems for the number of incomplete matches of Spatterns. Theory of Probability and its Applications, 47(2): 343–351, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97970.

Mikhailov:2009:PTL

[Mik09] V. G. Mikhailov. A Poissontype limit theorem for the number of pairs of matching sequences. Theory of Probability and its Applications, 53(1): 106–116, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Minozzo:2000:PGT

[Min00] M. Minozzo. Purely gametheoretic random sequences: I. Strong law of large numbers and law of the iterated logarithm. Theory of Probability and its Applications, 44(3):511– 522, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97776.

Minozzo:2001:PGT

[Min01] M. Minozzo. Purely gametheoretic random sequences: II. Limiting empirical distributions and strong central limit theorem. Theory of Probability and its Applications, 45(2): 233–245, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97819.

Mishchenko:2006:DBP

[Mis06a] A. S. Mishchenko. A discrete Bessel process and its properties. Theory of Probability and its Applications, 50(4):700– 709, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mishchenko:2006:DSF

[Mis06b] A. S. Mishchenko. On a distribution of some functionals of a simple random walk. Theory of Probability and its Applications, 50(4):710–717, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mitrophanov:2006:SEF

[Mit06] A. Yu. Mitrophanov. Stability estimates for finite homogeneous continuous-time Markov chains. Theory of Probability and its Applications, 50(2): 319–326, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mladenovic:2006:GMH

[Mla06] P. Mladenovic. A generalization of the Mejzler–de Haan Theorem. Theory of Probability and its Applications, 50(1): 141–153, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malyshev:2006:PTT

[MM06] V. Malyshev and A. Manita. Phase transitions in the time synchronization model. Theory of Probability and its Applications, 50(1):134–141, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Malyshev:2009:SMC

[MM09] V. A. Malyshev and A. D.

Manita. Stochastic micromodel of the Couette flow. Theory of Probability and its Applications, 53(4):716–727, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Makagon:2008:SDP

[MMSS08] A. Makagon, A. G. Miamee, H. Salehi, and A. R. Soltani. On spectral domain of periodically correlated processes. Theory of Probability and its Applications, 52(2):353–361, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Milstein:2003:SRW

[MmVT03] G. N. Milstein and m. V. Tretyakov.

The simplest random walks for the Dirichlet problem. Theory of Probability and its Applications, 47(1):53–68, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97943.

Malyshev:2007:HCR

[MMZ07] V. A. Malyshev, A. D. Manita, and A. A. Zamyatin. Homeostasis in chemical reaction pathways. Theory of Probability and its Applications, 51(4):714–723, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mishura:2001:COS

[MO01] Yu. S. Mishura and Ya. A. Oltsik. Choosing an optimal switching moment on the financial market with alternative strategies (semimartingale approach). Theory of Probability and its Applications, 45(3):480– 493, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97841.

Molchan:2000:MFB

[Mol00] G. M. Molchan. On the maximum of a fractional Brownian motion. Theory of Probability and its Applications, 44(1): 97–102, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97737.

Molchan:2001:MSL

[Mol01] G. M. Molchan. On a maximum of stable Lévy processes. Theory of Probability and its Applications, 45(2):343–349, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97829.

Molchan:2003:LPF

[Mol03] G. M. Molchan. Linear problems for a fractional Brownian motion: Group approach. Theory of Probability and its Applications, 47(1):69–78, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97944.

Mordecki:2000:NCS

[Mor00] E. Mordecki. Necessary conditions for stable Convergence of semimartingales. Theory of Probability and its Applications, 44(1):217–221, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97750.

Mordecki:2004:RPL

[Mor04]  E. Mordecki.  Ruin prob-

abilities for Levy processes with mixed-exponential negative jumps. Theory of Probability and its Applications, 48(1): 170–176, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98017.

Mancino:2001:SRS

[MP01] M. E. Mancino and L. Pratelli.

Some results of stable convergence for exchangeable random variables in Hilbert spaces. Theory of Probability and its Applications, 45(2):329–337, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97827.

Mazalov:2004:APO

[MP04] V. V. Mazalov and N. V. Peshkov. On asymptotic properties of optimal stopping time. Theory of Probability and its Applications, 48(3):549– 555, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98058.

Mokkadem:2008:CLI

[MP08] A. Mokkadem and M. Pelletier. Compact law of the iterated logarithm for matrix-normalized sums of random vectors. Theory of Probability and its Applications, 52(4):636–650, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mokkadem:2006:LDU

[MPW06] A. Mokkadem, M. Pelletier, and J. Worms. A large deviations upper bound for the kernel mode estimator. Theory of Probability and its Applications, 50(1):153–165, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mikulevicius:2000:FHE

[MR00] R. Mikulevicius and B. Rozovskii. Fourier–Hermite expansions for nonlinear filtering. Theory of Probability and its Applications, 44(3):606– 612, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97783.

Mirzaev:2007:LTM

[MS07a] M. Mirzaev and A. N. Startsev. Limit theorems for a model of interaction of particles of two types generalizing the Bartlett–McKendrick epidemic process. Theory of Probability and its Applications, 51 (2):362–367, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mishura:2007:ASS

[MS07b] Yu. S. Mishura and G. M.

Shevchenko. Approximation schemes for stochastic differential equations in Hilbert space. Theory of Probability and its Applications, 51(3):442–458, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Mania:2001:SFC

[MT01] M. Mania and R. Tevzadae.

Semimartingale functions of a class of diffusion processes. Theory of Probability and its Applications, 45(2):337–343, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97828.

Mason:2002:SPP

[MY02] J. D. Mason and Xiao Yimin.

Sample path properties of operator-self-similar Gaussian random fields. Theory of Probability and its Applications, 46 (1):58–78, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97874.

Nadtochiy:2007:ABR

[Nad07] S. A. Nadtochiy. Asymptotic behavior of a random walk with interaction. Theory of Probability and its Applications, 51 (1):182–188, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:2001:PMI

[Nag01]  S. V. Nagaev.  On prob-

ability and moment inequalities for dependent random variables. Theory of Probability and its Applications, 45(1):152– 160, March 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97814.

Nagaev:2002:LBP

[Nag02a] S. V. Nagaev. Lower bounds for probabilities of large deviations of sums of independent random variables. Theory of Probability and its Applications, 46(4):728– 735, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97933.

Nagaev:2002:LBL

[Nag02b] S. V. Nagaev. Lower bounds on large deviation probabilities for sums of independent random variables. Theory of Probability and its Applications, 46(1): 79–102, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97872.

Nagaev:2005:LDS

[Nag05] S. V. Nagaev. On large deviations of a self-normalized sum. Theory of Probability and its Applications, 49(4):704– 713, December 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98139.

Nagaev:2007:APM

[Nag07a] A. V. Nagaev. Asymptotic properties of multidimensional stable densities and asymmetric problems of large deviations. Theory of Probability and its Applications, 51(4):626–644, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:2007:PMI

[Nag07b] S. V. Nagaev. On probability and moment inequalities for supermartingales and martingales. Theory of Probability and its Applications, 51(2):367–377, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nasyrov:2000:RCC

[Nas00] F. S. Nasyrov. On the representation of certain classes of stochastic Ito integrals in the form of pathwise Lebesgue integrals. Theory of Probability and its Applications, 44(2): 404–409, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97765.

Nasyrov:2007:SIS

[Nas07] F. S. Nasyrov. Symmetric integrals and stochastic analysis. Theory of Probability and its Applications, 51(3):486–503, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nechaev:2000:MVH

[Nec00] M. L. Nechaev. On the meanvariance hedging in the Ho– Lee diffusion model. Theory of Probability and its Applications, 44(1):102–106, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97738.

Neuenschwander:2000:CRI

[Neu00] D. Neuenschwander. The class I_{0} for random increasing upper semicontinuous functions. Theory of Probability and its Applications, 44(1):106–109, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97745.

Nirenberg:2003:MNA

[Nir03] Louis Nirenberg. In memoriam:

Natascha Artin Brunswick, Editor Emeritus. Theory of Probability and its Applications, 47 (2):189, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97997.

Nadarajah:2009:CPD

[NK09] S. Nadarajah and S. Kotz. A class of t posterior distributions with application. Theory of Probability and its Applications, 53(1):161–169, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nazarov:2005:LSB

[NN05] A. I. Nazarov and Ya. Yu.

Nikitin. Logarithmic L_{2}-small ball asymptotics for some fractional Gaussian processes. Theory of Probability and its Applications, 49(4):645–658, December 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98131.

Noquet:2000:ITV

[Noq00] C. Noquet. Inequalities for the total variation between the distributions of a sequence and its translate. Theory of Probability and its Applications, 44(3):561– 569, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97781.

Nosko:2002:HRC

[Nos02] V. P. Nosko. The horizon of a random cone field under a trend: One-dimensional distributions. Theory of Probability and its Applications, 46(4):736– 744, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97934.

Novak:2000:GKD

[Nov00a] S. Y. Novak. Generalized kernel density estimator. Theory of Probability and its Applications, 44(3):570–583, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97778.

Novak:2000:MUP

[Nov00b] S. Yu. Novak. On the mode of an unknown probability distribution. Theory of Probability and its Applications, 44(1): 109–113, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97739.

Novikov:2004:MFP

[Nov04] A. Novikov. Martingales and first-passage times for Ornstein– Uhlenbeck processes with a jump component. Theory of Probability and its Applications, 48(2):288–303, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98040.

Novak:2005:SNS

[Nov05] S. Y. Novak. On self-normalized sums and Student’s statistic. Theory of Probability and its Applications, 49(2):336–344, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98108.

Novikov:2009:DFP

[Nov09] A. A. Novikov. On distributions of first passage times and optimal stopping of AR(1) sequences. Theory of Probability and its Applications, 53(3): 419–429, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nowak:2005:ACB

[Now05] E. Nowak. Absolute continuity between a Gibbs measure and its translate. Theory of Probability and its Applications, 49(4):713–724, December 2005. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98142.

Novikov:2005:ESO

[NS05] A. A. Novikov and A. N. Shiryaev. On an effective solution of the optimal stopping problem for random walks. Theory of Probability and its Applications, 49(2):344–354, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98109.

Nagaev:2006:PIC

[NV06a] S. V. Nagaev and V. Vakhtel. Probability inequalities for a critical Galton–Watson process. Theory of Probability and its Applications, 50(2):225– 247, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nagaev:2006:LLT

[NV06b] S. V. Nagaev and V. I. Vakhtel. On the local limit theorem for a critical Galton–Watson process. Theory of Probability and its Applications, 50(3):400– 419, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Nazarov:2009:WLE

[NZ09] A. Nazarov and D. Zaporozhets. What is the least expected number of real roots of a random polynomial? Theory of Probability and its Applications, 53 (1):117–133, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ouknine:2009:TTI

[OE09] Y. Ouknine and M. Erraoui.

Transformations of two independent Brownian motions and orthogonal decompositions of Brownian filtrations. Theory of Probability and its Applications, 53(4):610–625, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Oseledets:2002:STN

[OK02] V. I. Oseledets and D. V.

Khmelev. Stochastic transportation networks and stability of dynamical systems. Theory of Probability and its Applications, 46(1):154–161, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97878.

Orlov:2009:TER

[Orl09] D. Orlov. On two estimates of a risk measure. Theory of Probability and its Applications, 53 (1):169–173, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Osmolovskii:2009:SCM

[Osm09] I. Yu. Osmolovskii. On some characteristics of the multivariate analogues of the Chebyshev– Hermite polynomials. Theory of Probability and its Applications, 53(2):340–346, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ostrovskii:2002:AMC

[Ost02] E. I. Ostrovskii. The anisotropic moduli of continuity of random fields. Theory of Probability and its Applications, 46(1): 161–167, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97879.

Panfilova:2006:NBa

[Pan06a] E. V. Panfilova. New books. Theory of Probability and its Applications, 50(2):351– 352, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Panfilova:2006:NBb

[Pan06b] E. V. Panfilova. New books. Theory of Probability and its Applications, 50(4):723, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pastukhov:2005:SPS

[Pas05] S. V. Pastukhov. On some probabilistic-statistical methods in technical analysis. Theory of Probability and its Applications, 49(2):245–260, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98102.

Porcu:2009:SOP

[PCMGM09] E. Porcu, R. Crujeiras, J. Mateu, and W. Gonzalez-Manteiga. On the second order properties of the multidimensional periodogram for regularly spaced data. Theory of Probability and its Applications, 53(2):349–356, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pechinkin:2001:QSS

[Pec01] A. V. Pechinkin. The MAP/G/1/∞ queue with SRPT service discipline. Theory of Probability and its Applications, 45(3):532– 539, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97849.

Peskir:2009:OSG

[Pes09]  G. Peskir.  Optimal stop-

ping games and Nash equilibrium. Theory of Probability and its Applications, 53(3):558–571, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Petrov:2002:LIL

[Pet02] V. V. Petrov. On the law of the iterated logarithm for a sequence of independent random variables. Theory of Probability and its Applications, 46(3):542– 544, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97915.

Petrov:2003:RCF

[Pet03] A. I. Petrov. Random contextfree grammars: Supercritical case with a nonzero extinction probability. Theory of Probability and its Applications, 47(4):709–718, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98011.

Petrov:2009:SLL

[Pet09] V. V. Petrov. On the strong law of large numbers for nonnegative random variables. Theory of Probability and its Applications, 53(2):346–349, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pingyan:2006:CTL

[Pin06] C. Pingyan. The Chover-type law of the iterated logarithm for certain power series. Theory of Probability and its Applications, 50(3):497–505, January 2006. CODEN TPRBAU.

ISSN 0040-585X (print), 1095-7219 (electronic).

Pinelis:2009:LE

[Pin09] I. Pinelis. Letter to the Editors. Theory of Probability and its Applications, 53(2):383, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piterbarg:2003:LJC

[PK03] V. I. Piterbarg and A. M. Kozlov. On large jumps of a Cramer random walk. Theory of Probability and its Applications, 47(4):719–729, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98012.

Prokhorov:2008:PCG

[PK08] Yu. V. Prokhorov and V. I.

Khokhlov. On polynomials of components of Gaussian random vectors. Theory of Probability and its Applications, 52 (4):720–724, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pancheva:2007:RTC

[PKJ07] E. I. Pancheva, E. T. Kolkovska, and P. K. Jordanova. Random time-changed extremal processes. Theory of Probability and its Applications, 51 (4):645–662, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pang:2007:NCT

[PL07] T.-X. Pang and Z.-Y. Lin. A nonclassical Chung-type law of the iterated logarithm for independent identically distributed random variables. Theory of Probability and its Applications, 51(4):723–729, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Peng:2003:JLD

[PN03] Z. Peng and S. Nadarajah. On the joint limiting distribution of sums and maxima of stationary normal sequence. Theory of Probability and its Applications, 47(4):706–709, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98014.

Presman:2009:ECB

[Pre09] E. L. Presman. Estimation of the constant in a Burkholder inequality for supermartingales and martingales. Theory of Probability and its Applications, 53(1):173–179, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Peskir:2002:NCP

[PS02] G. Peskir and A. N. Shiryaev. A note on the call-put parity and a call-put duality. Theory of Probability and its Applications, 46(1):167–170, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97884.

Piterbarg:2004:LTH

[PS04] V. Piterbarg and S. Stamatovic. Limit theorem for high level a-upcrossings by χprocess. Theory of Probability and its Applications, 48(4):734– 741, December 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98078.

Paulauskas:2008:RAL

[PS08] V. Paulauskas and D. Surgailis.

On the rate of approximation in limit theorems for sums of moving averages. Theory of Probability and its Applications, 52 (2):361–370, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:2009:SFY

[PS09] Yu. V. Prokhorov and A. N.

Shiryaev. Seventy-five years since publication of the monograph by A. N. Kolmogorov. Theory of Probability and its Applications, 53(2):191–193, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pellegrinotti:2000:SSD

[PSV00] A. Pellegrinotti, V. Sidoravicius, and M. E. Vares. Stationary state and diffusion for a charged particle in a onedimensional medium with lifetimes. Theory of Probability and its Applications, 44(4):697– 721, December 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97790.

Philippe:2008:TVF

[PSV08] A. Philippe, D. Surgailis, and M.-C. Viano. Time-varying fractionally integrated processes with nonstationary long memory. Theory of Probability and its Applications, 52(4):651–673, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Piterbarg:2001:MRC

[PT01] V. I. Piterbarg and Yu. N. Tyurin. Multivariate rank correlations: a Gaussian field on a direct product of spheres. Theory of Probability and its Applications, 45(2):246–257, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97820.

Pollak:2009:AED

[PT09] M. Pollak and A. G. Tartakovsky. Asymptotic exponentiality of the distribution of first exit times for a class of Markov processes with applications to quickest change detection. Theory of Probability and its Applications, 53(3):430–442, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Prokhorov:2002:PRS

[PU02] A. V. Prokhorov and N. G.

Ushakov. On the problem of reconstructing a summands distribution by the distribution of their sum. Theory of Probability and its Applications, 46(3):420– 430, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97920.

Pugachev:2009:CSM

[Pug09a] O. V. Pugachev. Capacities and surface measures in locally convex spaces. Theory of Probability and its Applications, 53 (1):179–190, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pugachev:2009:MCD

[Pug09b] O. V. Pugachev. On Mosco convergence of diffusion Dirichlet forms. Theory of Probability and its Applications, 53(2):242–255, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Pustylnikov:2003:DWE

[Pus03] L. D. Pustylńikov. Discrete wave equations with random parameters and a discrete string with random masses. Theory of Probability and its Applications, 47(2):257–268, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97962.

Prokhorov:2002:BAC

[PVK02] Yu. V. Prokhorov, O. V. Viskov, and V. I. Khokhlov. Binomial analogues of the Chernoff inequality. Theory of Probability and its Applications, 46(3):544– 547, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97907.

Prokhorov:2006:ACI

[PVK06] Yu. V. Prokhorov, O. V. Viskov, and V. I. Khokhlov. Analogues of the Chernoff inequality for negative binomial distribution. Theory of Probability and its Applications, 50(2): 327–330, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Radchenko:2007:VMH

[Rad07] V. M. Radchenko. Varianceminimizing hedging in a model with jumps at deterministic times. Theory of Probability and its Applications, 51(3):536–545, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ragimov:2006:ILT

[Rag06a] F. G. Ragimov. Integral limit theorems for nonlinear boundary crossing time by sums of independent variables. Theory of Probability and its Applications, 50(1):165–168, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ragimov:2006:ABO

[Rag06b] F. G. Ragimov. On the asymptotical behavior of one class of the first exit time of a multidimensional random walk. Theory of Probability and its Applications, 50(3):505–510, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rogozin:2000:CDS

[Rog00] B. A. Rogozin. On the constant in the definition of subexponential distributions. Theory of Probability and its Applications, 44(2):409–412, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97766.

Rogozin:2002:RLT

[Rog02] B. A. Rogozin. Ratio limit theorem for concentration functions. Theory of Probability and its Applications, 46(4):744– 746, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97935.

Rogozin:2003:TTD

[Rog03] B. A. Rogozin. Tauberian theorems for decreasing functions of dominated variation. Theory of Probability and its Applications, 47(2):351–357, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97971.

Rogozin:2006:CCF

[Rog06] B. A. Rogozin. Compactness and the concentration functions of the convolutions of a distribution. Theory of Probability and its Applications, 50(1): 99–111, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rokhlin:2005:EVD

[Rok05] D. B. Rokhlin. An extended version of the Dalang–Morton– Willinger theorem under portfolio constraints. Theory of Probability and its Applications, 49(3):429–443, September 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98118.

Rokhlin:2006:MSP

[Rok06] D. B. Rokhlin. A martingale selection problem in the finite discrete-time case. Theory of Probability and its Applications, 50(3):420–435, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rokhlin:2008:CNA

[Rok08] D. B. Rokhlin. Constructive noarbitrage criterion under transaction costs in the case of finite discrete time. Theory of Probability and its Applications, 52 (1):93–107, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rokhlin:2009:ESD

[Rok09] D. B. Rokhlin. Equivalent supermartingale densities and measures in discrete time infinite horizon market models. Theory of Probability and its Applications, 53(4):626–647, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Romanova:2001:NAP

[Rom01] T. Romanova. Numerical analysis of Piterbarg–Tyurin procedures for testing homogeneity and independence. Theory of Probability and its Applications, 45(4):681–687, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97861.

Roos:2001:BAP

[Roo01] B. Roos. Binomial approximation to the Poisson binomial distribution: The Krawtchouk expansion. Theory of Probability and its Applications, 45(2): 258–272, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97821.

Roos:2002:MKA

[Roo02] B. Roos. Multinomial and Krawtchouk approximations to the generalized multinomial distribution. Theory of Probability and its Applications, 46(1): 103–117, March 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97875.

Roos:2003:KMM

[Roo03] B. Roos. Kerstan’s method in the multivariate Poisson approximation: An expansion in the exponent. Theory of Probability and its Applications, 47(2):358–363, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97977.

Roos:2004:PAC

[Roo04] B. Roos. Poisson approximation via the convolution with Kornya–Presman signed measures. Theory of Probability and its Applications, 48(3):555– 560, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98064.

Rotar:2008:EED

[Rot08]  V. Rotar.  On Edgeworth

expansions for dependencyneighborhoods chain structures with strong mixing characteristics. Theory of Probability and its Applications, 52(1):108–124, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovsky:2000:EPU

[Roz00a] L. V. Rozovsky. An extremal property of the uniform distribution and some of its consequences. Theory of Probability and its Applications, 44(3):583– 588, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97779.

Rozovsky:2000:FT

[Roz00b] L. V. Rozovsky. On a Feller theorem. Theory of Probability and its Applications, 44(1): 119–124, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97741.

Rozovsky:2000:JDI

[Roz00c]  L. V. Rozovsky.  On the

J. Doob inequality for characteristic functions. Theory of Probability and its Applications, 44(3):588–590, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97780.

Rozanov:2002:MEL

[Roz02] Yu. A. Rozanov. On Monotone extension of linear continuous functionals. Theory of Probability and its Applications, 46(4):747–749, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97938.

Rozovsky:2003:EBR

[Roz03a] L. V. Rozovsky. An estimate from below of the remainder in the Central Limit Theorem for a sum of independent random variables with finite moments of a high order. Theory of Probability and its Applications, 47(1): 174–183, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97955.

Rozovsky:2003:RKL

[Roz03b] L. V. Rozovsky. A remark on the Kolmogorov law of the iterated logarithm. Theory of Probability and its Applications, 47(1):

171–173, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97954.

Rozovsky:2004:EAW

[Roz04a]  L. V. Rozovsky.  On ex-

act asymptotics in the weak law of large numbers for sums of independent random variables with a common distribution function from the domain of attraction of a stable law. Theory of Probability and its Applications, 48(3):561– 568, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98059.

Rozovsky:2004:SDS

[Roz04b] L. V. Rozovsky. Superlarge deviations of a sum of independent random variables having a common absolutely continuous distribution under the Cramer condition. Theory of Probability and its Applications, 48(1): 108–130, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98023.

Rozovsky:2005:EAW

[Roz05] L. V. Rozovsky. On exact asymptotics in the Weak Law of Large Numbers for sums of independent random variables with a common distribution function from the domain of attraction of a stable law. II. Theory of Probability and its Applications, 49(4):724– 734, December 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98140.

Rozovsky:2008:SDP

[Roz08] L. V. Rozovsky. Superlarge deviation probabilities for sums of independent random variables with exponential decreasing distribution. Theory of Probability and its Applications, 52 (1):167–171, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovsky:2009:CTL

[Roz09a] L. V. Rozovsky. The Chernofftype large deviations for sums with a random normalization. Theory of Probability and its Applications, 53(4):727–735, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rozovsky:2009:GMB

[Roz09b] L. V. Rozovsky. On Gaussian measure of balls in a Hilbert space. Theory of Probability and its Applications, 53(2):357–364, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Rinott:2000:SBR

[RR00] Y. Rinott and V. Rotar. Some bounds on the rate of convergence in the CLT for martingales. II. Theory of Probability and its Applications, 44(3):523– 536, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97774.

Ramachandran:2001:MKD

[RR01] D. Ramachandran and L. Rüschendorf.

On the Monge-kantorovitch duality theorem. Theory of Probability and its Applications, 45(2):350–356, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97830.

Rudzkis:2005:AED

[RR05] R. Rudzkis and M. Radavicius.

Adaptive estimation of distribution density in the basis of algebraic polynomials. Theory of Probability and its Applications, 49(1):93–109, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98089.

Rozanov:2000:SBP

[RS00] Yu. A. Rozanov and F. Sanso.

On stochastic boundary problems for harmonic functions in Banach spaces. Theory of Probability and its Applications, 44(1):114–119, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97740.

Regazzini:2001:ALR

[RS01] E. Regazzini and V. V. Sazonov. Approximation of laws of random probabilities by mixtures of Dirichlet distributions with applications to nonparametric Bayesian inference. Theory of Probability and its Applications, 45(1):93–110, March 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97806.

Rackauskas:2005:CLT

[RS05] A. Rackauskas and C. Suquet. Central Limit Theorems in Hölder topologies for Banach space valued random fields. Theory of Probability and its Applications, 49(1): 77–92, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98088.

Rukhin:2001:TRS

[Ruk01] A. L. Rukhin. Testing randomness: a suite of statistical procedures. Theory of Probability and its Applications, 45(1): 111–132, March 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97808.

Rukhin:2007:PCM

[Ruk07] A. L. Rukhin. Pattern correlation matrices for Markov sequences and tests of randomness. Theory of Probability and its Applications, 51(4):663–679, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Ruzankin:2002:COJ

[Ruz02] P. S. Ruzankin. Construction of the optimal joint distribution of two random variables. Theory of Probability and its Applications, 46(2): 316–334, June 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97890.

Ryabinin:2001:CFP

[Rya01] A. A. Ryabinin. On characteristic functions of probability distributions of sums with random permutations of signs. Theory of Probability and its Applications, 45(4):687– 690, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97863.

Rybakov:2001:AOB

[Ryb01a] A. S. Rybakov. On the asymptotic optimality of the Bayesian decision rule in the problem of multiple classification of hypotheses. Theory of Probability and its Applications, 45(4):690– 695, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97862.

Rybunikova:2001:LRF

[Ryb01b] T. S. Rybunikova. On linear row-finite systems of stochastic differential equations. Theory of Probability and its Applications, 45(3):539–545, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97848.

Rychlik:2003:SMV

[Ryc03]  T. Rychlik.  Sharp mean-

variance inequalities for quantiles of distributions determined by convex and star orders. Theory of Probability and its Applications, 47(2):269–285, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97966.

Sakhanenko:2005:TPR

[Sak05] A. I. Sakhanenko. On transient phenomena in random walks. Theory of Probability and its Applications, 49(2): 354–367, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98110.

Salikhov:2003:OST

[Sal03] N. P. Salikhov. Optimal sequences of tests for several polynomial schemes of trials. Theory of Probability and its Applications, 47(2):286–298, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97963.

Shiryaev:2000:SDP

[SC00] A. N. Shiryaev and A. S.

Cherny. Some distributional properties of a Brownian motion with a drift and an extension of P. Lévy’s theorem. Theory of Probability and its Applications, 44(2):412–418, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97768.

Schreiber:2003:LTC

[Sch03]  T. Schreiber.  Limit theo-

rems for certain functionals of unions of random closed sets. Theory of Probability and its Applications, 47(1):79– 90, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97949.

Selivanov:2005:MME

[Sel05]  A. V. Selivanov.  On the

martingale measures in exponential Lévy models. Theory of Probability and its Applications, 49(2):261–274, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98103.

Senatov:2007:AEC

[Sen07] V. V. Senatov. On asymptotic expansions in the Central Limit Theorem with explicit estimates of remainder terms. Theory of Probability and its Applications, 51(4):729–736, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:2008:OPA

[Sen08] V. V. Senatov. On a one polynomial analogue of the Chebyshev expansion. Theory of Probability and its Applications, 52 (3):531–538, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Senatov:2009:SAD

[Sen09] V. V. Senatov. Some asymptotic decompositions in the Central Limit Theorem in the multidimensional case. Theory of Probability and its Applications, 53(2):256–268, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Serdobolskii:2000:ADE

[Ser00] V. I. Serdobolskii. Asymptotically dominating estimation of expectation value vectors. Theory of Probability and its Applications, 44(1):124– 130, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97742.

Sergeev:2006:LTO

[Ser06] A. A. Sergeev. Limit theorems for one class of polling models. Theory of Probability and its Applications, 50(3): 510–518, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sharipov:2003:IPS

[Sha03] O. Sh. Sharipov. Invariance principle for U-statistics and von Mises’ functionals of weakly dependent observations. Theory of Probability and its Applications, 47(4):730–733, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98013.

Sharipov:2005:BLI

[Sha05] O. Sh. Sharipov. The bounded law of the iterated logarithm for weakly dependent random variables with values in a type 2 Banach space. Theory of Probability and its Applications, 49(3):444–457, September 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98119.

Shashkin:2006:NCL

[Sha06a] A. P. Shashkin. On Newman’s Central Limit Theorem. Theory of Probability and its Applications, 50(2):330–337, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shatskikh:2006:SLL

[Sha06b] S. Ya. Shatskikh. The strong law of large numbers for a triangular array scheme of conditional distributions of stable elliptically contoured measures. Theory of Probability and its Applications, 50(2):248–264, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shervashidze:2000:LLT

[She00] T. Shervashidze. Local limit theorems for conditionally independent random variables controlled by a finite Markov chain. Theory of Probability and its Applications, 44(1):131– 135, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97744.

Sherstnev:2003:DUD

[She03] V. I. Sherstnev. Decompositions of the uniform distribution on a finite group. Theory of Probability and its Applications, 47(3):550–555, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97993.

Shevtsova:2007:SUB

[She07] I. G. Shevtsova. Sharpening of the upper bound of the absolute constant in the Berry– Esseen inequality. Theory of Probability and its Applications, 51(3):549–553, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2000:HFR

[Shi00] A. N. Shiryaev. On the history of the foundation of the Russian Academy of Sciences and about the first articles on probability theory in Russian publications. Theory of Probability and its Applications, 44(2): 225–230, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97752.

Shipileva:2001:EDE

[Shi01] A. V. Shipileva. Estimates of the distribution of the extinction moment of a Markov branching process. Theory of Probability and its Applications, 45(4):695– 698, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97864.

Shixin:2002:SLL

[Shi02] Gan Shixin. Strong laws of large numbers for B-valued L_{q}-mixingale sequences and the q-smoothness of Banach space. Theory of Probability and its Applications, 46(4):717– 721, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97937.

Shirokov:2008:PQC

[Shi08] M. E. Shirokov. On properties of quantum channels related to their classical capacity. Theory of Probability and its Applications, 52(2):250–276, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shirokov:2009:CFH

[Shi09a] M. E. Shirokov. On channels with finite Holevo capacity. Theory of Probability and its Applications, 53(4):648–662, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2009:NSL

[Shi09b] A. N. Shiryaev. News of scientific life. Theory of Probability and its Applications, 53 (2):380–382, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2009:CEP

[Shi09c] A. N. Shiryaev. On conditionalextremal problems of the quickest detection of nonpredictable times of the observable Brownian motion. Theory of Probability and its Applications, 53 (4):663–678, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2009:SMO

[Shi09d] A. N. Shiryaev. On stochastic models and optimal methods in the quickest detection problems. Theory of Probability and its Applications, 53(3):385–401, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sholomitskii:2005:NCP

[Sho05] A. G. Sholomitskii. On the necessary conditions of Poisson convergence for martingales. Theory of Probability and its Applications, 49(4):735– 737, December 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98141.

Shur:2001:RLT

[Shu01]  M. G. Shur.  Ratio limit

theorems for self-adjoint operators and symmetric Markov chains. Theory of Probability and its Applications, 45(2): 273–288, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97822.

Shur:2006:UIS

[Shu06] M. G. Shur. Uniform integrability for strong ratio limit theorems. I. Theory of Probability and its Applications, 50(3): 436–447, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2008:NSL

[SM08] A. N. Shiryaev and V. V. Mazalov. News of scientific life. Theory of Probability and its Applications, 52(1):178–179, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Smorodina:2000:AED

[Smo00] N. V. Smorodina. Asymptotic expansion of the distribution of a homogeneous functional of a strictly stable random vector. II. Theory of Probability and its Applications, 44(2): 419–427, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97767.

Sobolev:2008:AEC

[Sob08] V. N. Sobolev. On asymptotic expansions in the Central Limit Theorem. Theory of Probability and its Applications, 52 (3):456–469, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sofronov:2002:AOT

[Sof02]  G. Yu. Sofronov.  Asymp-

totically d-optimal test of a change-point detection. Theory of Probability and its Applications, 46(3):547–548, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97916.

Sofronov:2005:AOT

[Sof05]  G. Yu. Sofronov.  Asymp-

totically d-optimal test of a posteriori change-point detection. Theory of Probability and its Applications, 49(2):367– 371, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98111.

Soloviov:2000:TTM

[Sol00] V. N. Soloviov. Towards the theory of minimax-Bayesian estimation. Theory of Probability and its Applications, 44(4):739– 754, December 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97787.

Soltani:2006:SRM

[SP06] A. R. Soltani and A. Parvardeh.

Simple random measures and simple processes. Theory of Probability and its Applications, 50(3):448–462, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Schachermayer:2000:TPC

[SS00] W. Schachermayer and W. Schachinger. Is there a predictable criterion for mutual singularity of two probability measures on a filtered space? Theory of Probability and its Applications, 44 (1):51–59, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97736.

Startsev:2002:MIP

[Sta02] A. N. Startsev. On a model of interacting particles of two types generalizing the Bartlett– McKendrick epidemic process. Theory of Probability and its Applications, 46(3):431–447, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97908.

Stepanov:2003:CDT

[Ste03]  A. Stepanov.  Conditional

distributions of the times of one-sided sequential approximations. Theory of Probability and its Applications, 47(2): 363–365, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97972.

Su:2002:CLT

[Su02] Z. G. Su. On Central Limit Theorems for vector random measures and measure-valued processes. Theory of Probability and its Applications, 46(3):448– 468, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97911.

Suidan:2002:CMR

[Sui02] T. M. Suidan. Convex minorants of random walks and Brownian motion. Theory of Probability and its Applications, 46(3):469–481, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97910.

Sunklodas:2008:NAS

[Sun08] J. Sunklodas. On normal approximation for strongly mixing random fields. Theory of Probability and its Applications, 52 (1):125–132, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Shiryaev:2004:PSI

[SY04] A. N. Shiryaev and M. Yor. On the problem of stochastic integral representations of functionals of the Brownian motion. I. Theory of Probability and its Applications, 48(2): 304–313, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98042.

Szewczak:2006:RLD

[Sze06] Z. S. Szewczak. A remark on a large deviation theorem for Markov chain with a finite number of states. Theory of Probability and its Applications, 50 (3):518–528, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tartakovsky:2009:AOB

[Tar09] A. G. Tartakovsky. Asymptotic optimality in Bayesian changepoint detection problems under global false alarm probability constraint. Theory of Probability and its Applications, 53 (3):443–466, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Sinai:2005:BS

[tbYGS05] translated by Ya. G. Sinai. On the 70th birthday of A. N. Shiryaev. Theory of Probability and its Applications, 49(4):691– 694, December 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98135.

Temido:2003:MSL

[TC03] M. G. Temido and L. Canto e Castro. Max-semistable laws in extremes of stationary random sequences. Theory of Probability and its Applications, 47(2): 365–374, June 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97978.

Tempelman:2000:DRF

[Tem00]  A. Tempelman.  Dimension

of random fractals in metric spaces. Theory of Probability and its Applications, 44(3):537– 557, September 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97775.

Editors:2003:IAI

[The03] The Editors. Ildar Abdullovich Ibragimov. Theory of Probability and its Applications, 47(3): 488, September 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97986.

Timashev:2001:LDA

[Tim01] A. N. Timashev. On the largedeviation asymptotics in an allocation scheme of particles into distinguishable cells with restrictions on the size of the cells. Theory of Probability and its Applications, 45(3):494– 506, September 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97842.

Timashev:2003:ILT

[Tim03] A. N. Timashev. Integral limit theorems on large deviations for multidimensional hypergeometric distribution. Theory of Probability and its Applications, 47 (1):91–98, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97945.

Timashev:2004:RMF

[Tim04] A. N. Timashev. Random mappings of finite sets with a known number of components. Theory of Probability and its Applications, 48(4):741– 751, December 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98079.

Timashev:2005:LTA

[Tim05] A. N. Timashev. Limit theorems for allocation of particles over different cells with restrictions to the size of the cells. Theory of Probability and its Applications, 49(4):659– 670, December 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98132.

Timushev:2007:RCP

[Tim07] D. A. Timushev. The rate of convergence in probability for spectra of the GUE. Theory of Probability and its Applications, 51(3):546–549, ???? 2007.

CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Tsilevich:2000:SRP

[Tsi00] N. V. Tsilevich. Stationary random partitions of positive integers. Theory of Probability and its Applications, 44(1): 60–74, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97733.

Tartakovsky:2005:GAB

[TV05] A. G. Tartakovsky and V. V.

Veeravalli. General asymptotic Bayesian theory of quickest change detection. Theory of Probability and its Applications, 49(3):458–497, September 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98120.

Urusov:2005:PMW

[Uru05] M. A. Urusov. On a property of the moment at which Brownian motion attains its maximum and some optimal stopping problems. Theory of Probability and its Applications, 49(1): 169–176, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98095.

Vakulenko:2002:EMS

[Vak02] N. V. Vakulenko. On estimating multimodal spectra of time series. Theory of Probability and its Applications, 46(3):482– 489, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97905.

Vatutin:2003:RBP

[Vat03] V. A. Vatutin. Reduced branching processes in random environment: The critical case. Theory of Probability and its Applications, 47(1):99–113, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97942.

Vatutin:2004:LTI

[Vat04] V. A. Vatutin. Limit theorem for an intermediate subcritical branching process in a random environment. Theory of Probability and its Applications, 48(3):481–492, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/98051.

Volodin:2006:CRD

[VCH06] A. Volodin, M. Ordóñez Cabrera, and T. C. Hu. Convergence rate of the dependent bootstrapped means. Theory of Probability and its Applications, 50(2):337–346, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:2004:GWB

[VD04] V. A. Vatutin and E. E.

Dyakonova. Galton–Watson branching processes in a random environment I: Limit theorems. Theory of Probability and its Applications, 48(2): 314–336, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98037.

Vatutin:2005:GWB

[VD05] V. A. Vatutin and E. E.

Dyakonova. Galton–Watson branching processes in a random environment. II: Finitedimensional distributions. Theory of Probability and its Applications, 49(2):275–309, June 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98100.

Vasudeva:2006:ASB

[VD06] R. Vasudeva and G. Divanji. On almost sure behavior of stable subordinators over rapidly increasing sequences. Theory of Probability and its Applications, 50(4):718–722, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:2007:BPR

[VD07] V. A. Vatutin and E. E. Dyakonova. Branching processes in a random environment and bottlenecks in the evolution of populations. Theory of Probability and its Applications, 51 (1):189–210, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:2008:LTR

[VD08] V. A. Vatutin and E. E. Dyakonova. Limit theorems for reduced branching processes in a random environment. Theory of Probability and its Applications, 52(2):277–302, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:2009:WRB

[VD09] V. A. Vatutin and E. E. Dyakonova. Waves in reduced branching processes in a random environment. Theory of Probability and its Applications, 53 (4):679–695, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Veretennikov:2000:PMC

[Ver00]  A. Yu. Veretennikov.  On

polynomial mixing and convergence rate for stochastic difference and differential equations. Theory of Probability and its Applications, 44(2): 361–374, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97755.

Veretennikov:2001:PMS

[Ver01] A. Yu. Veretennikov. On polynomial mixing for SDEs with a gradient-type drift. Theory of Probability and its Applications, 45(1):160–164, March 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97809.

Veretennikov:2003:SSA

[Ver03] A. Yu. Veretennikov. On SDE and semigroup approximations and large deviations. Theory of Probability and its Applications, 47(4):733–741, December 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98007.

Vershik:2004:KES

[Ver04] A. M. Vershik. Kolmogorov’s example (a survey of actions of infinite-dimensional groups with an invariant probability measure). Theory of Probability and its Applications, 48(2): 373–378, June 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98043.

Vershik:2008:SEF

[VG08] A. M. Vershik and A. D. Gorbulsky. Scaled entropy of filtrations of σ-fields. Theory of Probability and its Applications, 52 (3):493–508, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vinogradov:2007:JDR

[Vin07] O. P. Vinogradov. On a joint distribution of the ruin time and the number of the payment which leads to a ruin in a nonhomogeneous risk process. Theory of Probability and its Applications, 51(3):504–512, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Viskov:2001:RWS

[Vis01] O. V. Viskov. A random walk with a skip-free component and the Lagrange inversion formula. Theory of Probability and its Applications, 45(1): 164–172, March 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97810.

Vakhania:2007:UCW

[VK07] N. N. Vakhania and V. V. Kvaratskhelia. Unconditional convergence of weakly sub-Gaussian series in Banach spaces. Theory of Probability and its Applications, 51(2): 305–324, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Veretennikov:2009:CTE

[VK09] A. Yu. Veretennikov and M. L.

Kleptsyna. On continuous time ergodic filters with wrong initial data. Theory of Probability and its Applications, 53(2):269–300, ???? 2009. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vlasenko:2005:LVM

[Vla05] M. A. Vlasenko. Local visitation measures for some sequences of random variables with decreasing coefficients. Theory of Probability and its Applications, 49(1):176–186, March 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98091.

Valkeila:2000:MMS

[VM00] E. Valkeila and A. V. Melnikov.

Martingale models of stochastic approximation and their convergence. Theory of Probability and its Applications, 44(2): 333–360, June 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97754.

vanRooij:2000:AEI

[vRRvZ00] A. C. M. van Rooij, F. H.

Ruymgaart, and W. R. van Zwet. Asymptotic efficiency of inverse estimators. Theory of Probability and its Applications, 44(4):722–738, December 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97791.

Vatutin:2005:LTC

[VT05] V. A. Vatutin and V. A. Topchii. Limit theorem for critical catalytic branching random walks. Theory of Probability and its Applications, 49(3):498– 518, September 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98121.

Vershik:2007:MMY

[VT07] A. M. Vershik and N. V. Tsilevich. Markov measures on Young tableaux and induced representations of the infinite symmetric group. Theory of Probability and its Applications, 51(1):211–223, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vatutin:2008:CGW

[VWF08] V. A. Vatutin, V. Wachtel, and K. Fleischmann. Critical Galton–Watson processes: The maximum of total progenies within a large window. Theory of Probability and its Applications, 52(3):470–492, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Vysotsky:2006:ECS

[Vys06] V. V. Vysotsky. On energy and clusters in stochastic systems of sticky gravitating particles. Theory of Probability and its Applications, 50(2):265– 283, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Wachtel:2008:LTP

[Wac08] V. I. Wachtel. Limit theorems for probabilities of large deviations of a critical Galton– Watson process having power tails. Theory of Probability and its Applications, 52(4):674–688, ???? 2008. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Weber:2004:ERS

[Web04] M. Weber. Ergodic regularity of some Donsker classes. Theory of Probability and its Applications, 48(4):681–696, December 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98074.

Weber:2006:CMU

[Web06] M. Weber. On the CLT for means under the rotation action. I. Theory of Probability and its Applications, 50(4): 631–649, January 2006. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Weber:2007:CMU

[Web07] M. Weber. On the CLT for means under the rotation action. II. Theory of Probability and its Applications, 51(2): 377–387, ???? 2007. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yakymiv:2003:ADI

[Yak03] A. L. Yakymiv. On the asymptotics of the density of an infinitely divisible distribution at infinity. Theory of Probability and its Applications, 47(1): 114–122, March 2003. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97946.

Yakymiv:2004:TTA

[Yak04] A. L. Yakymiv. Tauberian theorems and asymptotics of infinitely divisible distributions in a cone. Theory of Probability and its Applications, 48(3):493– 505, September 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98053.

Yakymiv:2008:LTG

[Yak08] A. L. Yakymiv. Limit theorem for the general number of cycles in a random Apermutation. Theory of Probability and its Applications, 52 (1):133–146, ???? 2008. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yarykin:2007:SNS

[Yar07] P. N. Yarykin. Stability of the nonlinear stochastic process that approximates the system of interacting Brownian particles. Theory of Probability and its Applications, 51(2):387–396, ???? 2007. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yaroslavtseva:2009:NEB

[Yar09] L. S. Yaroslavtseva. Nonclassical error bounds for asymptotic expansions in the Central Limit Theorem. Theory of Probability and its Applications, 53 (2):365–367, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Yengibarian:2005:REM

[Yen05] N. B. Yengibarian. A renewal equation in a multidimensional space. Theory of Probability and its Applications, 49(4):737– 744, December 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/98136.

Zaitsev:2001:MVR

[Zai01] A. Yu. Zaitsev. Multidimensional version of a result of Sakhanenko in the invariance principle for vectors with finite exponential moments. I. Theory of Probability and its Applications, 45(4):624–641, December 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97855.

Zaitsev:2002:MVRa

[Zai02a] A. Yu. Zaitsev. Multidimensional version of a result of Sakhanenko in the invariance principle for vectors with finite exponential moments. II. Theory of Probability and its Applications, 46(3):490–514, September 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97912.

Zaitsev:2002:MVRb

[Zai02b] A. Yu. Zaitsev. Multidimensional version of a result of Sakhanenko in the invariance principle for vectors with finite exponential moments. III. Theory of Probability and its Applications, 46(4):676–698, December 2002. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97930.

Zanzotto:2004:RCS

[Zan04] P. A. Zanzotto. Representation of a class of semimartingales as stable Integrals.Part 2. Theory of Probability and its Applications, 48(1):181–188, March 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/98016.

Zaporozhets:2006:ERP

[Zap06] D. N. Zaporozhets. An example of a random polynomial with unusual zeros behavior. Theory of Probability and its Applications, 50(3):529–535, January 2006. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhdanov:2009:FAM

[Zhd09] D. A. Zhdanov. Filtered arithmetic mean measure and its applications. Theory of Probability and its Applications, 53 (2):368–379, ???? 2009. CO-DEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

Zhukov:2000:ANA

[Zhu00a] Yu. V. Zhukov. On the accuracy of normal approximation for the densities of sums of independent identically distributed random variables. Theory of Probability and its Applications, 44(4):785– 793, December 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97794.

Zhulenev:2000:LD

[Zhu00b] S. V. Zhulenev. On large deviations, I. Theory of Probability and its Applications, 44(1): 75–92, March 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/97732.

Zhulenev:2005:LDI

[Zhu05] S. V. Zhulenev. On large deviations, II. Theory of Probability and its Applications, 49(4):671– 690, December 2005. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98130.

Zielinski:2001:RSS

[Zie01] R. Zielinski. A reparameterization of the symmetric αstable distributions and their dispersive ordering. Theory of Probability and its Applications, 45(2):357–358, June 2001. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/97831.

Zaigraev:2004:ATL

[ZN04] A. Yu. Zaigraev and A. V. Nagaev. Abelian theorems, limit properties of conjugate distributions, and large deviations for sums of independent random vectors. Theory of Probability and its Applications, 48(4):664– 680, December 2004. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

                                        URL http://epubs.siam.org/
                                        sam-bin/dbq/article/98071.

Zubkov:2000:IBS

[Zub00] A. M. Zubkov. Isomorphisms between sets of probability distributions. Theory of Probability and its Applications, 44(4):794– 797, December 2000. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic).

URL http://epubs.siam.org/
sam-bin/dbq/article/97795.